### v1

# Neural Audio Codecs for Prompt-Driven Universal Source Separation

Abstract
[…]
1 Introduction
In this manuscript, we propose CodecSep, a novel framework for text-conditioned source separation that combines the interpretability of prompt-driven separation with the compute efficiency of neural audio codecs (NACs). CodecSep is the first model to bridge NACs with universal source separation (USS), enabling flexible extraction of arbitrary sound sources based on natural language queries. By conditioning a transformer-based masker on text embeddings generated via the text branch of CLAP—a multimodal encoder trained on audio-text pairs Wu et al. (2023)—CodecSep performs state-of-the-art (SOTA) separation directly in the NAC encoder latent space, introducing semantic control while maintaining very low compute cost.
The growing demand for flexible, real-time audio processing in bandwidth-constrained or on-device scenarios calls for generalizable and computationally efficient methods. Traditional separation models tackle the challenge of disentangling sound sources from complex mixtures Vincent et al. (2018) but are typically specialized for narrow domains such as speech or music. USS broadens this scope by separating arbitrary sources without reliance on predefined labels. Within this setting, text-guided source separation has emerged as a compelling approach, enabling users to specify the target source via natural language prompts—offering greater transparency, flexibility, and accessibility in real-world applications.
Recent advances have embraced this direction by integrating textual prompts into separation pipelines. Architectures based on the Conv-TasNet encoder-masker-decoder blueprint Luo and Mesgarani (2019) have achieved strong results in domain-specific tasks Chen et al. (2020b); Subakan et al. (2021); Li et al. (2023). Systems like AudioSep Liu et al. (2024) extend this framework to the text-conditioned setting by encoding text queries with models such as BERT Devlin et al. (2019) or CLAP Wu et al. (2023) and injecting semantic information through Feature-wise Linear Modulation (FiLM) layers Perez et al. (2018). These modifications allow intermediate representations to adapt dynamically to prompt semantics, steering the model towards the desired source.
However, such models, trained with signal-level objectives like scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019), are not only highly sensitive to compression artifacts but also tend to be large and computationally intensive. This limits their practicality for on-device or real-time applications where resources are constrained. As a result, they are often deployed via the cloud, introducing additional latency and infrastructure costs.
Neural audio codecs (NACs) offer a promising alternative. Models like SoundStream Zeghidour et al. (2022), Encodec Défossez et al. (2022), and DAC Kumar et al. (2023) compress audio into discrete token sequences using encoder-quantizer-decoder pipelines. Leveraging Residual Vector Quantization (RVQ), they enable adaptive bitrate control and efficient representation. These codecs are trained on diverse audio data using a combination of adversarial and perceptual losses, which promotes both fidelity and robustness under compression. The resulting token representations have proven useful in tasks such as audio generation for multimodal models Wang et al. (2024); Du et al. (2024) and language-conditioned synthesis Borsos et al. (2023); Wang et al. (2023). Building on this foundation, recent efforts like CodecFormer Yip et al. (2024b) and SDCodec Bie et al. (2024) have explored incorporating NACs into source separation pipelines. While these approaches achieve lightweight, high-fidelity performance, they remain limited to separating fixed types of sources (e.g., speech separation).
CodecSep addresses this gap by combining the computational efficiency of NACs with the flexibility of language-guided separation. It follows a NAC-based encoder-decoder architecture and integrates a transformer-based masker conditioned on CLAP-derived text embeddings. These embeddings generate FiLM parameters that modulate each transformer layer, enabling the model to selectively extract sources that align with the user’s query. Operating entirely in the codec latent space, CodecSep delivers interpretable, high-performance separation under strict computational constraints.
We evaluate CodecSep for universal source separation under a comprehensive set of protocols: (i) dnr-v2 for in-domain text-guided separation; (ii) five open-domain corpora—AudioCaps, ESC-50, Clotho-v2, AudioSet-eval, and VGGSound—for cross-domain generalization; (iii) three prompt granularities (fixed-stem, generic three-stem {“speech”, “music”, “sfx”}, and a universal setting with fine-grained, compositional SFX descriptions); and (iv) paraphrase robustness via unseen synonyms for speech/music. We also include an architectural ablation contrasting decoder-style source generation (CodecFormer Yip et al. (2024b)) with our Transformer masker over codec latents, plus a reconstruction diagnostic (Appendix) to probe source leakage and mixture consistency.
[…]
2 Related Work
2.1 Audio Source Separation
Notable approaches like DPTNet Chen et al. (2020b), SepFormer Subakan et al. (2021), and TDANet Li et al. (2023) follow the encoder-masker-decoder framework. In this setup, the encoder generates STFT-like latent representations. At the same time, a masker network predicts source masks, which are then applied elementwise to the encoder outputs to obtain source-specific latent representations. These estimates are further processed by a decoder, which mirrors the encoder architecture, to reconstruct the source audio. TDANet Li et al. (2023), the latest iteration, enhances this framework with a top-down attention mechanism, using global and local attention modules to hierarchically model multi-scale acoustic features. CodecSep also adopts this encoder-masker-decoder structure, extending it to operate in the NAC latent space with text guidance.
Beyond the encoder-masker-decoder paradigm, a variety of alternative architectures have been explored. Models like Wave-UNet Stoller et al. (2018) and Demucs Défossez et al. (2019, 2021) utilize encoder-decoder designs that estimate source waveforms directly, omitting the masking stage altogether. Other approaches operate in the frequency domain, notably, models like MM-DenseLSTM Takahashi et al. (2018), DCCRN Hu et al. (2020), and UNet-based architectures such as Spleeter Hennequin et al. (2020) perform separation directly in the complex STFT domain, jointly modelling magnitude and phase components. These diverse strategies highlight the broad design space of source separation, spanning both time and frequency domains, with and without masking.
2.2 Universal Source Separation
Prior work on source separation has concentrated mainly on speech or music-specific tasks. More recent advances have shifted toward USS, which aims to disentangle arbitrary sources from a mixture without assuming predefined source categories. Supervised USS models often employ Permutation Invariant Training (PIT) Yu et al. (2017); Kavalerov et al. (2019), while unsupervised approaches like Mixture Invariant Training (MixIT) Wisdom et al. (2020) learn directly from mixtures without access to clean references. However, both approaches assume a fixed maximum number of sources in the mixture and produce all outputs indiscriminately. As a result, an additional source identification step is needed post-separation to determine which of the estimated signals corresponds to the target source.
Query-Guided Source Separation (QSS) extends USS by enabling source extraction through external queries, including visual input, exemplar audio, class labels, or text. Text queries are particularly effective, offering a compact and expressive way to convey high-level semantics and encode spatial or temporal cues of the query source—all without the need for expensive reference data. Recent work such as AudioSep Liu et al. (2024) explores text-guided separation by integrating embeddings from models like BERT Devlin et al. (2019) and CLAP Wu et al. (2023), the latter of which aligns audio and text in a shared embedding space. AudioSep employs FiLM conditioning into the separation pipeline at intermediate layers of the masker. CodecSep adopts a similar strategy but operates in the NAC encoder latent space instead.
2.3 Source Separation with Neural Audio Codecs
Recent work has explored integrating neural audio codecs into source separation pipelines. CodecFormer Yip et al. (2024b) and its extension CodecFormer-EL Yip et al. (2024a) both leverage DAC as the encoder-decoder backbone, performing separation directly in the codec’s latent embedding space via a transformer-based separation module. The original CodecFormer is trained using a negative SI-SDR loss to supervise the predicted outputs. At the same time, CodecFormer-EL introduces an embedding-level objective that minimizes the distance between the separator’s output and the DAC encoder’s ground truth latent embeddings for each source. In contrast, SD-Codec Bie et al. (2024) integrates source separation more tightly into the codec itself. It is explicitly designed to disentangle mixtures into speech, music, and SFX. To achieve this, it introduces three dedicated RVQ branches—each trained to encode a specific source category. During encoding, the outputs of these domain-specific quantizers are summed across time to form a mixture-level representation, which is then decoded to reconstruct both the whole mixture and individual sources. This design enables SD-Codec to embed separation behavior directly into the codec’s structure. In this work, we repurpose the CodecFormer transformer separator into a text-conditioned mask predictor, enabling prompt-guided control over the separation process.
Figure 1: An overview of CodecSep. (Left) The full pipeline for text-guided USS. (Right) The integration of text conditioning into intermediate layers of transformer masker via FiLM layers.
3 Method
CodecSep is based on the 16kHz variant of CodecFormer Yip et al. (2024b), which we adapt to enable text-driven source separation. Specifically, we modify the transformer module to function as a source mask estimator and incorporate Feature-wise Linear Modulation (FiLM) Perez et al. (2018) to inject text-based conditioning into the separation process. FiLM layers are applied to the transformer’s intermediate activations, allowing the model to adjust its internal representations based on the semantics of the input query. To produce the FiLM parameters, a lightweight query network consumes embeddings produced by a pretrained CLAP encoder Wu et al. (2023) from the input prompt.
Figure 1 (Left) illustrates our text-guided USS pipeline. Unlike AudioSep’s raw STFT-based, compute-heavy architecture, CodecSep operates on compact codec representations, offering significantly greater compute efficiency for resource-constrained edge deployment. Figure 1 (Right) illustrates how text-conditioning is integrated into the intermediate encoder layers of the transformer masker.
3.1 Model Architecture
CodecSep utilizes DAC Kumar et al. (2023) for the encoder-decoder modules and a transformer-based masker network with text conditioning to disentangle query sources within the DAC encoder representation space. We will discuss each component in detail and present how we incorporated text conditioning within the transformer masker architecture.
3.1.1 Codec Backbone
DAC Kumar et al. (2023) builds upon the Encodec Défossez et al. (2022) and SoundStream Zeghidour et al. (2022) architectures and utilizes fully convolutional encoder and decoder networks. Leaky ReLU activations in the architectures are substituted with the periodic Snake Activation function Ziyin et al. (2020) expressed as $x+\sin^{2}{x}$. This enhances the modeling of periodic audio signals by incorporating a periodic inductive bias. Residual vector quantization is integrated into the encoder output, which enables audio compression at varying bit rates. In addition, factorized codes—where code assignment occurs in a lower-dimensional space than the codebook vectors—and L2-normalized codebook vectors are utilized to enhance codebook efficiency.
Let us consider compressing an audio recording $x(t)$ of duration $1$s, sampled at $F_{s}=24$kHz to target bit-rate of R=6000R=6000bps. The encoder network $Enc(\cdot)$ down-samples $x(t)$ by a factor of $M=320$ and outputs $T=\frac{F_{s}}{M}=75$ frames of latent representations, denoted as $Z=[z_{t}\in R^{d}|t\in[T]]$. Each frame $z_{t}$ is allocated $r=\frac{R}{T}=80$ bits. Residual vector quantization (RVQ) efficiently allocates the bit-rate budget across a cascade of $N_{q}=8$ vector quantizers. The first vector quantizer in the cascade quantizes the frame $z_{t}$, and the residuals (the quantization error) are computed. The residuals are then iteratively quantized by the remaining $N_{q}-1$ quantizers in the cascade. Each quantizer is assigned $r_{i}=\frac{r}{N_{q}}=10$ bits, maintaining a manageable code-book size of $2^{10}=1024$. The number of quantizers in the cascade $N_{q}$ is key in balancing computational complexity and compression for a given bit-rate budget $r$.
The RVQ network $Quant(\cdot)$ produces a code matrix $A=[a_{t}\in[1024]^{8}|t\in[T]]$ given encoder output $Z$. Each frame $z_{t}\in Z$ is quantized into a code vector $a_{t}\in[1024]^{8}$, which can be mapped to an embedding via code-book lookup as $e_{t}=\sum_{i=1}^{8}e^{i}_{t}$, where $e^{i}_{t}$ corresponds to the embedding associated with code $a^{i}_{t}\in[1024]$ from the $i^{th}$ RVQ code-book. The RVQ output embedding sequence $E=[e_{t}|t\in[T]]$ is then processed by the decoder network $Dec(.)$, which up-samples it to generate $y(t)$, a reconstruction of the original audio $x(t)$.
3.1.2 Transformer Masker
CodecSep masker module $Mask(.)$ predicts source mask given DAC Kumar et al. (2023) encoder $Enc(.)$ representation $Z$. These masks are applied element-wise to representation $Z$ to extract the corresponding source representation estimate. It follows CodecFormer Yip et al. (2024b) Transformer separator module architecture and is designed as a stack of $L=16$ transformer layers $Mask^{l\in[L]}(.)$ of embedding size $d=256$. The output of the final transformer layer, $H^{L}=Mask^{L}(H^{L-1})$, is processed through $K$ distinct 1D convolutional layers to generate source masks $\{M_{k}|k\in[K]\}$. To preserve the periodic inductive bias across the entire model, the Snake activation function, employed in DAC Kumar et al. (2023), replaces tanh and LeakyReLU activations throughout the masker module. The resulting source representation estimates are subsequently mapped to time domain $\{\tilde{y}_{k}=Dec(M_{k}\odot Z)|k\in[K]\}$ using DAC Kumar et al. (2023) decoder $Dec(.)$ directly, without being processed by RVQ network $Quant(.)$.
3.1.3 Text-guided Source Separation
We consider a single-channel audio mixture $x(t)\in\mathbb{R}^{1\times N}$, which consists of an unknown number of overlapping open-domain sound sources. The mixture is expressed as the following sum:
$x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$ (1)
where $\mathcal{S}$ is the (potentially unbounded) set of sound source classes or instances present in the mixture. Our objective is to recover a target source waveform or stem $y_{s}(t)$ from the mixture $x(t)$, using a natural language query $\tau_{s}$ that semantically describes the desired source. These prompts may reference specific sources (e.g., “dog barking,” “violin solo,” “crowd cheering”) or combinations of sources (e.g., “speech and music,” “footsteps and glass breaking”), allowing for fine-grained control.
To integrate text-based conditioning into CodecSep, we leverage CLAP Wu et al. (2023) and utilize a masker module $Mask(.)$ with $K=1$ final convolution to generate a single source mask $M_{s}$. Given a query text $\tau$ that describes the target source, CLAP Wu et al. (2023) generates a corresponding text embedding $e_{\tau}\in\mathbb{R}^{d}$. This embedding is used to modulate the intermediate $l\in[2,...,L-1]$ transformer layers $Mask^{l}(.)$ via feature-wise linear modulation (FiLM) Perez et al. (2018), enabling the network to estimate the desired source conditioned on the text input.
At transformer layer $l$, the layer output $H^{l}\in\mathbb{R}^{d\times T}$ is modulated along the feature dimension using parameters $\gamma^{l},\beta^{l}\in\mathbb{R}^{d}$. These modulation parameters are generated by a query network $query(.)$, which processes text embedding $e_{\tau}$ to produce $(\gamma^{l},\beta^{l})_{l=1}^{L}=query(e_{\tau})$. The modulation transformation is then carried out as follows:
$\tilde{H}^{l}_{s}=FiLM(H^{l}_{i},\gamma^{l}_{i},\beta^{l}_{i})=\gamma^{l}_{i}H^{l}_{i}+\beta^{l}_{i},\forall i\in[d]$ (2)
The decoder network $Dec(.)$ subsequently processes the source representation estimate $\tilde{Z}_{s}=M_{s}\odot Z$ and generates an estimate of the query source audio $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$.
3.2 Training Objective
We train CodecSep to reconstruct each individual source from mixtures using textual prompts corresponding to broad categories such as music, speech, and a wide range of sound effects (SFX). We also reconstruct the input mixture $x$ by decoding the summed source latent estimates: $\tilde{x}=g(\sum_{s}\tilde{Z}_{s})$. The model is optimized using the negative scale-invariant signal-to-distortion ratio (SI-SDR) loss Luo and Mesgarani (2019); Le Roux et al. (2019), encouraging the faithful separation of sources aligned with the given queries.
To support generalization across varied SFX types, prompts associated with SFX are not limited to atomic events—they may describe composite scenarios involving multiple overlapping sources (e.g., “footsteps, background traffic,” “crowd cheering, fireworks”), requiring the model to isolate multi-faceted acoustic scenes based on complex queries.
SI-SDR is designed to measure the similarity between the predicted $\tilde{v}$ and ground truth $v$ signals, focusing on the signal-to-noise ratio (SNR) while being invariant to scale differences. It is given as,
[…]
The overall loss function is then given as:
[…]
To account for compression artifacts and improve training stability, we further compute the negative codec SI-SDR (cSI-SDR) Yip et al. (2024b), which is a variant of SI-SDR calculated with a codec-processed version of ground truth audio. During training, we keep the DAC Kumar et al. (2023) backbone and CLAP Devlin et al. (2019) text encoder weights frozen, updating only the parameters of $Mask(.)$ and $query(.)$ modules.
3.3 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
Problem setup and pipeline contrast.
Let $x(t)\in\mathbb{R}$ be a mono mixture with sources $\{y_{s}(t)|s\in\mathcal{S}\}$, where $x(t)$ is expressed as equation 1.
Spectrogram-domain (AudioSep):
$x\;\xrightarrow{\text{STFT}}\;X\!\in\!\mathbb{C}^{F\times T}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\odot|X|\exp(\angle X+\angle\hat{M}_{s})\;\xrightarrow{\text{ISTFT}}\;\tilde{y}_{s}(t),$
where $g(\cdot,e_{\tau})$ denotes a FiLM-conditioned, complex-domain spectrogram separator that predicts a magnitude mask $|\hat{M}_{s}|\in[0,1]^{F\times T}$ and a phase residual $\angle\hat{M}_{s}$, conditioned jointly on the mixture spectrogram $X$ (obtained by the STFT of $x$) and the text embedding $e_{\tau}$. The predicted magnitude mask and phase residual are then applied element-wise to form the source spectrogram $\tilde{Y}_{s}$ which is subsequently transformed back to the time domain via ISTFT to obtain $\tilde{y}_{s}(t)$.
CodecSep (NAC latent-domain):
$x\;\xrightarrow{\,Enc(\cdot)\,}\;Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}\;=\;M_{s}\odot Z\;\xrightarrow{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
with a frozen DAC backbone $Enc(\cdot),Dec(\cdot)$, and a FiLM-conditioned transformer masker $Mask(\cdot,e_{\tau})$ that estimates a soft mask $M_{s}\in[0,1]^{d\times T}$ conditioned on both the codec latents $Z=Enc(x)$ and the text embedding $e_{\tau}$. The predicted mask $M_{s}$ is applied element-wise to $Z$ to produce source-specific latents $\tilde{Z}_{s}=M_{s}\odot Z$, which are subsequently decoded by $Dec(\cdot)$ to obtain the separated waveform $\tilde{y}_{s}(t)$.
Dimensionality reduction and compression.
For a typical 1 s audio, 32 kHz STFT used by spectrogram-domain baselines, the mixture spectrogram $X\in\mathbb{C}^{F\times T}$ (e.g., $F=2\cdot 1024$, $T\approx 100$) contains about $|X|\approx 204{,}800$ complex coefficients. In contrast, a neural audio codec latent $Z\in\mathbb{R}^{d\times T_{c}}$ is far more compact: at 16 kHz with $d=64$ and $T_{c}=\tfrac{16000}{320}\approx 50$, |Z|=3,200|Z|=3{,}200; at 32 kHz with $T_{c}=\tfrac{32000}{320}\approx 100$, |Z|=6,400|Z|=6{,}400. Thus, the ratio $|Z|/|X|$ is $\approx 1.5625\times 10^{-2}$ for 16 kHz and $\approx 3.125\times 10^{-2}$ for 32 kHz , implying a $\sim 32$$64\times$ reduction in representation size. Operating directly in $Z$ therefore cuts memory bandwidth and MACs for the separator $Mask(\cdot)$ while retaining perceptually salient structure imparted by the codec, making latent-domain separation both compute-efficient and semantically aligned. Moreover, the compact channel dimension ($d=64$) enables a smaller masker (fewer parameters in $Q/K/V$ and MLP projections) and more effective self-attention.
3.3.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
STFT vs. NAC encoding. The STFT is a linear projection from $x(t)$ to $X\in\mathbb{C}^{F\times T}$ and does not explicitly preserve or organize the intrinsic semantic structure of audio. Consequently, spectrogram-based separators (e.g., AudioSep) require an additional learned encoder–decoder (typically CNN/ResUNet) to first compress $X$ into high-dimensional latent features and then decode spectrogram masks that are discriminatively trained for separation. This couples semantic abstraction and separation into one network, increasing parameters and MACs and forcing the model to learn structure “from scratch” in a redundant, noisy representation.
NAC encoder as a semantic prior. Neural audio codecs (DAC) are trained with perceptual, adversarial, and codebook objectives that encourage the encoder $Enc(\cdot)$ to map $x$ into compact, structured latents:
$Enc(.):x\mapsto Z\in\mathbb{R}^{d\times T_{c}}.$
These latents lie on a discriminative compressed manifold $\mathcal{M}_{\text{latent}}$ in which semantically meaningful factors (e.g., pitch, timbre, transients) are disentangled and aligned for downstream use. In our system, we operate directly on the continuous latents $Z$ (from $Enc$), use a FiLM-conditioned transformer masker to predict a separation mask $M_{k}$ from a textual query $\tau_{k}$, and form the estimate
$\tilde{Z}_{k}\;=\;M_{k}\odot Z,\qquad\tilde{y}_{k}\;=\;Dec(\tilde{Z}_{k}),$
thereby leveraging the codec’s structured manifold for masking instead of learning a new representation.
Separation mapping in latent vs. spectrogram space. In CodecSep, the separator learns
$Mask(.,.):\;Z\in\mathcal{M}_{\text{latent}}\;\to\;\hat{Z}_{k},$
which is easier to optimize because the input is denoised, compressed, and semantically organized by the codec. In contrast, spectrogram models must learn
$g(.,.):\;X\in\mathbb{C}^{F\times T}\;\to\;\hat{Y}_{k},$
over a noisier, higher-dimensional space without semantic compression.
Hierarchical (RVQ) structure that benefits separation. The codec applies Residual Vector Quantization (RVQ) to $Z$, producing discrete codes $A=[a_{t}\in[K]^{N_{q}}]_{t=1}^{T_{c}}$ with $K=1024$ and $N_{q}$ quantizers. Codebook lookup yields embeddings
$e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}(a_{t}^{(i)}),\qquad E=[e_{t}|t\in[T]]\approx Z,$
and $Dec(E)$ reconstructs the waveform. The RVQ cascade induces a natural coarse-to-fine hierarchy: the first quantizer captures coarse structure (e.g., low-frequency content, speaker/instrument timbre, global acoustic traits), while later quantizers refine residual details (e.g., high-frequency components, onsets/transients, background textures). This hierarchy mirrors the discriminative cues needed for separation and is directly exploitable by a mask-based transformer.
Loss-induced organization of the latent space. The DAC objective shapes $Z$ using complementary terms:
  • Multi-scale spectral loss $\mathcal{L}_{\text{mel}}$ to preserve perceptually relevant frequency content at multiple time scales;
  • Time-domain reconstruction loss $\mathcal{L}_{\text{feat}}=\|x(t)-\tilde{y}(t)\|_{1}$ for fidelity and stability;
  • Multi-resolution adversarial loss $\mathcal{L}_{\text{adv}}$ with (i) multi-period waveform discriminators (pitch/periodicity) and (ii) multi-band STFT discriminators (fine spectral detail), plus a feature-matching term $\mathcal{L}_{\text{feat}}^{G}$;
  • Codebook loss $\mathcal{L}_{\text{code}}$ to ensure compact, diverse, well-utilized codes and reinforce RVQ’s coarse-to-fine disentanglement.
  • Quantizer dropout (RVQ stage dropout): randomly disabling a subset of RVQ stages during training to discourage over-reliance on late codebooks and encourage smoother coarse-to-fine residual allocation, yielding more bitrate-robust and well-structured RVQ representations.
The overall loss is
$\mathcal{L}_{\text{DAC}}\;=\;\lambda_{\text{mel}}\mathcal{L}_{\text{mel}}\;+\;\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}\;+\;\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}\;+\;\lambda_{\text{code}}\mathcal{L}_{\text{code}},$
yielding latents that are (i) denoised (robust to low-level artifacts), (ii) semantic (preserve pitch, timbre, temporal structure), (iii) disentangled (coarse-to-fine RVQ), and (iv) efficient (bitrate-aware constraints).
Separation benefits from the codec prior (and contrast to spectrogram baselines). Because $Z$ is already semantically organized, the FiLM-conditioned masker operates on a representation that encodes the right factors for selection, not generation. This (i) reduces compute and memory (the separator acts on compact $Z$ instead of $X$), (ii) accelerates convergence and improves stability (masking over a clean manifold), and (iii) improves robustness to prompt variation (text FiLM modulates semantically aligned channels). In contrast, spectrogram systems must learn a task-specific latent from $X$ and perform separation jointly, which increases parameter count and MACs, slows convergence, and can lead to overfitting in the absence of the codec’s inductive bias.
3.3.2 Transformer-based masker vs. decoder-style generation (specific design choice).
In CodecSep, we replace decoder-style source generation (as in CodecFormer, which directly predicts the target waveform/spectrogram) with a Transformer-based, FiLM-conditioned masker that outputs a soft mask over the codec latents. Concretely,
$M_{s}\;=\;Mask(Z,e_{\tau})\in[0,1]^{d\times T},\qquad\tilde{Z}_{s}\;=\;M_{s}\odot Z,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{Z}_{s}).$
This choice has the following concrete advantages (all in the NAC latent domain):
  • Efficient and stable training. Predicting $M_{s}$ to modulate existing latent content is a simpler, more constrained learning problem than end-to-end generation of $\tilde{Z}_{s}$ or $\tilde{y}_{s}(t)$ with a decoder head. Working in $Z$ avoids the instability commonly observed in direct waveform/spectrogram prediction, leading to faster convergence and lower training variance under the same optimization settings.
  • Minimal distortion through modulation (no hallucination). The separator does not synthesize new content; it rescales and selects what is already encoded in $Z$. Forming $\tilde{Z}_{s}=M_{s}\odot Z$ preserves the mixture’s latent structure and reduces artifacts relative to encoder–decoder separation pipelines that generate source signals from scratch, thereby limiting hallucinations and leakage.
  • Preservation of long-term temporal/spectral structure. The NAC encoder has already organized periodicity, timbre, and transient structure in $Z$. Masking retains this organization across long contexts, whereas fully convolutional decoders trained to generate sources often exhibit long-term inconsistencies (e.g., drift over time or loss of periodic cues) when reconstructing from scratch.
  • Efficient use of Transformer capacity. Instead of synthesizing source signals, the masker learns to gate semantically organized channels in $Z$, which is a substantially lighter optimization problem than decoder-style generation. FiLM conditioning steers the transformer to decide where/how much information to pass—not what to generate—so parameters and compute in the $Q/K/V$ and MLP projections are concentrated on selection and attenuation.
3.3.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
We deliberately integrate Feature-wise Linear Modulation (FiLM) Perez et al. (2018) inside the transformer masker to inject text semantics while preserving the codec manifold and keeping $Enc(\cdot)$$Dec(\cdot)$ frozen.
Targeted placement (masker, mid-layers).
Given a CLAP text embedding $e_{\tau}$, a lightweight query network $query(\cdot)$ produces per-layer affine parameters
$(\gamma^{l},\beta^{l})_{l=1}^{L}\;=\;query(e_{\tau}),\quad\gamma^{l},\beta^{l}\in\mathbb{R}^{d},$
which modulate intermediate activations $H^{l}\in\mathbb{R}^{d\times T}$ for $l\in\{2,\dots,L-1\}$:
$\tilde{H}^{l}_{s}\;=\;\mathrm{FiLM}\!\big{(}H^{l};\gamma^{l},\beta^{l}\big{)}\;=\;\gamma^{l}\odot H^{l}+\beta^{l}.$
Placing FiLM in the masker (not in $Enc$ or $Dec$) keeps the codec latent distribution intact and confines conditioning to the selection step.
Lightweight computation (overhead and parameterization).
FiLM adds only small per-layer vectors $(\gamma^{l},\beta^{l})$ and a compact $query(\cdot)$ MLP; it does not increase sequence length, attention heads, or the quadratic attention cost. The extra FLOPs/params are negligible relative to multi-head attention and MLP blocks, aligning with CodecSep’s efficiency goals.
Non-iterative inference (single forward pass).
FiLM applies in a single pass through the masker. Unlike iterative conditioning mechanisms (e.g., flow-matching-based sampling), there are no sampling steps, thus preserving low latency for edge and hybrid deployments.
Manifold preservation and stability.
Because FiLM scales/shifts existing channels of $H^{l}$ rather than rewriting $Z$ or generating $\tilde{Z}_{s}$ from scratch, the NAC manifold structure (periodicity, timbre, transients) is preserved. Empirically this reduces training variance and mitigates long-horizon inconsistencies common with generator-style heads.
3.3.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
Why we evaluate on continuous latents $Z$.
We choose to perform—and therefore report—separation on the continuous DAC encoder latents $Z=Enc(x)\in\mathbb{R}^{d\times T_{c}}$ for the following concrete reasons:
  1. 1.
    End-to-end gradient flow for separation. Our training updates only the masker and query networks while keeping the codec frozen. Using $Z$ allows straightforward backpropagation through $Mask(\cdot,e_{\tau})$ and $Dec(\cdot)$ without dealing with discrete indices or straight-through estimators; gradients are well-behaved and convergence is consistently stable.
  2. 2.
    Representational fidelity and disentanglement. During codec pretraining, the RVQ cascade regularizes $Z$ so that pitch, timbre, onsets/transients, and background textures are cleanly and hierarchically organized. $Z$ therefore provides a richer, more disentangled signal for FiLM-conditioned masking than hard code indices, which are subject to quantization coarsening.
  3. 3.
    Training stability and variance. In our setting (frozen codec), operating on $Z$ avoids variability from codebook utilization dynamics (e.g., late-stage RVQ sensitivity, bitrate truncation). Empirically, this reduces run-to-run variance and removes the need for specialized regularizers when training the separator.
How to extend the same model to discrete code streams.
For deployment scenarios where the input is a compressed bitstream, we operate on the reconstructed embeddings $E$ obtained from the codes $A$ via codebook lookup, and train the masker on $E$ instead of $Z$:
$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T_{c}]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big{(}a_{t}^{(i)}\big{)},$ (5)
$E$ $=[e_{t}]_{t=1}^{T_{c}}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (6)
Compressed bitstream path (codes-in, codes-out). Given $E\approx Z$ and the element-wise masking operation, the estimated source embeddings satisfy
$\tilde{E}_{s}\;=\;M_{s}\odot E\;\approx\;M_{s}\odot Z\;=\;Z_{s}$
When a uniform codec pathway (or codes-out interface) is required, we re-quantize the masked embeddings and optionally decode via the codec:
$\hat{A}_{s}\;=\;Quant(\tilde{E}_{s}),\qquad\hat{E}_{s}\;=\;\mathrm{lookup}(\hat{A}_{s}),\qquad\tilde{y}_{s}(t)\;=\;Dec\!\big{(}\hat{E}_{s}\big{)}$
In deployments that only need to return a bitstream, the server can emit $\hat{A}_{s}$ directly and defer decoding to the client; otherwise, decoding $Dec(\tilde{E}_{s})$ or $Dec(\hat{E}_{s})$ yields the waveform on-device or server-side, respectively.
Embedding alignment (CodecFormer-EL Yip et al. (2024a)). To tighten $\tilde{E}_{s}\approx Z_{s}$, we can optimize an embedding-level consistency loss as in CodecFormer-EL Yip et al. (2024a)instead of our SI-SDR objective:
$\mathcal{L}_{\mathrm{emb}}\;=\;\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$
Why this works. By design of the codec, $E$ approximates $Z$ at the operating bitrate; the decoder already reconstructs $\hat{x}(t)=Dec(E)$ with high fidelity. Since our separator is a masker (selective modulation) rather than a generator, replacing $Z$ with $E$ preserves the semantics needed for separation while enabling direct operation on bitstreams. In practice, the extension amounts to training (or fine-tuning) the same FiLM-conditioned transformer on $E$ with unchanged objectives.
Summary of scope. We present results on $Z$ to (i) isolate separator performance without discrete-index training complications, (ii) leverage continuous gradient flow and stable optimization, and (iii) align FiLM conditioning with a smooth latent manifold. The deployment-ready path to discrete code streams is immediate: switch the masker input to $E$ via lookup from $A$ and train/fine-tune accordingly, with optional distribution alignment and bitrate-robustness augmentation.
3.3.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
Central motivation. In realistic deployments, edge devices already run a neural audio codec; they transmit code streams rather than raw waveforms. CodecSep treats the codec backbone as part of the separation stack and operates directly on codec representations, thereby eliminating redundant decode $\rightarrow$ separate $\rightarrow$ re-encode cycles required by spectrogram-domain systems.
Pipeline comparison (server-side separation).
Traditional (spectrogram) pipeline:
$\underbrace{\text{Edge: }Enc(.)}_{\text{codec}}\ \Rightarrow\ \text{Code}\ \Rightarrow\ \underbrace{\text{Server: }Dec(.)+g(\cdot,e_{\tau})+Enc(.)}_{\text{decode + separate + re-encode}}\ \Rightarrow\ \text{Code}\ \Rightarrow\ \text{Edge: }Dec(.)$
CodecSep pipeline:
$\underbrace{\text{Edge: }Enc(.)}_{\text{codec}}\ \Rightarrow\ \text{Code}\ \Rightarrow\ \underbrace{\text{Server: }Mask(\cdot,e_{\tau})}_{\text{mask on }E\approx Z}\ \Rightarrow\ \text{Code}\ \Rightarrow\ \text{Edge: }Dec(.)$
CodecSep performs FiLM-modulated masking in the codec latent domain and returns separated code streams for edge-side decoding; spectrogram systems must decode to waveform (or magnitude/phase), separate in $X$, and re-encode.
Complexity accounting.
Let $C_{\mathrm{enc}}$ and $C_{\mathrm{dec}}$ be codec encode/decode costs, $C_{\mathrm{spec}}$ the spectrogram separator cost, and $C_{\mathrm{mask}}$ the CodecSep masker cost.
Code-stream input (typical): $\text{AudioSep: }C_{\mathrm{dec}}+C_{\mathrm{spec}}+C_{\mathrm{enc}},\qquad\text{CodecSep: }C_{\mathrm{mask}}.$
Audio-stream input (edge-only): $\text{AudioSep: }C_{\mathrm{spec}},\qquad\text{CodecSep: }C_{\mathrm{enc}}+C_{\mathrm{mask}}+C_{\mathrm{dec}}.$
In the common (code-stream) case, CodecSep removes both decode and re-encode on the server. Moreover, within the separator, CodecSep operates on $Z/E$ with $|Z|\ll|X|$ (cf. Sec. 3.3), reducing activation memory and bandwidth throughout attention and MLP blocks.
Interface compatibility with codec bitstreams.
When only quantized codes $A=[a_{t}|t\in[T]]$ are available, we reconstruct embeddings by codebook lookup $E=[e_{t}|t\in[T]]$ with $e_{t}=\sum_{i=1}^{N_{q}}\text{lookup}(a_{t}^{(i)})\approx Z$ and apply the same masker:
$A\ \Rightarrow\ E\approx Z\ \xrightarrow{\ Mask(\cdot,e_{\tau})\ }\ \tilde{E}_{s}\ \Rightarrow\ \text{Code stream out}.$
No architectural change is required; separation remains in the codec latent domain and stays fully compatible with streaming/edge ecosystems.
Why spectrogram-domain systems incur extra overhead.
Spectrogram separators (e.g., AudioSep) are defined on $X=\mathrm{STFT}(x)$. Given code-stream inputs, they must first run $Dec(.)$ to obtain a waveform, compute $X$, perform separation, and then run $Enc(.)$ to return codes. This decode + separate + re-encode loop adds latency, memory traffic, and energy cost on the server path and scales poorly with concurrent streams.
Operational advantages of CodecSep.
  • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec. 3.3.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
  • Smaller working representation during separation: masking in $Z/E$ (e.g., d=64d{=}64) reduces intermediate activations, lowering memory bandwidth and enabling tighter batching.
  • Non-iterative, single-pass conditioning: FiLM-conditioning within the masker adds negligible overhead and preserves low latency (no iterative sampling).
  • Seamless edge/server hybrid and edge-only modes: identical separator logic serves both modalities; with code-stream inputs, the server path remains separator-only.
  • Maintains codec manifold structure: by modulating $Z/E$ rather than rebuilding $X$, CodecSep preserves periodicity/timbre/transients already organized by the codec, supporting stable, high-fidelity stems at deployment.
4 Experiments
4.1 Datasets
4.1.1 Divide and Remaster v 2.0 (dnr-v2)
dnr-v2 Petermann et al. (2022) dataset consists of $60s$-duration artificial mixtures of speech, music, and SFX sampled from LibriSpeech Panayotov et al. (2015), Free Music Archive (FMA) Defferrard et al. (2017), and Freesound Dataset $50K$ (FSD50K50K) Fonseca et al. (2022), respectively. It includes $3,406$ ($56.7hrs$) training, $487$ ($8.13hrs$) validation, and $973$ ($16.22hrs$) test mixtures, each provided with its three individual source audios. The mixtures are generated by normalizing each source to fixed Loudness Units Full-Scale (LUFS) levels: 17-17 dB (speech), 24-24 dB (music), and 21-21 dB (SFX), with $\pm 2$ dB random perturbations. Any source exceeding a peak threshold is normalized to $0.5$ dB. The sources are mixed and normalized to 27-27 dB LUFS with additional random perturbations. The validation and test sets are trimmed for silence and split into $5s$ or $10s$ segments. Segments where sources are present for less than $50\%$ of the duration are removed, resulting in $2,852$ ($\approx 3.96hrs$) validation and $1,840$ ($\approx 5.11hrs$) test mixtures.
While originally developed for $3$-stem separation, we adapt dnr-v2 to the USS setting by replacing fixed source labels with natural language descriptions. For speech or music stem, we use broad, category-level prompts (e.g., “speech,” “music”), reflecting realistic usage in production workflows. In contrast, SFX sources are more complex—often containing three or more overlapping events. We generate prompts to query the SFX stem using FSD50K’s hierarchical annotations, combining fine-grained class labels with their parent categories. This results in long-form, compositional queries that reflect the structure of the mixture (e.g., “dog barking, Animal, engine rumbling, motor vehicle”).
4.1.2 Open-Domain Benchmarks
We benchmark on five open-domain datasets spanning captioned audio, environmental sounds, and large multi-event corpora: AudioCaps Kim et al. (2019), an AudioSet-derived collection of >46>46k $10$ s YouTube clips paired with human-written captions describing the dominant sound events (used by us to synthesize training and test mixtures); ESC-50 Piczak , a curated environmental sound dataset of 2,0002{,}000 clips ($5$ s each) organized into $50$ classes with $40$ examples per class across five meta-categories (animals, natural, human non-speech, domestic, exterior/urban); Clotho-v2 Drossos et al. (2020), 6,9746{,}974 audio samples (15153030 s) each annotated with five human captions (8–20 words) covering open-domain events; AudioSet-eval Gemmeke et al. (2017), the evaluation split of AudioSet comprising human-labeled $10$ s YouTube clips over an ontology of $632$ audio event classes in a multi-label setting; and VGGSound Chen et al. (2020a), an AudioSet-derived audio–visual corpus with 550+550{+} hours of $10$ s segments covering a wide variety of everyday sound categories. For AudioCaps we form both training and testing mixtures by summing three clips (validation segmented into $5$ s, test preserves clips up to $20$ s), while for ESC-50, Clotho-v2, AudioSet-eval, and VGGSound we construct test-only mixtures using the same three-clip protocol.
4.2 Evaluation
We evaluate CodecSep alongside TDANet Li et al. (2023); Pons et al. (2024), CodecFormer Yip et al. (2024b), SDCodec Bie et al. (2024), and AudioSep Liu et al. (2024) on universal source separation conducted on the dnr-v2 and AudioCaps test splits. We report the scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019) for signal quality assessment and Virtual Speech Quality Objective Listener (ViSQOL) Chinen et al. (2020) for perceptual quality assessment. ViSQOL Chinen et al. (2020) evaluates the perceptual similarity between the ground truth $x$ and the estimated signal $\tilde{x}$ using a spectro-temporal approach. It then predicts a Mean Opinion Score - Listening Quality Objective (MOS-LQO), which ranges from $1$ (lowest quality) to $5$ (highest quality). In addition, we evaluate the inference efficiency of our models in a hardware-agnostic setting by reporting Multiply-and-Accumulate operations (MACs) using Torchprofile  [^1] https://github.com/zhijian-liu/torchprofile.git averaged over 8 audios of $2s$ duration each.
4.3 Training
The complete model, including the query module $query(.)$, is trained for $400K$ iterations with DAC Kumar et al. (2023) and CLAP Wu et al. (2023) modules frozen. Validation is conducted every $5K$ iterations and test every $10K$ iterations. We use ADAM Kingma and Ba (2017) as our optimizer and train with a batch size of $4$ examples, each $2$ seconds in duration, and a learning rate of $1.5e^{-4}$ on a single 2424GB NVIDIA A-$30$ GPU. Training employs a $ReduceLRonPlateau$ Mukherjee et al. (2019) scheduler, which reduces the learning rate by a factor of $0.5$ if the validation loss does not improve for two consecutive validation checks. We train two versions of CodecSep, one using the dnr-v2 dataset and the other using AudioCaps, to evaluate performance across different training distributions. We refer to these models using the prefixes +dnr-v2 and +AudioCaps, respectively, to indicate which dataset each model was trained on.
Since TDANet and CodecFormer were originally developed for speech separation, we re-trained newly initialized versions on the dnr-v2 training set using the same configuration as CodecSep. For AudioSep, we evaluate the publicly available pre-trained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps using our setup for consistency. We also include SDCodec, using the official pre-trained models made available by the authors. To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing during training and inference. This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
Table 1: Results: Separation Performance, Universal Source Separation (dnr-v2-test)
Model Metric ($\uparrow$) Separation
Music Speech Sfx
AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
CodecSep + dnr-v2 SI-SDR $6.72^{\pm 2.77}$ $1.95^{\pm 2.84}$ $6.75^{\pm 3.83}$
(ablate Masker) ViSQOL $2.48^{\pm 0.58}$ $2.58^{\pm 0.50}$ $2.08^{\pm 0.74}$
AudioSep + AudioCaps SI-SDR $14.86^{\pm 23.08}$ $7.11^{\pm 25.80}$ $14.63^{\pm 23.26}$
(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
CodecSep + AudioCaps SI-SDR $8.46^{\pm 2.78}$ $2.47^{\pm 2.91}$ $5.93^{\pm 4.33}$
(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
4.4 Results and Discussions
Tables 16 present universal source separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols. Table 1 reports dnr-v2 test results for speech, music, and SFX: text-guided models use generic prompts for speech/music and ground-truth compositional captions for SFX; we also include a masker ablation to isolate the role of the Transformer masker. Table 2 studies SFX prompt granularity across three regimes—(i) fixed-stem, non-text baselines (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts ({music, speech, sfx}), and (iii) a universal setup with fine-grained, compositional SFX prompts—thereby aligning input conditions for fair comparison with fixed-head systems. Table 3 isolates architecture: decoder-style generation (CodecFormer) vs. an unguided 3-stem masker variant and its text-guided counterpart, all operating in the codec latent space. To assess out-of-domain generalization, Table 4 evaluates on AudioCaps-test with mixtures of three randomly sampled sources and prompts drawn from captions (not tied to fixed labels). Beyond AudioCaps/dnr-v2, Table 5 benchmarks on four additional open-domain corpora (ESC-50, Clotho-v2, AudioSet-eval, VGGSound) under the same prompt protocol. Table 6 probes prompt paraphrasing robustness by replacing the generic speech/music prompts with unseen synonymic variants at inference (zero-shot paraphrase test). Finally, Table 7 compares end-to-end and architecture-only GMACs for spectrogram-domain separation versus codec-latent masking, including the practical code-stream case. All models are evaluated against the original (uncompressed) ground truth; our methods are highlighted in bold; and we report mean and standard deviation ($1\sigma$) over each evaluation set.
Source Separation Performance on dnr-v2 (Table 1). CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot scenario) and retrained AudioSep+dnr-v2 across all categories, with significant SI-SDR gains in speech ($\mathbf{9.97}$ vs. 4.924.92/ $7.68$ dB), music ($\mathbf{1.15}$ vs. -2.462.46/ -$5.55$ dB), and SFX ($\mathbf{0.89}$ vs. -0.340.34/ -$4.66$ dB). Regarding ViSQOL, CodecSep matches or exceeds AudioSep in speech and music but slightly underperforms in SFX. While AudioSep is trained on a diverse dataset with sfx prompts covering isolated events and complex mixtures, CodecSep is trained exclusively on long-form, compositional SFX prompts from dnr-v2. Despite this narrower prompt exposure, CodecSep remains robust and achieves the highest overall SI-SDR. As part of our ablation study, we evaluate a lightweight variant—CodecSep+dnr-v2 (ablate Masker)—that removes the transformer masker and applies text conditioning directly to the encoder. While it performs comparable to AudioSep+dnr-v2 in SI-SDR, it achieves better perceptual quality for speech. The drop in separation quality stems from FiLM distorting the encoder’s latent space, which holds the mixture representation.
Effect of SFX Prompt Granularity during training (Table 2). We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts (“music/speech/sfx”), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions. Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality, despite those baselines being explicitly supervised per head. Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses the spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts. Moreover, replacing the single “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement. While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
Table 2: Results: Impact of SFX Prompt Granularity on Universal Source Separation (dnr-v2-test)
Model Metric ($\uparrow$) Separation
Music Speech Sfx
3-Stem: Fixed stem baselines (no text-guidance)
TDANet SI-SDR $1.84^{\pm 3.55}$ $10.18^{\pm 2.91}$ $1.36^{\pm 4.90}$
ViSQOL $2.85^{\pm 0.58}$ $3.09^{\pm 0.43}$ $2.44^{\pm 0.72}$
CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
SDCodec SI-SDR $\mathbf{1.85^{\pm 3.68}}$ $\mathbf{11.32^{\pm 2.98}}$ $\mathbf{1.77^{\pm 4.08}}$
ViSQOL $\mathbf{2.96^{\pm 0.56}}$ $\mathbf{3.49^{\pm 0.40}}$ $\mathbf{2.64^{\pm 0.73}}$
3-Stem: {“music”,“speech”,“sfx”} as generic prompt
AudioSep SI-SDR $\mathbf{-2.46^{\pm 4.06}}$ $4.92^{\pm 4.21}$ $6.65^{\pm 4.73}$
(zero-shot) ViSQOL $\mathbf{2.86^{\pm 0.63}}$ $\mathbf{3.11^{\pm 0.56}}$ $2.08^{\pm 0.68}$
AudioSep + dnr-v2 SI-SDR $6.22^{\pm 2.77}$ $\mathbf{7.71^{\pm 3.11}}$ $2.11^{\pm 3.90}$
ViSQOL $2.55^{\pm 0.57}$ $2.47^{\pm 0.37}$ $2.39^{\pm 0.74}$
CodecSep + dnr-v2 SI-SDR $7.71^{\pm 2.84}$ $4.63^{\pm 2.48}$ $\mathbf{0.58^{\pm 4.15}}$
ViSQOL $2.45^{\pm 0.55}$ $2.70^{\pm 0.49}$ $\mathbf{2.39^{\pm 0.70}}$
Universal: “music”,“speech”, fine-grained sfx prompt
AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
Why use a Transformer masker instead of a decoder?(Table 3). We compare (i) CodecFormer (decoder-style source generation), (ii) CodecSep (unguided, 3-stem) which uses the CodecFormer Transformer as a masker over codec latents, and (iii) CodecSep (text-guided). The results on dnr-v2-test exhibit two clear trends. First, replacing decoder-style generation with masking consistently strengthens separation across music, speech, and SFX. This aligns with our design rationale: in the DAC latent domain, the masker modulates existing, semantically structured content instead of synthesizing new signals, which (a) reduces artifacts and cross-talk leakage, (b) preserves long-range periodicity/timbre and transient organization already encoded by the codec, and (c) stabilizes optimization compared to end-to-end generation. Second, adding text guidance yields a further uniform improvement. The masker formulation concentrates Transformer capacity on selection (“where/how much” to pass) rather than generation (“what” to produce). Taken together, these findings support our design choice: in the codec latent space, a FiLM-conditioned Transformer masker is a more effective and robust allocator of model capacity than a decoder, with text guidance further sharpening that allocation.
Table 3: Results: Architectural advantages in using CodecFormer decoder as masker (dnr-v2-test)
Model Metric ($\uparrow$) Separation
Music Speech Sfx
CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
CodecSep + dnr-v2 SI-SDR $1.15^{\pm 3.35}$ $9.90^{\pm 2.91}$ $0.82^{\pm 4.18}$
(unguided, 3-stem) ViSQOL $2.75^{\pm 0.55}$ $3.09^{\pm 0.45}$ $\mathbf{2.48^{\pm 0.72}}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
(text-guided) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
Table 4: Results: Generalization, Universal Source Separation (AudioCaps-test)
Model Separation
SI-SDR ($\uparrow$) ViSQOL ($\uparrow$)
AudioSep $\mathbf{-2.51^{\pm 12.14}}$ $\mathbf{2.44^{\pm 1.08}}$
AudioSep + dnr-v2 (zero-shot) $6.44^{\pm 11.48}$ $2.33^{\pm 1.08}$
CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.09^{\pm 11.62}}$ $2.24^{\pm 1.16}$
AudioSep + AudioCaps $9.17^{\pm 18.71}$ $2.29^{\pm 1.11}$
CodecSep + AudioCaps $\mathbf{-6.19^{\pm 10.58}}$ $2.14^{\pm 1.00}$
Table 5: Results: Further Benchmarking on ESC-50, Clotho-v2, AudioSet-eval, VGGSound
Model Metric ($\uparrow$) Separation
ESC-50 Clotho-v2 AudioSet-eval VGGSound
AudioSep + dnr-v2 SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$
ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$
CodecSep + dnr-v2 SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$
ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$
Generalization across open-domain datasets.
Table 4 reports generalization results on the AudioCaps-test set derived from AudioSet. AudioSep benefits from distributional alignment, having been trained on diverse datasets, including AudioSet, and consequently achieves the strongest separation performance. CodecSep+dnr-v2 generalizes well to AudioCaps and outperforms AudioSep+dnr-v2 in SI-SDR while maintaining competitive ViSQOL scores. When retrained on AudioCaps, CodecSep again outperforms AudioSep in separation quality, demonstrating strong cross-domain robustness. Table 1 further supports this trend on the more challenging dnr-v2 test set, where CodecSep+AudioCaps outperforms AudioSep+AudioCaps in SI-SDR across all sources while maintaining comparable perceptual quality. However, both models experience a performance drop on dnr-v2 due to its increased mixture complexity, often containing speech, music, and three or more overlapping SFX sources—making it significantly more challenging than the simpler mixtures seen during AudioCaps training.
Further benchmarking on ESC-50, Clotho-v2, AudioSet-eval, and VGGSound (Table 5).
Extending beyond AudioCaps and dnr-v2, we evaluate both systems on four additional open-domain benchmarks spanning environmental sound classification (ESC-50), audio captioning-style corpora (Clotho-v2), weakly labeled web-scale audio (AudioSet-eval), and visually grounded audio (VGGSound). Under matched training data and prompt protocols, CodecSep+dnr-v2 consistently outperforms AudioSep+dnr-v2 across all four datasets in both separation fidelity and perceptual quality. Taken together with the dnr-v2 analyses, these results indicate that CodecSep’s latent-domain masking—leveraging the codec’s structured, semantically aligned representations—confers robust cross-domain behavior while using $25\times$ less architecture-only compute than spectrogram separators. This combination of accuracy and efficiency is particularly attractive in practical deployments, where models must generalize to diverse sound taxonomies under strict compute and memory budgets.
Robustness to prompt paraphrasing (Table 6) . To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}. This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference. We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts. Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
Table 6: Results: Impact of prompt ambiguity (dnr-v2-test)
Model Metric ($\uparrow$) Separation
Music Speech
Using Generic Prompts for Speech and Music
AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$
ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$
Using ambiguous prompts for Speech and Music
AudioSep + dnr-v2 SI-SDR $6.43^{\pm 3.29}$ $4.14^{\pm 3.77}$
ViSQOL $2.53^{\pm 0.57}$ $2.64^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $\mathbf{-5.60^{\pm 3.61}}$ $\mathbf{4.19^{\pm 4.18}}$
ViSQOL $\mathbf{2.56^{\pm 0.58}}$ $\mathbf{2.74^{\pm 0.51}}$
Table 7: Full Inference Complexity (GMACs $\downarrow$)
Model Input: Audio Stream Input: Code Stream Architecture-only
AudioSep $33.5$ $73.6$ $33.5$
CodecSep $41.45$ $\mathbf{1.35}$ ($\downarrow 54\times$) $\mathbf{1.35}$ ($\downarrow 25\times$)
Codec GMACs used above: DAC Enc(.) =12.28=12.28, Dec(.) =27.82=27.82.
Full inference complexity (Table 7) . We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep). Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep. Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space. Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting. Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth. Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
Relative–gain summaries (Appendix A).
For a compact view of percent improvements, Appendix A Table 8 reports relative gains of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text. In brief: on dnr-v2, CodecSep yields large improvements in SI-SDR with clear perceptual lifts; under paraphrased prompts, gains are smaller but remain positive, indicating robustness to lexical variation; across ESC-50, Clotho-v2, AudioSet-eval, and VGGSound, SI-SDR gains are consistent while ViSQOL deltas are modest; and when trained on AudioCaps, CodecSep maintains an advantage on AudioCaps-test and shows strong improvements when evaluated on dnr-v2.
Reconstruction study (Appendix B).
We probe source leakage with a single-source reconstruction diagnostic on dnr-v2-test : each input contains one target source and models are either prompted with the matching class/label (text-guided) or routed to the corresponding fixed head (non–text-guided). We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original. This is a leakage/consistency check—not a primary separation metric; full setup/results are in the Appendix B (Table 9).
5 Conclusion
[…]
Limitations and clarifications. (1) Codec/bandwidth. Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic. The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning. (2) Data and prompts. Training data scale and prompt diversity are modest relative to open-domain audio. As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains. (3) Temporal prompting. While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction. (4) Perceptual SFX quality. In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work. (5) Perceptual evaluation. Following CodecFormer, we report ViSQOL as a proxy for listening quality (predicting MOS-LQO on a 1–5 scale). Although we did not conduct a formal multi-listener study, we include real-world audio outputs for both CodecSep and AudioSep in the supplementary material, together with a Markdown summary of an informal author listening pass.
Future Works. We will scale CodecSep with higher-bandwidth/stereo codecs, broaden prompt coverage (including temporal and referring-expression cues), explore stronger self-alignment objectives, and investigate alternative conditioning (audio-only or multimodal) within the same FiLM-masking framework. Combined with its large compute reduction and code-stream compatibility, these extensions position CodecSep as a practical foundation for fast, universal, and deployable prompt-guided source separation.
Appendix A Discussion of Relative–Gain Summaries.
Tables 8(a)8(d) consolidate relative improvements of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text. On DnR-v2 (Table 8(a)), CodecSep delivers large SI-SDR gains—especially for music and SFX—together with a clear perceptual lift. Under paraphrased prompts (Table 8(b)), gains are smaller but remain positive, indicating robustness to lexical variation. Across additional open-domain benchmarks (Table 8(c)), SI-SDR gains are consistent while ViSQOL deltas are modest, aligning with cross-domain trends reported earlier. Finally, when trained on AudioCaps (Table 8(d)), CodecSep maintains an advantage on AudioCaps-test and yields strong improvements on DnR-v2, supporting the claim that codec-latent masking generalizes well across datasets and prompt regimes.
Table 8: Relative gains (%) of CodecSep over AudioSep under matched training/prompt settings. Each sub-table reports percent improvements for a specific evaluation setup.
(a) DnR-v2 test set
Metric Relative Gain (%)
Speech Music SFX
SI-SDR $\mathbf{+29.8}$ $\mathbf{+120.7}$ $\mathbf{+119.1}$
ViSQOL $\mathbf{+26.1}$ $\mathbf{+10.5}$ $\mathbf{+0.5}$
(b) Ambiguous prompts (speech & music paraphrases)
Metric Relative Gain (%)
Speech Music
SI-SDR $\mathbf{+1.2}$ $\mathbf{+13.0}$
ViSQOL $\mathbf{+3.8}$ $\mathbf{+1.2}$
(c) Additional open-domain benchmarks
Metric AudioCaps ESC-50 Clotho-v2 AudioSet VGGSound
SI-SDR $\mathbf{+5.5}$ $\mathbf{+24.3}$ $\mathbf{+30.0}$ $\mathbf{+16.4}$ $\mathbf{+13.0}$
ViSQOL 4.3-4.3 $\mathbf{+2.2}$ $\mathbf{+2.4}$ $\mathbf{+2.9}$ $\mathbf{+2.9}$
(d) Training on AudioCaps
Metric AudioCaps-test dnr-v2
Music Speech SFX
SI-SDR $\mathbf{+32.5}$ $\mathbf{+43.1}$ $\mathbf{+134.7}$ $\mathbf{+59.5}$
ViSQOL 6.5-6.5 3.8-3.8 $\mathbf{+5.4}$ -4.2
Appendix B Further Studies: Reconstruction Performance.
Table 9 assesses performance under a single-source reconstruction setting on the dnr-v2-test set, where each model is prompted to reproduce the input source. In addition, we report mixture reconstruction scores obtained by summing the separated sources and comparing them to the original mixture.
Among non-text-guided models, TDANet yields the best single-source reconstruction, while SDCodec performs better on mixture reconstruction. Replacing decoder-style generation in CodecFormer with a Transformer masker over codec latents in CodecSep+ dnr-v2, (unguided 3-stem) markedly improves both per-stem and mixture reconstruction fidelity and perceptual quality. Masking modulates information already organized in the codec manifold ($Z$) rather than re-synthesizing it from scratch, avoiding collapse/artifacts and yielding tighter mixture consistency.
Among the text-guided models, CodecSep achieves reconstruction performance comparable to the pre-trained and retrained AudioSep variants across all source types. AudioSep consistently excels in reconstruction due to its STFT-based masking pipeline, which enables more controlled and artifact-free waveform synthesis. However, CodecSep surpasses the pretrained AudioSep in SFX reconstruction—across both SI-SDR and ViSQOL on dnr-v2. Notably, the masker ablated CodecSep variant delivers the best reconstruction performance on dnr-v2. With isolated single-source input and matching prompts, direct conditioning minimally disturbs the NAC latent space, allowing high-fidelity reconstruction in this variant. However, strong mixture reconstruction despite poor separation suggests source leakage across separated outputs.
Table 9: Results: Reconstruction Performance, Universal Source Separation (dnr-v2-test)
Model Metric ($\uparrow$) Reconstruction
Mixture Music Speech Sfx
3-Stem: Fixed Stem, Non Text-guided
TDANet SI-SDR $3.26^{\pm 7.78}$ $\mathbf{7.96^{\pm 5.29}}$ $\mathbf{11.07^{\pm 3.32}}$ $\mathbf{4.70^{\pm 5.11}}$
ViSQOL $3.93^{\pm 0.35}$ $\mathbf{4.20^{\pm 0.46}}$ $\mathbf{4.51^{\pm 0.32}}$ $\mathbf{4.09^{\pm 0.39}}$
CodecFormer SI-SDR $47.55^{\pm 9.51}$ $47.11^{\pm 10.97}$ $47.77^{\pm 9.65}$ $48.21^{\pm 9.87}$
ViSQOL $1.02^{\pm 0.07}$ $4^{\pm 0.12} 1.01±0$$^{\pm 0.06} 1.1$$16^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $3.42^{\pm 1.85}$ $4.12^{\pm 3.97}$ $6.15^{\pm 2.87}$ $0.83^{\pm 5.16}$
(unguided, 3-stem) ViSQOL $3.23^{\pm 0.20}$ $3.02^{\pm 0.33}$ $3.47^{\pm 0.24}$ $3.23^{\pm 0.46}$
SDCodec SI-SDR $\mathbf{6.98^{\pm 2.49}}$ $7.65^{\pm 4.60}$ $8.28^{\pm 3.26}$ $2.54^{\pm 5.65}$
ViSQOL $\mathbf{4.29^{\pm 0.15}}$ $03^{\pm 0.28} 4.4$$44^{\pm 0.15}$ $3.98^{\pm 0.34}$
Text-guided
AudioSep SI-SDR $5.53^{\pm 1.96}$ $4.69^{\pm 5.36}$ $10.97^{\pm 2.99}$ $1.99^{\pm 5.68}$
(zero-shot) ViSQOL $4.06^{\pm 0.38}$ $3.75^{\pm 0.65}$ $4.57^{\pm 0.13}$ $3.19^{\pm 0.77}$
AudioSep + dnr-v2 SI-SDR $6.47^{\pm 2.26}$ $7.99^{\pm 4.55}$ $8.13^{\pm 3.35}$ $2.29^{\pm 5.95}$
ViSQOL $4.20^{\pm 0.18}$ $4.12^{\pm 0.21}$ $3.03^{\pm 0.29}$ $3.81^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $4.06^{\pm 2.06}$ $3.91^{\pm 3.93}$ $6.10^{\pm 2.86}$ $0.67^{\pm 5.29}$
ViSQOL $74^{\pm 0.22} 3.3$$37^{\pm 0.33}$ $3.83^{\pm 0.24}$ $3.54^{\pm 0.44}$
CodecSep + dnr-v2 SI-SDR $\mathbf{12.24^{\pm 2.42}}$ $\mathbf{12.58^{\pm 3.81}}$ $\mathbf{13.59^{\pm 2.59}}$ $\mathbf{8.67^{\pm 4.17}}$
(ablate Masker) ViSQOL $\mathbf{4.44^{\pm 0.14}}$ $\mathbf{4.13^{\pm 0.31}}$ $\mathbf{3.85^{\pm 0.34}}$ $3.76^{\pm 0.54}$
AudioSep + AudioCaps SI-SDR $6.67^{\pm 2.52}$ $8.10^{\pm 4.63}$ $8.39^{\pm 3.21}$ $2.42^{\pm 6.12}$
(zero-shot) ViSQOL $4.23^{\pm 0.19}$ $4.12^{\pm 0.21}$ $4.17^{\pm 0.21}$ $\mathbf{3.82^{\pm 0.46}}$
CodecSep + AudioCaps SI-SDR $0.59^{\pm 1.89}$ $0.24^{\pm 5.15}$ $11.03^{\pm 5.21}$ $1.23^{\pm 4.84}$
(zero-shot) ViSQOL $3.33^{\pm 0.23}$ $2.91^{\pm 0.64}$ $1.69^{\pm 0.48}$ $3.44^{\pm 0.41}$

### v2

@@
 1 Introduction
-In this manuscript, we propose CodecSep, a novel framework for text-conditioned source separation that combines the interpretability of prompt-driven separation with the compute efficiency of neural audio codecs (NACs).
-CodecSep is the first model to bridge NACs with universal source separation (USS), enabling flexible extraction of arbitrary sound sources based on natural language queries.
-By conditioning a transformer-based masker on text embeddings generated via the text branch of CLAP—a multimodal encoder trained on audio-text pairs Wu et al.
-(2023)—CodecSep performs state-of-the-art (SOTA) separation directly in the NAC encoder latent space, introducing semantic control while maintaining very low compute cost.
-The growing demand for flexible, real-time audio processing in bandwidth-constrained or on-device scenarios calls for generalizable and computationally efficient methods.
-Traditional separation models tackle the challenge of disentangling sound sources from complex mixtures Vincent et al.
-(2018) but are typically specialized for narrow domains such as speech or music.
-USS broadens this scope by separating arbitrary sources without reliance on predefined labels.
-Within this setting, text-guided source separation has emerged as a compelling approach, enabling users to specify the target source via natural language prompts—offering greater transparency, flexibility, and accessibility in real-world applications.
-Recent advances have embraced this direction by integrating textual prompts into separation pipelines.
-Architectures based on the Conv-TasNet encoder-masker-decoder blueprint Luo and Mesgarani (2019) have achieved strong results in domain-specific tasks Chen et al.
-(2020b); Subakan et al.
-(2021); Li et al.
+We propose CodecSep, a text-conditioned universal source separation (USS) framework that marries the interpretability of prompt-driven extraction with the efficiency of neural audio codecs (NACs).
+To our knowledge, CodecSep is the first system to bridge NACs with USS: it conditions a transformer masker on CLAP text embeddings Wu et al.
+(2023) via Feature-wise Linear Modulation (FiLM) Perez et al.
+(2018), and performs separation directly in the codec encoder latent space.
+This design introduces semantic control while preserving the low compute footprint of codec representations, enabling on-device and real-time deployment.
+Flexible, real-time separation on bandwidth- or compute-constrained platforms remains challenging.
+Classic models disentangle sources from complex mixtures Vincent et al.
+(2018) but are often domain-specific (e.g., speech/music) and heavy.
+Recent text-guided systems like AudioSep Liu et al.
+(2024) extend encoder–masker–decoder designs (e.g., Conv-TasNet-style Luo and Mesgarani (2019)) by injecting semantics from BERT/CLAP through FiLM layers Devlin et al.
+(2019); Wu et al.
+(2023); Perez et al.
 (2018).
-Systems like AudioSep Liu et al.
-(2024) extend this framework to the text-conditioned setting by encoding text queries with models such as BERT Devlin et al.
-(2019) or CLAP Wu et al.
-(2023) and injecting semantic information through Feature-wise Linear Modulation (FiLM) layers Perez et al.
+However, spectrogram/waveform-domain separators trained with SI-SDR-style losses Luo and Mesgarani (2019); Le Roux et al.
+(2019) are compute-intensive and sensitive to compression artifacts, often pushing inference to the cloud.
+NACs such as SoundStream, Encodec, and DAC Zeghidour et al.
+(2022); Défossez et al.
+(2022); Kumar et al.
+(2023) compress audio to discrete tokens with Residual Vector Quantization (RVQ), providing compact, perceptually aligned latents useful for generation and conditioned synthesis Borsos et al.
+(2023); Wang et al.
+(2023, 2024); Du et al.
 (2024).
-These modifications allow intermediate representations to adapt dynamically to prompt semantics, steering the model towards the desired source.
-However, such models, trained with signal-level objectives like scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al.
-(2019), are not only highly sensitive to compression artifacts but also tend to be large and computationally intensive.
-This limits their practicality for on-device or real-time applications where resources are constrained.
-As a result, they are often deployed via the cloud, introducing additional latency and infrastructure costs.
-Neural audio codecs (NACs) offer a promising alternative.
-Models like SoundStream Zeghidour et al.
-(2022), Encodec Défossez et al.
-(2022), and DAC Kumar et al.
-(2023) compress audio into discrete token sequences using encoder-quantizer-decoder pipelines.
-Leveraging Residual Vector Quantization (RVQ), they enable adaptive bitrate control and efficient representation.
-These codecs are trained on diverse audio data using a combination of adversarial and perceptual losses, which promotes both fidelity and robustness under compression.
-The resulting token representations have proven useful in tasks such as audio generation for multimodal models Wang et al.
-(2024); Du et al.
-(2024) and language-conditioned synthesis Borsos et al.
-(2023); Wang et al.
-(2023).
-Building on this foundation, recent efforts like CodecFormer Yip et al.
-(2024b) and SDCodec Bie et al.
-(2024) have explored incorporating NACs into source separation pipelines.
-While these approaches achieve lightweight, high-fidelity performance, they remain limited to separating fixed types of sources (e.g., speech separation).
-CodecSep addresses this gap by combining the computational efficiency of NACs with the flexibility of language-guided separation.
-It follows a NAC-based encoder-decoder architecture and integrates a transformer-based masker conditioned on CLAP-derived text embeddings.
-These embeddings generate FiLM parameters that modulate each transformer layer, enabling the model to selectively extract sources that align with the user’s query.
-Operating entirely in the codec latent space, CodecSep delivers interpretable, high-performance separation under strict computational constraints.
+Prior codec–separation hybrids (CodecFormer, SDCodec) Yip et al.
+(2024b); Bie et al.
+(2024) demonstrate light-weight, high-fidelity separation but target fixed stems (e.g., speech separation or speech vs.
+music vs.
+SFX) rather than open-domain prompt-conditioned USS.
+CodecSep adopts a frozen DAC encoder–decoder backbone and inserts a FiLM-conditioned transformer masker that predicts a soft mask over codec latents.
+CLAP-derived text embeddings Wu et al.
+(2023) are mapped to per-layer FiLM parameters, modulating the masker’s intermediate activations to align the selected latent subspace with the query semantics.
+Operating on compact codec features cuts memory traffic and MACs compared to spectrogram-domain pipelines, while preserving the codec’s inductive biases (periodicity, timbre, transients).
+In doing so, CodecSep delivers interpretable, prompt-guided separation with markedly lower compute, without sacrificing separation fidelity.
 We evaluate CodecSep for universal source separation under a comprehensive set of protocols: (i) dnr-v2 for in-domain text-guided separation; (ii) five open-domain corpora—AudioCaps, ESC-50, Clotho-v2, AudioSet-eval, and VGGSound—for cross-domain generalization; (iii) three prompt granularities (fixed-stem, generic three-stem {“speech”, “music”, “sfx”}, and a universal setting with fine-grained, compositional SFX descriptions); and (iv) paraphrase robustness via unseen synonyms for speech/music.
@@
 2 Related Work
-2.1 Audio Source Separation
-Notable approaches like DPTNet Chen et al.
+Classical audio source separation systems frequently adopt an encoder–masker–decoder design in which an encoder produces STFT-like latents, a masker predicts source-specific masks, and a decoder reconstructs waveforms.
+Representative models include DPTNet Chen et al.
 (2020b), SepFormer Subakan et al.
 (2021), and TDANet Li et al.
-(2023) follow the encoder-masker-decoder framework.
-In this setup, the encoder generates STFT-like latent representations.
-At the same time, a masker network predicts source masks, which are then applied elementwise to the encoder outputs to obtain source-specific latent representations.
-These estimates are further processed by a decoder, which mirrors the encoder architecture, to reconstruct the source audio.
-TDANet Li et al.
-(2023), the latest iteration, enhances this framework with a top-down attention mechanism, using global and local attention modules to hierarchically model multi-scale acoustic features.
-CodecSep also adopts this encoder-masker-decoder structure, extending it to operate in the NAC latent space with text guidance.
-Beyond the encoder-masker-decoder paradigm, a variety of alternative architectures have been explored.
-Models like Wave-UNet Stoller et al.
-(2018) and Demucs Défossez et al.
-(2019, 2021) utilize encoder-decoder designs that estimate source waveforms directly, omitting the masking stage altogether.
-Other approaches operate in the frequency domain, notably, models like MM-DenseLSTM Takahashi et al.
+(2023), the last introducing a top-down attention scheme that blends global and local attention to capture multi-scale acoustic structure.
+Beyond masking pipelines, several works generate waveforms directly in the time domain (Wave-UNet Stoller et al.
+(2018), Demucs Défossez et al.
+(2019, 2021)) or operate fully in the complex STFT domain with joint magnitude–phase modeling (MM-DenseLSTM Takahashi et al.
 (2018), DCCRN Hu et al.
-(2020), and UNet-based architectures such as Spleeter Hennequin et al.
-(2020) perform separation directly in the complex STFT domain, jointly modelling magnitude and phase components.
-These diverse strategies highlight the broad design space of source separation, spanning both time and frequency domains, with and without masking.
-2.2 Universal Source Separation
-Prior work on source separation has concentrated mainly on speech or music-specific tasks.
-More recent advances have shifted toward USS, which aims to disentangle arbitrary sources from a mixture without assuming predefined source categories.
-Supervised USS models often employ Permutation Invariant Training (PIT) Yu et al.
+(2020), Spleeter Hennequin et al.
+(2020)), underscoring the breadth of design choices across time and frequency domains.
+Moving from domain-specific separation to universal source separation (USS), supervised systems often rely on Permutation Invariant Training (PIT) Yu et al.
 (2017); Kavalerov et al.
-(2019), while unsupervised approaches like Mixture Invariant Training (MixIT) Wisdom et al.
-(2020) learn directly from mixtures without access to clean references.
-However, both approaches assume a fixed maximum number of sources in the mixture and produce all outputs indiscriminately.
-As a result, an additional source identification step is needed post-separation to determine which of the estimated signals corresponds to the target source.
-Query-Guided Source Separation (QSS) extends USS by enabling source extraction through external queries, including visual input, exemplar audio, class labels, or text.
-Text queries are particularly effective, offering a compact and expressive way to convey high-level semantics and encode spatial or temporal cues of the query source—all without the need for expensive reference data.
-Recent work such as AudioSep Liu et al.
-(2024) explores text-guided separation by integrating embeddings from models like BERT Devlin et al.
-(2019) and CLAP Wu et al.
-(2023), the latter of which aligns audio and text in a shared embedding space.
-AudioSep employs FiLM conditioning into the separation pipeline at intermediate layers of the masker.
-CodecSep adopts a similar strategy but operates in the NAC encoder latent space instead.
-2.3 Source Separation with Neural Audio Codecs
-Recent work has explored integrating neural audio codecs into source separation pipelines.
+(2019), while unsupervised approaches such as MixIT Wisdom et al.
+(2020) learn from mixtures alone.
+However, both typically assume a fixed maximum number of sources and emit all estimates indiscriminately, necessitating a separate identification step.
+Query-Guided Source Separation (QSS) addresses this limitation by conditioning extraction on external queries—visual, exemplar audio, class labels, or text.
+Text queries are compact and expressive, capture high-level semantics, and avoid costly references.
+AudioSep Liu et al.
+(2024) exemplifies this direction by injecting BERT Devlin et al.
+(2019) or CLAP Wu et al.
+(2023) embeddings via FiLM Perez et al.
+(2018) at intermediate masker layers to steer separation toward the requested source.
+Neural audio codecs (NACs) have recently been integrated into separation pipelines to improve efficiency.
 CodecFormer Yip et al.
-(2024b) and its extension CodecFormer-EL Yip et al.
-(2024a) both leverage DAC as the encoder-decoder backbone, performing separation directly in the codec’s latent embedding space via a transformer-based separation module.
-The original CodecFormer is trained using a negative SI-SDR loss to supervise the predicted outputs.
-At the same time, CodecFormer-EL introduces an embedding-level objective that minimizes the distance between the separator’s output and the DAC encoder’s ground truth latent embeddings for each source.
-In contrast, SD-Codec Bie et al.
-(2024) integrates source separation more tightly into the codec itself.
-It is explicitly designed to disentangle mixtures into speech, music, and SFX.
-To achieve this, it introduces three dedicated RVQ branches—each trained to encode a specific source category.
-During encoding, the outputs of these domain-specific quantizers are summed across time to form a mixture-level representation, which is then decoded to reconstruct both the whole mixture and individual sources.
-This design enables SD-Codec to embed separation behavior directly into the codec’s structure.
-In this work, we repurpose the CodecFormer transformer separator into a text-conditioned mask predictor, enabling prompt-guided control over the separation process.
+(2024b) performs separation directly in DAC Kumar et al.
+(2023) latent space with a transformer separator trained using negative SI-SDR, and CodecFormer-EL Yip et al.
+(2024a) further adds an embedding-level objective that aligns separator outputs with DAC encoder latents.
+SD-Codec Bie et al.
+(2024) embeds separation behavior inside the codec by assigning dedicated RVQ branches to speech, music, and SFX, summing their codes to form mixture-level representations while enabling per-stem decoding.
+In contrast, our work repurposes the CodecFormer transformer into a text-conditioned mask predictor operating in the NAC latent domain: CLAP-driven FiLM conditioning modulates intermediate transformer activations to estimate a single, prompt-guided mask, enabling efficient, controllable USS without fixed source heads.
 Figure 1: An overview of CodecSep. (Left) The full pipeline for text-guided USS. (Right) The integration of text conditioning into intermediate layers of transformer masker via FiLM layers.
 3 Method
-CodecSep is based on the 16kHz variant of CodecFormer Yip et al.
-(2024b), which we adapt to enable text-driven source separation.
-Specifically, we modify the transformer module to function as a source mask estimator and incorporate Feature-wise Linear Modulation (FiLM) Perez et al.
-(2018) to inject text-based conditioning into the separation process.
-FiLM layers are applied to the transformer’s intermediate activations, allowing the model to adjust its internal representations based on the semantics of the input query.
-To produce the FiLM parameters, a lightweight query network consumes embeddings produced by a pretrained CLAP encoder Wu et al.
-(2023) from the input prompt.
-Figure 1 (Left) illustrates our text-guided USS pipeline.
-Unlike AudioSep’s raw STFT-based, compute-heavy architecture, CodecSep operates on compact codec representations, offering significantly greater compute efficiency for resource-constrained edge deployment.
-Figure 1 (Right) illustrates how text-conditioning is integrated into the intermediate encoder layers of the transformer masker.
-3.1 Model Architecture
-CodecSep utilizes DAC Kumar et al.
-(2023) for the encoder-decoder modules and a transformer-based masker network with text conditioning to disentangle query sources within the DAC encoder representation space.
-We will discuss each component in detail and present how we incorporated text conditioning within the transformer masker architecture.
-3.1.1 Codec Backbone
-DAC Kumar et al.
-(2023) builds upon the Encodec Défossez et al.
-(2022) and SoundStream Zeghidour et al.
-(2022) architectures and utilizes fully convolutional encoder and decoder networks.
-Leaky ReLU activations in the architectures are substituted with the periodic Snake Activation function Ziyin et al.
-(2020) expressed as $x+\sin^{2}{x}$.
-This enhances the modeling of periodic audio signals by incorporating a periodic inductive bias.
-Residual vector quantization is integrated into the encoder output, which enables audio compression at varying bit rates.
-In addition, factorized codes—where code assignment occurs in a lower-dimensional space than the codebook vectors—and L2-normalized codebook vectors are utilized to enhance codebook efficiency.
-Let us consider compressing an audio recording $x(t)$ of duration $1$s, sampled at $F_{s}=24$kHz to target bit-rate of R=6000R=6000bps.
-The encoder network $Enc(\cdot)$ down-samples $x(t)$ by a factor of $M=320$ and outputs $T=\frac{F_{s}}{M}=75$ frames of latent representations, denoted as $Z=[z_{t}\in R^{d}|t\in[T]]$.
-Each frame $z_{t}$ is allocated $r=\frac{R}{T}=80$ bits.
-Residual vector quantization (RVQ) efficiently allocates the bit-rate budget across a cascade of $N_{q}=8$ vector quantizers.
-The first vector quantizer in the cascade quantizes the frame $z_{t}$, and the residuals (the quantization error) are computed.
-The residuals are then iteratively quantized by the remaining $N_{q}-1$ quantizers in the cascade.
-Each quantizer is assigned $r_{i}=\frac{r}{N_{q}}=10$ bits, maintaining a manageable code-book size of $2^{10}=1024$.
-The number of quantizers in the cascade $N_{q}$ is key in balancing computational complexity and compression for a given bit-rate budget $r$.
-The RVQ network $Quant(\cdot)$ produces a code matrix $A=[a_{t}\in[1024]^{8}|t\in[T]]$ given encoder output $Z$.
-Each frame $z_{t}\in Z$ is quantized into a code vector $a_{t}\in[1024]^{8}$, which can be mapped to an embedding via code-book lookup as $e_{t}=\sum_{i=1}^{8}e^{i}_{t}$, where $e^{i}_{t}$ corresponds to the embedding associated with code $a^{i}_{t}\in[1024]$ from the $i^{th}$ RVQ code-book.
-The RVQ output embedding sequence $E=[e_{t}|t\in[T]]$ is then processed by the decoder network $Dec(.)$, which up-samples it to generate $y(t)$, a reconstruction of the original audio $x(t)$.
-3.1.2 Transformer Masker
-CodecSep masker module $Mask(.)$ predicts source mask given DAC Kumar et al.
-(2023) encoder $Enc(.)$ representation $Z$.
-These masks are applied element-wise to representation $Z$ to extract the corresponding source representation estimate.
-It follows CodecFormer Yip et al.
-(2024b) Transformer separator module architecture and is designed as a stack of $L=16$ transformer layers $Mask^{l\in[L]}(.)$ of embedding size $d=256$.
-The output of the final transformer layer, $H^{L}=Mask^{L}(H^{L-1})$, is processed through $K$ distinct 1D convolutional layers to generate source masks $\{M_{k}|k\in[K]\}$.
-To preserve the periodic inductive bias across the entire model, the Snake activation function, employed in DAC Kumar et al.
-(2023), replaces tanh and LeakyReLU activations throughout the masker module.
-The resulting source representation estimates are subsequently mapped to time domain $\{\tilde{y}_{k}=Dec(M_{k}\odot Z)|k\in[K]\}$ using DAC Kumar et al.
-(2023) decoder $Dec(.)$ directly, without being processed by RVQ network $Quant(.)$.
-3.1.3 Text-guided Source Separation
-We consider a single-channel audio mixture $x(t)\in\mathbb{R}^{1\times N}$, which consists of an unknown number of overlapping open-domain sound sources. The mixture is expressed as the following sum:
-$x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$ (1)
-where $\mathcal{S}$ is the (potentially unbounded) set of sound source classes or instances present in the mixture.
-Our objective is to recover a target source waveform or stem $y_{s}(t)$ from the mixture $x(t)$, using a natural language query $\tau_{s}$ that semantically describes the desired source.
-These prompts may reference specific sources (e.g., “dog barking,” “violin solo,” “crowd cheering”) or combinations of sources (e.g., “speech and music,” “footsteps and glass breaking”), allowing for fine-grained control.
-To integrate text-based conditioning into CodecSep, we leverage CLAP Wu et al.
-(2023) and utilize a masker module $Mask(.)$ with $K=1$ final convolution to generate a single source mask $M_{s}$.
-Given a query text $\tau$ that describes the target source, CLAP Wu et al.
-(2023) generates a corresponding text embedding $e_{\tau}\in\mathbb{R}^{d}$.
-This embedding is used to modulate the intermediate $l\in[2,...,L-1]$ transformer layers $Mask^{l}(.)$ via feature-wise linear modulation (FiLM) Perez et al.
-(2018), enabling the network to estimate the desired source conditioned on the text input.
-At transformer layer $l$, the layer output $H^{l}\in\mathbb{R}^{d\times T}$ is modulated along the feature dimension using parameters $\gamma^{l},\beta^{l}\in\mathbb{R}^{d}$.
-These modulation parameters are generated by a query network $query(.)$, which processes text embedding $e_{\tau}$ to produce $(\gamma^{l},\beta^{l})_{l=1}^{L}=query(e_{\tau})$.
-The modulation transformation is then carried out as follows:
-$\tilde{H}^{l}_{s}=FiLM(H^{l}_{i},\gamma^{l}_{i},\beta^{l}_{i})=\gamma^{l}_{i}H^{l}_{i}+\beta^{l}_{i},\forall i\in[d]$ (2)
-The decoder network $Dec(.)$ subsequently processes the source representation estimate $\tilde{Z}_{s}=M_{s}\odot Z$ and generates an estimate of the query source audio $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$.
-3.2 Training Objective
-We train CodecSep to reconstruct each individual source from mixtures using textual prompts corresponding to broad categories such as music, speech, and a wide range of sound effects (SFX).
-We also reconstruct the input mixture $x$ by decoding the summed source latent estimates: $\tilde{x}=g(\sum_{s}\tilde{Z}_{s})$.
-The model is optimized using the negative scale-invariant signal-to-distortion ratio (SI-SDR) loss Luo and Mesgarani (2019); Le Roux et al.
-(2019), encouraging the faithful separation of sources aligned with the given queries.
-To support generalization across varied SFX types, prompts associated with SFX are not limited to atomic events—they may describe composite scenarios involving multiple overlapping sources (e.g., “footsteps, background traffic,” “crowd cheering, fireworks”), requiring the model to isolate multi-faceted acoustic scenes based on complex queries.
-SI-SDR is designed to measure the similarity between the predicted $\tilde{v}$ and ground truth $v$ signals, focusing on the signal-to-noise ratio (SNR) while being invariant to scale differences.
-It is given as,
+CodecSep adapts the 16 kHz DAC Kumar et al.
+(2023) codec backbone for text-driven universal source separation (USS).
+We use a transformer masker that estimates soft masks over codec latents and inject text conditioning via Feature-wise Linear Modulation (FiLM) Perez et al.
+(2018) using CLAP text embeddings Wu et al.
+(2023).
+FiLM is applied to intermediate transformer activations so the query semantics steer separation.
+Figure 1 (Left) shows the overall text-guided pipeline; Figure 1 (Right) highlights the FiLM-conditioned masker.
+Compared to STFT-domain AudioSep, operating in compact codec latents yields markedly lower compute and is amenable to edge deployment.
+3.1 Our Model
+Backbone codec.
+We use DAC Kumar et al.
+(2023) as encoder–decoder.
+Following Encodec/SoundStream, DAC uses fully-convolutional encoder/decoder with the periodic Snake activation $x+\sin^{2}x$ (replacing LeakyReLU) to bias periodic audio modeling.
+Residual vector quantization (RVQ) compresses encoder outputs with factorized codes and $\ell_{2}$-normalized codebooks.
+For a 1 s audio $x(t)$ at $F_{s}{=}24$ kHz compressed to R=6000R{=}6000 bps, the encoder $Enc(\cdot)$ downsamples by M=320M{=}320 to $T{=}{F_{s}}/{M}{=}75$ frames $Z{=}[z_{t}\!\in\!\mathbb{R}^{d}]_{t=1}^{T}$ with r=R/T=80r{=}R/T{=}80 bits/frame.
+RVQ allocates $r_{i}{=}r/N_{q}{=}10$ bits across $N_{q}{=}8$ codebooks (size $2^{10}{=}1024$).
+Given $Z$, $Quant(\cdot)$ yields discrete codes $A{=}[a_{t}\!\in\![1024]^{8}]$, which map to embeddings $e_{t}{=}\sum_{i=1}^{8}e^{i}_{t}$; $Dec(\cdot)$ upsamples $E{=}[e_{t}]$ back to waveform $y(t)$.
+Transformer masker and text-guided separation.
+We consider a mono mixture as $x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$, where $x(t)$ is the observed waveform, $\mathcal{S}$ is the (unbounded) set of source classes/instances present, and $y_{s}(t)$ is the waveform of source $s$.
+Passing $x(t)$ through the frozen DAC encoder yields codec latents $Z\in\mathbb{R}^{d\times T}$ ($d$: channel width, $T$: latent frames).
+The masker $Mask(\cdot)$ operates in this latent space to predict an element-wise mask that selects the target source.
+We adopt a CodecFormer-style transformer with L=16L{=}16 layers, width d=256d{=}256, and Snake activations.
+Given a natural-language query $\tau$ (e.g., “dog barking”, “speech and music”), we compute a CLAP text embedding $e_{\tau}\in\mathbb{R}^{d}$.
+A lightweight query network $query(.)$ maps $e_{\tau}$ to per-layer FiLM parameters $(\gamma^{l},\beta^{l})\in\mathbb{R}^{d}$ for $l\in\{2,\ldots,L{-}1\}$, applied channel-wise to intermediate activations $H^{l}\in\mathbb{R}^{d\times T}$: $\tilde{H}^{l}=\mathrm{FiLM}(H^{l};\gamma^{l},\beta^{l})=\gamma^{l}\odot H^{l}+\beta^{l}$.
+The final transformer output $H^{L}$ is then passed through a single 1D convolutional head to produce the prompt-conditioned mask $M_{s}\in[0,1]^{d\times T}$.
+We obtain source latents $\tilde{Z}_{s}=M_{s}\odot Z$ and decode with the frozen codec decoder to get the waveform estimate $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$, bypassing RVQ lookup.
+Training Objective
 […]
-The overall loss function is then given as:
+3.2 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
+Spectrogram-domain (AudioSep):
+$x\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
+where $g(\cdot,e_{\tau})$ is a FiLM-conditioned separator that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$ from $X$ and text embedding $e_{\tau}$.
+CodecSep (NAC latent-domain):
+$x\xrightarrow{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
+with frozen DAC backbone $Enc(\cdot),Dec(\cdot)$ and a FiLM-conditioned transformer masker $Mask(\cdot,e_{\tau})$ that estimates $M_{s}\!\in\![0,1]^{d\times T}$.
+Why NAC latents vs. spectrograms.
+Operating on NAC latents $Z$ slashes dimensionality while preserving perceptual factors.
+For $1\,\mathrm{s}$ audio at $32\,\mathrm{kHz}$, complex STFT with N=1024N{=}1024 and hop size M=320M{=}320 samples has $T_{\mathrm{spec}}\!\approx\!100$ frames and $F=2\times 1024$ (Re+Im) scalars per frame, so $F\cdot T_{spec}\!\approx\!204{,}800$.
+A $16\,\mathrm{kHz}$ DAC with width d=64d{=}64 and the same $M$ yields $T\!\approx\!50$ and $d\cdot T{=}64\times 50{=}3{,}200$ ($\sim 64{\times}$ smaller), shrinking $Q/K/V$ and MLP sizes and easing self-attention.
+Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs.
+Crucially, $Enc(\cdot)$ organizes $Z$ on a discriminative, perceptually aligned manifold , making selection (masking) easier than representation learning from raw $X$.
+In spectrogram systems, the separator must first learn a high-level latent from $X$ (via CNN/UNet) and then separate, coupling abstraction and masking and inflating parameters/compute.
+Masking over codec latents (leveraging the codec prior).
+Because the DAC codec induces a strong semantic prior in its latent space via residual vector quantization (RVQ) and perceptual/adversarial training, we mask the codec latents rather than generate sources from scratch as in CodecFormer.
+RVQ creates a coarse-to-fine hierarchy in $Z{=}Enc(x)\in\mathbb{R}^{d\times T}$: early stages capture coarse structure (e.g., low-frequency content, timbre), while later stages refine residual detail (e.g., high-frequency components, transients) Wang et al.
+(2023).
+We exploit this organization with a FiLM-conditioned transformer masker that predicts a soft mask $M_{s}\in[0,1]^{d\times T}$ and applies it element-wise, yielding source latent estimate $\tilde{Z}_{s}=M_{s}\odot Z$.
+In contrast to learning a generator $Sep(Z,e_{\tau}):Z\!\to\!\tilde{Z}_{s}$ as in CodecFormer, learning a mask $Mask(Z,e_{\tau}):Z\!\to\!M_{s}$ on the compact, semantically organized codec manifold both exploits the codec prior more effectively and yields a more stable optimization that converges faster.
+Moreover, masking in the denoised, low-dimensional codec space is fundamentally easier than masking in the high-dimensional, noisy spectrogram domain.
+This selection-centric design (i) constrains learning to modulation of existing latent content, (ii) avoids hallucination and reduces leakage because no new signals are synthesized, and (iii) preserves long-horizon structure (periodicity, timbre, transients) already organized by the codec, yielding stable, low-artifact separations.
+Why FiLM inside the masker.
+Placing FiLM in the masker (not in $Enc/Dec$) targets the selection step, preserves the codec manifold, and adds negligible overhead (two vectors per layer).
+Conditioning is non-iterative (single forward pass), maintaining low latency for edge/server workflows.
+Continuous latents $Z$ vs. discrete codes $A$ (and deployment path).
+For training and analysis, we operate on continuous latents $Z{=}Enc(x)\in\mathbb{R}^{d\times T}$: (i) gradients flow cleanly through $Mask(\cdot,e_{\tau})$ and $Dec(\cdot)$ with a frozen codec (no straight‐through estimators), yielding stable convergence; (ii) RVQ pretraining regularizes $Z$ so pitch, timbre, onsets/transients, and textures are hierarchically organized, providing a richer, more disentangled signal for FiLM; and (iii) $Z$ avoids run‐to‐run variance from codebook utilization (e.g., late RVQ sensitivity, bitrate truncation), reducing the need for special regularizers.
+For deployments with compressed bitstreams, we reconstruct embeddings by codebook lookup and use the same masker:
+$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big(a_{t}^{(i)}\big),$ (1)
+$E$ $=[e_{t}]_{t=1}^{T}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (2)
+When a codes-out interface is desired, we re-quantize masked embeddings and optionally decode:
+$\hat{A}_{s}=Quant(\tilde{E}_{s}),\qquad\hat{E}_{s}=\mathrm{lookup}(\hat{A}_{s}),\qquad\tilde{y}_{s}(t)=Dec(\hat{E}_{s}).$
+By design, $E\!\approx\!Z$ at the operating bitrate and $Dec(E)$ already yields high-fidelity reconstructions; because our separator is a masker (selective modulation) rather than a generator, swapping $Z\!\to\!E$ preserves the semantics needed for separation with no architectural change.
+While we report results on $Z$ to isolate separator performance and maintain stable optimization, we also evaluate the bitstream path by feeding reconstructed embeddings $E$ (codes-in) to the same trained masker without any fine-tuning; performance remains competitive relative to the $Z$ path.
+The residual gap can be narrowed with light fine-tuning the masker on $E$ or optimizing an embedding-consistency loss (cf.
+CodecFormer-EL) in place of, or alongside, SI-SDR: $\mathcal{L}_{\mathrm{emb}}=\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$.
+In deployment, the variant simply replaces the masker input with $E$ and optionally re-quantizes for codes-out.
+Deployment advantages over spectrogram-domain separators.
+In realistic pipelines, edge devices already run a codec and send code streams.
+Spectrogram models must decode $\to$ separate on $X$ $\to$ re-encode, adding latency and energy; CodecSep masks in the codec domain and can return separated code streams directly:
+$\text{Edge: }Enc(\cdot)\Rightarrow\text{Code}\Rightarrow\text{Server: reconstruct }E\!\approx\!Z\ \&\ Mask(\cdot,e_{\tau})\Rightarrow\text{Code}\Rightarrow\text{Edge: }Dec(\cdot).$
+Concretely, with codec costs $C_{\mathrm{enc}},C_{\mathrm{dec}}$, spectrogram separator cost $C_{\mathrm{spec}}$, and CodecSep masker cost $C_{\mathrm{mask}}$:
+$\text{Code-stream input: }\ \text{AudioSep}=C_{\mathrm{dec}}{+}C_{\mathrm{spec}}{+}C_{\mathrm{enc}},\quad\text{CodecSep}=C_{\mathrm{mask}}.$
+Within the separator, working on $Z/E$ (small $d,T$ with $|Z|\!\ll\!|X|$) shrinks attention/MLP activations, enabling tighter batching and lower memory bandwidth.
+Interface compatibility is immediate when only codes $A$ are available: perform lookup to $E$, apply the same FiLM-conditioned masker, and optionally re-quantize to emit $\hat{A}_{s}$.
+Overall, CodecSep (i) eliminates redundant decode/re-encode in server workflows; (ii) preserves the codec manifold (periodicity, timbre, transients) by modulating $Z/E$ instead of rebuilding $X$; and (iii) keeps conditioning non-iterative (single forward pass), yielding low-latency, high-fidelity stems at scale.
+See Appendix A for a full rationale covering latent-domain masking vs.
+spectrogram-domain generation, the RVQ-induced semantic prior, FiLM placement inside the masker, continuous $Z$ vs.
+discrete $A{\to}E$ operation (and codes-in/out), and complexity/deployment comparisons.
+4 Experiments
+Datasets.
+We evaluate across a controlled multi-stem corpus and multiple open-domain benchmarks.
+For in-domain experiments, we adapt Divide and Remaster v2 (dnr-v2) Petermann et al.
+(2022) from fixed-label three-stem separation to universal, prompt-driven separation by replacing source labels with natural-language queries.
+Speech and music are queried with broad category prompts (e.g., “speech”, “music”), while SFX stems are queried using long-form, compositional prompts (>1>1 overlapping sources) synthesized from FSD50K’s hierarchical annotations Fonseca et al.
+(2022), combining fine-grained classes with parent categories (e.g., “dog barking, Animal, engine rumbling, motor vehicle”).
+To assess cross-domain generalization, we form three-source mixtures on AudioCaps Kim et al.
+(2019) (used for both training and testing in our open-domain setting) and construct test-only three-source mixtures from ESC-50 Piczak , Clotho-v2 Drossos et al.
+(2020), AudioSet-eval Gemmeke et al.
+(2017), and VGGSound Chen et al.
+(2020a).
+Dataset construction details, clip durations, split statistics, and segmentation rules are provided in the Appendix B.
+Evaluation.
+We compare CodecSep against representative spectrogram- and codec-domain baselines—TDANet Li et al.
+(2023); Pons et al.
+(2024), CodecFormer Yip et al.
+(2024b), SDCodec Bie et al.
+(2024), and the text-guided spectrogram model AudioSep Liu et al.
+(2024).
+We report objective signal fidelity via scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al.
+(2019) and perceptual quality via ViSQOL Chinen et al.
+(2020), which measures spectro–temporal similarity between the estimate $\tilde{x}$ and reference $x$ and maps it to a 1–5 MOS-LQO scale.
+Following prior work (e.g., SDCodec Bie et al.
+(2024)), we use ViSQOL as a proxy for perceptual listening quality and complement it with a human MOS-LQS study comparing real-world outputs from CodecSep (trained on dnr-v2) and the publicly released AudioSep.
+To quantify efficiency in a hardware-agnostic manner, we report multiply–accumulate operations (MACs) using torchprofile [^1] https://github.com/zhijian-liu/torchprofile under matched input durations (2 s) and batching (batch size 2).
+Training.
+Unless otherwise stated, the DAC codec Kumar et al.
+(2023) and CLAP text encoder Wu et al.
+(2023) remain frozen; we train the FiLM-conditioned transformer masker and the lightweight query network end-to-end with an Adam optimizer Kingma and Ba (2017) and a plateau-based learning-rate schedule Mukherjee et al.
+(2019).
+We produce two variants of CodecSep, trained separately on dnr-v2 and on AudioCaps, to study distributional effects; we denote them with the suffixes “+dnr-v2” and “+AudioCaps”.
+For fair comparison, TDANet and CodecFormer are re-trained from scratch on dnr-v2 using our setup, AudioSep is evaluated both as a public checkpoint and when re-trained under matched protocols, and SDCodec is used as released by the authors.
+To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through the codec during both training and inference.
+Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix C.
+Table 1: Results: Separation Performance, Universal Source Separation (dnr-v2-test)
+Model Metric ($\uparrow$) Music Speech Sfx
+AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
+(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
+AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
+ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
+ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
+CodecSep + dnr-v2 SI-SDR $0.16^{\pm 3.55}$ $8.26^{\pm 2.60}$ $0.96^{\pm 4.20}$
+(codes in : codes out, zero-shot) ViSQOL $2.52^{\pm 0.52}$ $2.97^{\pm 0.44}$ $2.33^{\pm 0.67}$
+CodecSep + dnr-v2 SI-SDR $6.72^{\pm 2.77}$ $1.95^{\pm 2.84}$ $6.75^{\pm 3.83}$
+(ablate Masker) ViSQOL $2.48^{\pm 0.58}$ $2.58^{\pm 0.50}$ $2.08^{\pm 0.74}$
+4.1 Results and Discussions
+Tables 15 present universal source separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols.
+Table 1 reports dnr-v2 test results for speech, music, and SFX: text-guided models use generic prompts for speech/music and ground-truth compositional captions for SFX; we also include a masker ablation to isolate the role of the Transformer masker.
+Table 2 studies SFX prompt granularity across three regimes—(i) fixed-stem, non-text baselines (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts ({music, speech, sfx}), and (iii) a universal setup with fine-grained, compositional SFX prompts—thereby aligning input conditions for fair comparison with fixed-head systems.
+Table 3 isolates architecture: decoder-style generation (CodecFormer) vs.
+an unguided 3-stem masker variant and its text-guided counterpart, all operating in the codec latent space.
+To assess out-of-domain generalization, Table 4 benchmarks on five additional open-domain corpora (ESC-50, Clotho-v2, AudioSet, VGGSound, AudioCaps) with mixtures of three randomly sampled sources and prompts drawn from captions (not tied to fixed labels)..
+Table 5 probes prompt paraphrasing robustness by replacing the generic speech/music prompts with unseen synonymic variants at inference (zero-shot paraphrase test).
+Finally, Table 6 compares end-to-end and architecture-only GMACs for spectrogram-domain separation versus codec-latent masking, including the practical code-stream case.
+All models are evaluated against the original (uncompressed) ground truth; our methods are highlighted in bold; and we report mean and standard deviation ($1\sigma$).
+Source Separation Performance on dnr-v2 (cf.
+Table 1).
+CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot) and retrained AudioSep+dnr-v2 across all categories, with sizable SI-SDR gains in speech (9.97 vs.
+4.92/7.68 dB), music (1.15 vs.
+2.46-2.46/5.55-5.55 dB), and SFX (0.89 vs.
+0.34-0.34/4.66-4.66 dB).
+In ViSQOL, CodecSep matches or exceeds AudioSep in speech and music while slightly trailing on SFX, likely reflecting differences in SFX prompt distributions (AudioSep’s diverse training vs.
+CodecSep’s compositional SFX prompts from dnr-v2).
+Importantly, our bitstream-native variant—CodecSep+dnr-v2 (codes in: codes out, zero-shot)—evaluates the CodecSep+dnr-v2 masker directly on reconstructed embeddings $E$ from code streams (§3.2, para.
+6677) without any fine-tuning: it incurs a modest drop relative to the continuous-latent path (about $1{\text{–}}2$ dB SI-SDR across sources; small ViSQOL deltas for music/speech and parity on SFX), yet still surpasses AudioSep+dnr-v2 on SI-SDR for all sources (e.g., music: 0.16-0.16 vs.
+5.55-5.55 dB; speech: $8.26$ vs.
+$7.68$ dB; SFX: 0.96-0.96 vs.
+4.66-4.66 dB).
+Compared to pretrained AudioSep (zero-shot), the codes-in:codes-out variant improves SI-SDR on speech/music but lags on SFX SI-SDR and ViSQOL.
+These results indicate that a deployment-friendly, no-finetuning bitstream path is already competitive; as discussed in §3.2, para.
+$6$ , light fine-tuning on $E$ or an embedding-consistency loss can close the residual gap.
+As part of our ablation, the lightweight CodecSep+dnr-v2 (ablate Masker) removes the transformer masker and applies FiLM directly to the encoder; it attains SI-SDR comparable to AudioSep+dnr-v2 with better perceptual speech quality, but overall separation quality drops due to FiLM perturbing the mixture latents held by the encoder.
+Effect of SFX Prompt Granularity during training.
+We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec) (cf.
+Table 2), (ii) generic 3-stem prompts (“music/speech/sfx”) (cf.
+Table 2), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions (cf.
+Table 1).
+For (ii) and (iii), we train and evaluate separate versions of both CodecSep and AudioSep on each prompt setting.
+Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality (cf.
+Table 2).
+Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts (cf.
+Table  1– 2).
+Moreover, replacing the “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement (cf.
+Table 1).
+While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
+Table 2: Results: Impact of SFX Prompt Granularity on Universal Source Separation (dnr-v2-test)
+Model Metric ($\uparrow$) Music Speech Sfx
+3-Stem: Fixed stem baselines (no text-guidance)
+TDANet SI-SDR $1.84^{\pm 3.55}$ $10.18^{\pm 2.91}$ $1.36^{\pm 4.90}$
+ViSQOL $2.85^{\pm 0.58}$ $3.09^{\pm 0.43}$ $2.44^{\pm 0.72}$
+CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
+ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
+SDCodec SI-SDR $\mathbf{1.85^{\pm 3.68}}$ $\mathbf{11.32^{\pm 2.98}}$ $\mathbf{1.77^{\pm 4.08}}$
+ViSQOL $\mathbf{2.96^{\pm 0.56}}$ $\mathbf{3.49^{\pm 0.40}}$ $\mathbf{2.64^{\pm 0.73}}$
+3-Stem: {“music”,“speech”,“sfx”} as generic prompt
+AudioSep SI-SDR $\mathbf{-2.46^{\pm 4.06}}$ $4.92^{\pm 4.21}$ $6.65^{\pm 4.73}$
+(zero-shot) ViSQOL $\mathbf{2.86^{\pm 0.63}}$ $\mathbf{3.11^{\pm 0.56}}$ $2.08^{\pm 0.68}$
+AudioSep + dnr-v2 SI-SDR $6.22^{\pm 2.77}$ $\mathbf{7.71^{\pm 3.11}}$ $2.11^{\pm 3.90}$
+ViSQOL $2.55^{\pm 0.57}$ $2.47^{\pm 0.37}$ $2.39^{\pm 0.74}$
+CodecSep + dnr-v2 SI-SDR $7.71^{\pm 2.84}$ $4.63^{\pm 2.48}$ $\mathbf{0.58^{\pm 4.15}}$
+ViSQOL $2.45^{\pm 0.55}$ $2.70^{\pm 0.49}$ $\mathbf{2.39^{\pm 0.70}}$
+Why use a Transformer masker instead of a decoder?(cf.
+Table 3).
+We compare (i) CodecFormer (decoder-style source generation), (ii) CodecSep (unguided, 3-stem) which uses the CodecFormer Transformer as a masker over codec latents, and (iii) CodecSep (text-guided).
+The results on dnr-v2-test exhibit two clear trends.
+First, replacing decoder-style generation with masking consistently strengthens separation across music, speech, and SFX.
+This aligns with our design rationale: in the DAC latent domain, the masker modulates existing, semantically structured content instead of synthesizing new signals, which (a) reduces artifacts and cross-talk leakage, (b) preserves long-range periodicity/timbre and transient organization already encoded by the codec, and (c) stabilizes optimization compared to end-to-end generation.
+Second, adding text guidance yields a further uniform improvement.
+The masker formulation concentrates Transformer capacity on selection (“where/how much” to pass) rather than generation (“what” to produce).
+Table 3: Results: Architectural advantages in using CodecFormer decoder as masker (dnr-v2-test)
+Model Metric ($\uparrow$) Music Speech Sfx
+CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
+ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
+CodecSep + dnr-v2 SI-SDR $1.15^{\pm 3.35}$ $9.90^{\pm 2.91}$ $0.82^{\pm 4.18}$
+(unguided, 3-stem) ViSQOL $2.75^{\pm 0.55}$ $3.09^{\pm 0.45}$ $\mathbf{2.48^{\pm 0.72}}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
+(text-guided) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
+Further benchmarking on ESC-50, Clotho-v2, AudioSet, VGGSound, & AudioCaps (cf. Table 4).
+Extending beyond dnr-v2, we evaluate both systems on five additional open-domain benchmarks spanning environmental sounds (ESC-50), audio captioning-style corpora (Clotho-v2, AudioCaps), weakly labeled web-scale audio (AudioSet), and visually grounded audio (VGGSound).
+Under matched training data and prompt protocols, CodecSep+dnr-v2 consistently outperforms AudioSep+dnr-v2 across all five datasets in both separation fidelity and perceptual quality.
+Table 4: Results: Benchmarking on ESC-50, Clotho-v2, AudioSet, VGGSound, AudioCaps
+Model Metric ($\uparrow$) ESC-50 Clotho-v2 AudioSet VGGSound AudioCaps
+AudioSep SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$ $6.44^{\pm 11.48}$
++ dnr-v2 ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$ $\mathbf{2.33^{\pm 1.08}}$
+CodecSep SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$ $\mathbf{-6.09^{\pm 11.62}}$
++ dnr-v2 ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$ $2.24^{\pm 1.16}$
+Robustness to prompt paraphrasing (cf.
+Table 5) .
+To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}.
+This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference.
+We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts.
+Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
+Full inference complexity (cf.
+Table 6) .
+We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep).
+Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep.
+Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space.
+Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting.
+Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth.
+Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
+Additional experiments (cf. Appendix DF) .
+For readability, three extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs.
+universal and paraphrased variants), training/evaluation splits, and metrics.
+(i) Generalization across open-domain datasets (cf.
+Appendix D): On AudioCaps (derived from AudioSet), the spectrogram baseline AudioSep benefits from distributional alignment and attains the strongest absolute scores, yet CodecSep+dnr-v2 generalizes competitively—surpassing AudioSep+dnr-v2 in SI-SDR at comparable ViSQOL.
+When both are retrained on AudioCaps, CodecSep+AudioCaps again outperforms AudioSep+AudioCaps in separation quality.
+The same trend holds on the more challenging dnr-v2 test set, where mixtures often contain speech, music, and multiple overlapping SFX; both models drop in absolute performance, but CodecSep retains its advantage in SI-SDR.
+(ii) Relative–gain summaries (cf.
+Appendix E): We report percent improvements of CodecSep over AudioSep under matched training data and prompt settings; CodecSep shows large SI-SDR gains on dnr-v2 with positive (though smaller) gains under paraphrased prompts, and consistent improvements across ESC-50, Clotho-v2, AudioSet, VGGSound, and AudioCaps, with modest ViSQOL deltas (cf.
+Table 8).
+(iii) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided).
+We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original.
+This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix F (cf.
+Table 9).
+Table 5: Results: Using ambiguous prompts for Speech and Music (dnr-v2-test)
+Model Metric ($\uparrow$) Music Speech
+AudioSep + dnr-v2 SI-SDR $6.43^{\pm 3.29}$ $4.14^{\pm 3.77}$
+ViSQOL $2.53^{\pm 0.57}$ $2.64^{\pm 0.47}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{-5.60^{\pm 3.61}}$ $\mathbf{4.19^{\pm 4.18}}$
+ViSQOL $\mathbf{2.56^{\pm 0.58}}$ $\mathbf{2.74^{\pm 0.51}}$
+Table 6: Full Inference Complexity (GMACs $\downarrow$)
+Model Input: Audio Stream Input: Code Stream Architecture-only
+AudioSep $33.5$ $73.6$ $33.5$
+CodecSep $41.45$ $\mathbf{1.35}$ ($\downarrow 54\times$) $\mathbf{1.35}$ ($\downarrow 25\times$)
+Codec GMACs used above: DAC Enc(.) =12.28=12.28, Dec(.) =27.82=27.82.
+Subjective evaluation (MOS–LQS).
+We ran a human evaluation test with n=20n{=}20 participants on $20$ dnr-v2 $3$-stem test mixtures, comparing paired outputs from CodecSep+dnr-v2 and the official AudioSep model using fixed speech/music prompts and per-clip sfx prompts.
+Raters scored each stem independently in randomized order on the MOS–LQS scale (1=bad, 5=excellent); we report mean $\pm$ $1\sigma$.
+Overall, CodecSep scored $3.34^{\pm 1.00}$} vs.
+AudioSep $2.61^{\pm 1.04}$.
+By source, CodecSep achieved $3.17^{\pm 1.01}$} (music), $3.37^{\pm 0.97}$} (sfx), and $3.49^{\pm 1.00}$} (speech), while AudioSep obtained $2.49^{\pm 0.95}$, $2.84^{\pm 1.16}$, and $2.50^{\pm 1.02}$, respectively.
+These outcomes align with objective trends (SI-SDR/ViSQOL) and indicate consistent perceptual gains for CodecSep.
+Paired model outputs and reference stems are included in the supplementary materials for side-by-side listening.
+Extensions: Multimodal prompting and high-bandwidth codecs.
+Because conditioning enters only via a fixed-dimensional query embedding $e_{\tau}$ that drives FiLM in the masker, the architecture is agnostic to the prompt modality.
+Concretely, one can replace the text encoder with (i) an audio encoder to accept exemplar audio prompts $\big(e_{\tau}^{\text{aud}}\big)$, (ii) an image/vision–language encoder (e.g., CLIP-like) to accept image prompts $\big(e_{\tau}^{\text{vis}}\big)$, or (iii) a lightweight fusion (e.g., gated additive or attention pooling) of $\big(e_{\tau}^{\text{text}},e_{\tau}^{\text{aud}},e_{\tau}^{\text{vis}}\big)$ to support mixed prompts—all without modifying the masker or the codec.
+Likewise, the separation stack is codec-agnostic: upgrading the frozen backbone from 16 kHz to full-band 44.1/48 kHz simply changes the latent shape $(d,T)$, primarily along the feature axis rather than the time dimension; the FiLM-conditioned masker and training objective remain unchanged.
+5 Conclusion
 […]
-To account for compression artifacts and improve training stability, we further compute the negative codec SI-SDR (cSI-SDR) Yip et al.
-(2024b), which is a variant of SI-SDR calculated with a codec-processed version of ground truth audio.
-During training, we keep the DAC Kumar et al.
-(2023) backbone and CLAP Devlin et al.
-(2019) text encoder weights frozen, updating only the parameters of $Mask(.)$ and $query(.)$ modules.
-3.3 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
+Limitations and clarifications.
+(1) Codec/bandwidth.
+Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic.
+The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning.
+(2) Data and prompts.
+Training data scale and prompt diversity are modest relative to open-domain audio.
+As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains.
+(3) Temporal prompting.
+While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction.
+(4) Perceptual SFX quality.
+In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work.
+Future Work.
+We will (i) scale the frozen codec backbone to higher bandwidths and stereo (44.1/48 kHz), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
+Appendix A Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
 Problem setup and pipeline contrast.
-Let $x(t)\in\mathbb{R}$ be a mono mixture with sources $\{y_{s}(t)|s\in\mathcal{S}\}$, where $x(t)$ is expressed as equation 1.
+Let $x(t)\in\mathbb{R}$ be a mono mixture with sources $\{y_{s}(t)|s\in\mathcal{S}\}$, where $x(t)$ is expressed as $x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$.
 Spectrogram-domain (AudioSep):
-$x\;\xrightarrow{\text{STFT}}\;X\!\in\!\mathbb{C}^{F\times T}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\odot|X|\exp(\angle X+\angle\hat{M}_{s})\;\xrightarrow{\text{ISTFT}}\;\tilde{y}_{s}(t),$
-where $g(\cdot,e_{\tau})$ denotes a FiLM-conditioned, complex-domain spectrogram separator that predicts a magnitude mask $|\hat{M}_{s}|\in[0,1]^{F\times T}$ and a phase residual $\angle\hat{M}_{s}$, conditioned jointly on the mixture spectrogram $X$ (obtained by the STFT of $x$) and the text embedding $e_{\tau}$.
+$x\;\xrightarrow{\text{STFT}}\;X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\odot|X|\exp(\angle X+\angle\hat{M}_{s})\;\xrightarrow{\text{ISTFT}}\;\tilde{y}_{s}(t),$
+where $g(\cdot,e_{\tau})$ denotes a FiLM-conditioned, complex-domain spectrogram separator that predicts a magnitude mask $|\hat{M}_{s}|\in[0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$, conditioned jointly on the mixture spectrogram $X$ (obtained by the STFT of $x$) and the text embedding $e_{\tau}$.
 The predicted magnitude mask and phase residual are then applied element-wise to form the source spectrogram $\tilde{Y}_{s}$ which is subsequently transformed back to the time domain via ISTFT to obtain $\tilde{y}_{s}(t)$.
@@
 Dimensionality reduction and compression.
-For a typical 1 s audio, 32 kHz STFT used by spectrogram-domain baselines, the mixture spectrogram $X\in\mathbb{C}^{F\times T}$ (e.g., $F=2\cdot 1024$, $T\approx 100$) contains about $|X|\approx 204{,}800$ complex coefficients.
-In contrast, a neural audio codec latent $Z\in\mathbb{R}^{d\times T_{c}}$ is far more compact: at 16 kHz with $d=64$ and $T_{c}=\tfrac{16000}{320}\approx 50$, |Z|=3,200|Z|=3{,}200; at 32 kHz with $T_{c}=\tfrac{32000}{320}\approx 100$, |Z|=6,400|Z|=6{,}400.
-Thus, the ratio $|Z|/|X|$ is $\approx 1.5625\times 10^{-2}$ for 16 kHz and $\approx 3.125\times 10^{-2}$ for 32 kHz , implying a $\sim 32$$64\times$ reduction in representation size.
-Operating directly in $Z$ therefore cuts memory bandwidth and MACs for the separator $Mask(\cdot)$ while retaining perceptually salient structure imparted by the codec, making latent-domain separation both compute-efficient and semantically aligned.
-Moreover, the compact channel dimension ($d=64$) enables a smaller masker (fewer parameters in $Q/K/V$ and MLP projections) and more effective self-attention.
-3.3.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
+Operating on NAC latents $Z$ slashes dimensionality while preserving perceptual factors.
+For $1\,\mathrm{s}$ audio at $32\,\mathrm{kHz}$, complex STFT with N=1024N{=}1024 and hop size M=320M{=}320 samples has $T_{\mathrm{spec}}\!\approx\!100$ frames and $F=2\times 1024$ (Re+Im) scalars per frame, so $F\cdot T_{spec}\!\approx\!204{,}800$.
+A $16\,\mathrm{kHz}$ DAC with width d=64d{=}64 and the same $M$ yields $T\!\approx\!50$ and $d\cdot T{=}64\times 50{=}3{,}200$ ($\sim 64{\times}$ smaller), shrinking $Q/K/V$ and MLP sizes and easing self-attention.
+Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs.
+A.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
 STFT vs.
@@
 Neural audio codecs (DAC) are trained with perceptual, adversarial, and codebook objectives that encourage the encoder $Enc(\cdot)$ to map $x$ into compact, structured latents:
-$Enc(.):x\mapsto Z\in\mathbb{R}^{d\times T_{c}}.$
+$Enc(.):x\mapsto Z\in\mathbb{R}^{d\times T}.$
 These latents lie on a discriminative compressed manifold $\mathcal{M}_{\text{latent}}$ in which semantically meaningful factors (e.g., pitch, timbre, transients) are disentangled and aligned for downstream use.
@@
 Hierarchical (RVQ) structure that benefits separation.
-The codec applies Residual Vector Quantization (RVQ) to $Z$, producing discrete codes $A=[a_{t}\in[K]^{N_{q}}]_{t=1}^{T_{c}}$ with $K=1024$ and $N_{q}$ quantizers.
+The codec applies Residual Vector Quantization (RVQ) to $Z$, producing discrete codes $A=[a_{t}\in[K]^{N_{q}}]_{t=1}^{T}$ with $K=1024$ and $N_{q}$ quantizers.
 Codebook lookup yields embeddings
@@
 In contrast, spectrogram systems must learn a task-specific latent from $X$ and perform separation jointly, which increases parameter count and MACs, slows convergence, and can lead to overfitting in the absence of the codec’s inductive bias.
-3.3.2 Transformer-based masker vs. decoder-style generation (specific design choice).
+A.2 Transformer-based masker vs. decoder-style generation (specific design choice).
 In CodecSep, we replace decoder-style source generation (as in CodecFormer, which directly predicts the target waveform/spectrogram) with a Transformer-based, FiLM-conditioned masker that outputs a soft mask over the codec latents.
@@
 Working in $Z$ avoids the instability commonly observed in direct waveform/spectrogram prediction, leading to faster convergence and lower training variance under the same optimization settings.
+  • Direct leverage of codec-disentangled latents.
+DAC/RVQ pretraining organizes $Z$ into a coarse-to-fine, semantically disentangled manifold (periodicity, timbre, transients).
+The masker exploits this structure by gating along these axes to isolate sources, rather than using $Z$ merely as input to a decoder that generates new latents/waveforms.
+This turns separation into selection on a well-structured space, improving identifiability while reducing parameters and data demands.
   • Minimal distortion through modulation (no hallucination).
@@
 FiLM conditioning steers the transformer to decide where/how much information to pass—not what to generate—so parameters and compute in the $Q/K/V$ and MLP projections are concentrated on selection and attenuation.
-3.3.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
+A.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
 We deliberately integrate Feature-wise Linear Modulation (FiLM) Perez et al.
-(2018) inside the transformer masker to inject text semantics while preserving the codec manifold and keeping $Enc(\cdot)$$Dec(\cdot)$ frozen.
+[2018] inside the transformer masker to inject text semantics while preserving the codec manifold and keeping $Enc(\cdot)$$Dec(\cdot)$ frozen.
 Targeted placement (masker, mid-layers).
@@
 Empirically this reduces training variance and mitigates long-horizon inconsistencies common with generator-style heads.
-3.3.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
+A.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
 Why we evaluate on continuous latents $Z$.
-We choose to perform—and therefore report—separation on the continuous DAC encoder latents $Z=Enc(x)\in\mathbb{R}^{d\times T_{c}}$ for the following concrete reasons:
+We choose to perform—and therefore report—separation on the continuous DAC encoder latents $Z=Enc(x)\in\mathbb{R}^{d\times T}$ for the following concrete reasons:
   1. 1.
@@
 For deployment scenarios where the input is a compressed bitstream, we operate on the reconstructed embeddings $E$ obtained from the codes $A$ via codebook lookup, and train the masker on $E$ instead of $Z$:
-$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T_{c}]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big{(}a_{t}^{(i)}\big{)},$ (5)
-$E$ $=[e_{t}]_{t=1}^{T_{c}}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (6)
+$A$ $=[a_{t}\in[1024]^{N_{q}}\mid t\in[T]],\qquad e_{t}\;=\;\sum_{i=1}^{N_{q}}\mathrm{lookup}\!\big(a_{t}^{(i)}\big),$ (3)
+$E$ $=[e_{t}]_{t=1}^{T}\;\approx\;Z,\qquad\tilde{E}_{s}\;=\;M_{s}\odot E,\qquad\tilde{y}_{s}(t)\;=\;Dec(\tilde{E}_{s})$ (4)
 Compressed bitstream path (codes-in, codes-out). Given $E\approx Z$ and the element-wise masking operation, the estimated source embeddings satisfy
@@
 Embedding alignment (CodecFormer-EL Yip et al.
-(2024a)).
+[2024a]).
 To tighten $\tilde{E}_{s}\approx Z_{s}$, we can optimize an embedding-level consistency loss as in CodecFormer-EL Yip et al.
-(2024a)instead of our SI-SDR objective:
+[2024a]instead of our SI-SDR objective:
 $\mathcal{L}_{\mathrm{emb}}\;=\;\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$
@@
 The deployment-ready path to discrete code streams is immediate: switch the masker input to $E$ via lookup from $A$ and train/fine-tune accordingly, with optional distribution alignment and bitrate-robustness augmentation.
-3.3.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
+A.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
 Central motivation.
@@
 Sec.
-3.3), reducing activation memory and bandwidth throughout attention and MLP blocks.
+A), reducing activation memory and bandwidth throughout attention and MLP blocks.
 Interface compatibility with codec bitstreams.
@@
   • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec.
-3.3.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
+A.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
   • Smaller working representation during separation: masking in $Z/E$ (e.g., d=64d{=}64) reduces intermediate activations, lowering memory bandwidth and enabling tighter batching.
@@
   • Maintains codec manifold structure: by modulating $Z/E$ rather than rebuilding $X$, CodecSep preserves periodicity/timbre/transients already organized by the codec, supporting stable, high-fidelity stems at deployment.
-4 Experiments
-4.1 Datasets
-4.1.1 Divide and Remaster v 2.0 (dnr-v2)
+Appendix B Dataset Details
+B.1 Divide and Remaster v 2.0 (dnr-v2)
 dnr-v2 Petermann et al.
-(2022) dataset consists of $60s$-duration artificial mixtures of speech, music, and SFX sampled from LibriSpeech Panayotov et al.
-(2015), Free Music Archive (FMA) Defferrard et al.
-(2017), and Freesound Dataset $50K$ (FSD50K50K) Fonseca et al.
-(2022), respectively.
+[2022] dataset consists of $60s$-duration artificial mixtures of speech, music, and SFX sampled from LibriSpeech Panayotov et al.
+[2015], Free Music Archive (FMA) Defferrard et al.
+[2017], and Freesound Dataset $50K$ (FSD50K50K) Fonseca et al.
+[2022], respectively.
 It includes $3,406$ ($56.7hrs$) training, $487$ ($8.13hrs$) validation, and $973$ ($16.22hrs$) test mixtures, each provided with its three individual source audios.
@@
 This results in long-form, compositional queries that reflect the structure of the mixture (e.g., “dog barking, Animal, engine rumbling, motor vehicle”).
-4.1.2 Open-Domain Benchmarks
+B.2 Open-Domain Benchmarks
 We benchmark on five open-domain datasets spanning captioned audio, environmental sounds, and large multi-event corpora: AudioCaps Kim et al.
-(2019), an AudioSet-derived collection of >46>46k $10$ s YouTube clips paired with human-written captions describing the dominant sound events (used by us to synthesize training and test mixtures); ESC-50 Piczak , a curated environmental sound dataset of 2,0002{,}000 clips ($5$ s each) organized into $50$ classes with $40$ examples per class across five meta-categories (animals, natural, human non-speech, domestic, exterior/urban); Clotho-v2 Drossos et al.
-(2020), 6,9746{,}974 audio samples (15153030 s) each annotated with five human captions (8–20 words) covering open-domain events; AudioSet-eval Gemmeke et al.
-(2017), the evaluation split of AudioSet comprising human-labeled $10$ s YouTube clips over an ontology of $632$ audio event classes in a multi-label setting; and VGGSound Chen et al.
-(2020a), an AudioSet-derived audio–visual corpus with 550+550{+} hours of $10$ s segments covering a wide variety of everyday sound categories.
-For AudioCaps we form both training and testing mixtures by summing three clips (validation segmented into $5$ s, test preserves clips up to $20$ s), while for ESC-50, Clotho-v2, AudioSet-eval, and VGGSound we construct test-only mixtures using the same three-clip protocol.
-4.2 Evaluation
-We evaluate CodecSep alongside TDANet Li et al.
-(2023); Pons et al.
-(2024), CodecFormer Yip et al.
-(2024b), SDCodec Bie et al.
-(2024), and AudioSep Liu et al.
-(2024) on universal source separation conducted on the dnr-v2 and AudioCaps test splits.
-We report the scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al.
-(2019) for signal quality assessment and Virtual Speech Quality Objective Listener (ViSQOL) Chinen et al.
-(2020) for perceptual quality assessment.
-ViSQOL Chinen et al.
-(2020) evaluates the perceptual similarity between the ground truth $x$ and the estimated signal $\tilde{x}$ using a spectro-temporal approach.
-It then predicts a Mean Opinion Score - Listening Quality Objective (MOS-LQO), which ranges from $1$ (lowest quality) to $5$ (highest quality).
-In addition, we evaluate the inference efficiency of our models in a hardware-agnostic setting by reporting Multiply-and-Accumulate operations (MACs) using Torchprofile  [^1] https://github.com/zhijian-liu/torchprofile.git averaged over 8 audios of $2s$ duration each.
-4.3 Training
+[2019], an AudioSet-derived collection of >46>46k $10$ s YouTube clips paired with human-written captions describing the dominant sound events (used by us to synthesize training and test mixtures); ESC-50 Piczak , a curated environmental sound dataset of 2,0002{,}000 clips ($5$ s each) organized into $50$ classes with $40$ examples per class across five meta-categories (animals, natural, human non-speech, domestic, exterior/urban); Clotho-v2 Drossos et al.
+[2020], 6,9746{,}974 audio samples (15153030 s) each annotated with five human captions (8–20 words) covering open-domain events; AudioSet Gemmeke et al.
+[2017], the evaluation split of AudioSet comprising human-labeled $10$ s YouTube clips over an ontology of $632$ audio event classes in a multi-label setting; and VGGSound Chen et al.
+[2020a], an AudioSet-derived audio–visual corpus with 550+550{+} hours of $10$ s segments covering a wide variety of everyday sound categories.
+For AudioCaps we form both training and testing mixtures (same scale of test data as dnr-v2) by summing three clips (validation segmented into $5$ s, test preserves clips up to $20$ s), while for ESC-50, Clotho-v2, AudioSet-eval, and VGGSound we construct test-only mixtures using the same three-clip protocol.
+Appendix C Training Details
 The complete model, including the query module $query(.)$, is trained for $400K$ iterations with DAC Kumar et al.
-(2023) and CLAP Wu et al.
-(2023) modules frozen.
+[2023] and CLAP Wu et al.
+[2023] modules frozen.
 Validation is conducted every $5K$ iterations and test every $10K$ iterations.
-We use ADAM Kingma and Ba (2017) as our optimizer and train with a batch size of $4$ examples, each $2$ seconds in duration, and a learning rate of $1.5e^{-4}$ on a single 2424GB NVIDIA A-$30$ GPU.
+We use ADAM Kingma and Ba [2017] as our optimizer and train with a batch size of $4$ examples, each $2$ seconds in duration, and a learning rate of $1.5e^{-4}$ on a single 2424GB NVIDIA A-$30$ GPU.
 Training employs a $ReduceLRonPlateau$ Mukherjee et al.
-(2019) scheduler, which reduces the learning rate by a factor of $0.5$ if the validation loss does not improve for two consecutive validation checks.
+[2019] scheduler, which reduces the learning rate by a factor of $0.5$ if the validation loss does not improve for two consecutive validation checks.
 We train two versions of CodecSep, one using the dnr-v2 dataset and the other using AudioCaps, to evaluate performance across different training distributions.
-We refer to these models using the prefixes +dnr-v2 and +AudioCaps, respectively, to indicate which dataset each model was trained on.
+We refer to these models using the suffixes +dnr-v2 and +AudioCaps, respectively, to indicate which dataset each model was trained on.
 Since TDANet and CodecFormer were originally developed for speech separation, we re-trained newly initialized versions on the dnr-v2 training set using the same configuration as CodecSep.
@@
 This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
-Table 1: Results: Separation Performance, Universal Source Separation (dnr-v2-test)
-Model Metric ($\uparrow$) Separation
-Music Speech Sfx
-AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
-(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
-AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
-ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
-ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
-CodecSep + dnr-v2 SI-SDR $6.72^{\pm 2.77}$ $1.95^{\pm 2.84}$ $6.75^{\pm 3.83}$
-(ablate Masker) ViSQOL $2.48^{\pm 0.58}$ $2.58^{\pm 0.50}$ $2.08^{\pm 0.74}$
-AudioSep + AudioCaps SI-SDR $14.86^{\pm 23.08}$ $7.11^{\pm 25.80}$ $14.63^{\pm 23.26}$
-(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
-CodecSep + AudioCaps SI-SDR $8.46^{\pm 2.78}$ $2.47^{\pm 2.91}$ $5.93^{\pm 4.33}$
-(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
-4.4 Results and Discussions
-Tables 16 present universal source separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols.
-Table 1 reports dnr-v2 test results for speech, music, and SFX: text-guided models use generic prompts for speech/music and ground-truth compositional captions for SFX; we also include a masker ablation to isolate the role of the Transformer masker.
-Table 2 studies SFX prompt granularity across three regimes—(i) fixed-stem, non-text baselines (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts ({music, speech, sfx}), and (iii) a universal setup with fine-grained, compositional SFX prompts—thereby aligning input conditions for fair comparison with fixed-head systems.
-Table 3 isolates architecture: decoder-style generation (CodecFormer) vs.
-an unguided 3-stem masker variant and its text-guided counterpart, all operating in the codec latent space.
-To assess out-of-domain generalization, Table 4 evaluates on AudioCaps-test with mixtures of three randomly sampled sources and prompts drawn from captions (not tied to fixed labels).
-Beyond AudioCaps/dnr-v2, Table 5 benchmarks on four additional open-domain corpora (ESC-50, Clotho-v2, AudioSet-eval, VGGSound) under the same prompt protocol.
-Table 6 probes prompt paraphrasing robustness by replacing the generic speech/music prompts with unseen synonymic variants at inference (zero-shot paraphrase test).
-Finally, Table 7 compares end-to-end and architecture-only GMACs for spectrogram-domain separation versus codec-latent masking, including the practical code-stream case.
-All models are evaluated against the original (uncompressed) ground truth; our methods are highlighted in bold; and we report mean and standard deviation ($1\sigma$) over each evaluation set.
-Source Separation Performance on dnr-v2 (Table 1).
-CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot scenario) and retrained AudioSep+dnr-v2 across all categories, with significant SI-SDR gains in speech ($\mathbf{9.97}$ vs.
-4.924.92/ $7.68$ dB), music ($\mathbf{1.15}$ vs.
--2.462.46/ -$5.55$ dB), and SFX ($\mathbf{0.89}$ vs.
--0.340.34/ -$4.66$ dB).
-Regarding ViSQOL, CodecSep matches or exceeds AudioSep in speech and music but slightly underperforms in SFX.
-While AudioSep is trained on a diverse dataset with sfx prompts covering isolated events and complex mixtures, CodecSep is trained exclusively on long-form, compositional SFX prompts from dnr-v2.
-Despite this narrower prompt exposure, CodecSep remains robust and achieves the highest overall SI-SDR.
-As part of our ablation study, we evaluate a lightweight variant—CodecSep+dnr-v2 (ablate Masker)—that removes the transformer masker and applies text conditioning directly to the encoder.
-While it performs comparable to AudioSep+dnr-v2 in SI-SDR, it achieves better perceptual quality for speech.
-The drop in separation quality stems from FiLM distorting the encoder’s latent space, which holds the mixture representation.
-Effect of SFX Prompt Granularity during training (Table 2).
-We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec), (ii) generic 3-stem prompts (“music/speech/sfx”), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions.
-Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality, despite those baselines being explicitly supervised per head.
-Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses the spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts.
-Moreover, replacing the single “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement.
-While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
-Table 2: Results: Impact of SFX Prompt Granularity on Universal Source Separation (dnr-v2-test)
-Model Metric ($\uparrow$) Separation
-Music Speech Sfx
-3-Stem: Fixed stem baselines (no text-guidance)
-TDANet SI-SDR $1.84^{\pm 3.55}$ $10.18^{\pm 2.91}$ $1.36^{\pm 4.90}$
-ViSQOL $2.85^{\pm 0.58}$ $3.09^{\pm 0.43}$ $2.44^{\pm 0.72}$
-CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
-ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
-SDCodec SI-SDR $\mathbf{1.85^{\pm 3.68}}$ $\mathbf{11.32^{\pm 2.98}}$ $\mathbf{1.77^{\pm 4.08}}$
-ViSQOL $\mathbf{2.96^{\pm 0.56}}$ $\mathbf{3.49^{\pm 0.40}}$ $\mathbf{2.64^{\pm 0.73}}$
-3-Stem: {“music”,“speech”,“sfx”} as generic prompt
-AudioSep SI-SDR $\mathbf{-2.46^{\pm 4.06}}$ $4.92^{\pm 4.21}$ $6.65^{\pm 4.73}$
-(zero-shot) ViSQOL $\mathbf{2.86^{\pm 0.63}}$ $\mathbf{3.11^{\pm 0.56}}$ $2.08^{\pm 0.68}$
-AudioSep + dnr-v2 SI-SDR $6.22^{\pm 2.77}$ $\mathbf{7.71^{\pm 3.11}}$ $2.11^{\pm 3.90}$
-ViSQOL $2.55^{\pm 0.57}$ $2.47^{\pm 0.37}$ $2.39^{\pm 0.74}$
-CodecSep + dnr-v2 SI-SDR $7.71^{\pm 2.84}$ $4.63^{\pm 2.48}$ $\mathbf{0.58^{\pm 4.15}}$
-ViSQOL $2.45^{\pm 0.55}$ $2.70^{\pm 0.49}$ $\mathbf{2.39^{\pm 0.70}}$
-Universal: “music”,“speech”, fine-grained sfx prompt
-AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
-(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
-AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
-ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
-ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
-Why use a Transformer masker instead of a decoder?(Table 3).
-We compare (i) CodecFormer (decoder-style source generation), (ii) CodecSep (unguided, 3-stem) which uses the CodecFormer Transformer as a masker over codec latents, and (iii) CodecSep (text-guided).
-The results on dnr-v2-test exhibit two clear trends.
-First, replacing decoder-style generation with masking consistently strengthens separation across music, speech, and SFX.
-This aligns with our design rationale: in the DAC latent domain, the masker modulates existing, semantically structured content instead of synthesizing new signals, which (a) reduces artifacts and cross-talk leakage, (b) preserves long-range periodicity/timbre and transient organization already encoded by the codec, and (c) stabilizes optimization compared to end-to-end generation.
-Second, adding text guidance yields a further uniform improvement.
-The masker formulation concentrates Transformer capacity on selection (“where/how much” to pass) rather than generation (“what” to produce).
-Taken together, these findings support our design choice: in the codec latent space, a FiLM-conditioned Transformer masker is a more effective and robust allocator of model capacity than a decoder, with text guidance further sharpening that allocation.
-Table 3: Results: Architectural advantages in using CodecFormer decoder as masker (dnr-v2-test)
-Model Metric ($\uparrow$) Separation
-Music Speech Sfx
-CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
-ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
-CodecSep + dnr-v2 SI-SDR $1.15^{\pm 3.35}$ $9.90^{\pm 2.91}$ $0.82^{\pm 4.18}$
-(unguided, 3-stem) ViSQOL $2.75^{\pm 0.55}$ $3.09^{\pm 0.45}$ $\mathbf{2.48^{\pm 0.72}}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
-(text-guided) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
-Table 4: Results: Generalization, Universal Source Separation (AudioCaps-test)
+Appendix D Cross-Benchmark Performance: AudioCaps and dnr-v2.
+Table 7(a) reports generalization results on the AudioCaps-test set derived from AudioSet.
+AudioSep benefits from distributional alignment, having been trained on diverse datasets, including AudioSet, and consequently achieves the strongest separation performance.
+CodecSep+dnr-v2 generalizes well to AudioCaps and outperforms AudioSep+dnr-v2 in SI-SDR while maintaining competitive ViSQOL scores.
+When retrained on AudioCaps, CodecSep again outperforms AudioSep in separation quality, demonstrating strong cross-domain robustness.
+Table 7(b) further supports this trend on the more challenging dnr-v2 test set, where CodecSep+AudioCaps outperforms AudioSep+AudioCaps in SI-SDR across all sources while maintaining comparable perceptual quality.
+However, both models experience a performance drop on dnr-v2 due to its increased mixture complexity, often containing speech, music, and three or more overlapping SFX sources—making it significantly more challenging than the simpler mixtures seen during AudioCaps training.
+Table 7: Generalization and transfer results for universal source separation.
+(a) Generalization on AudioCaps-test.
 Model Separation
@@
 CodecSep + AudioCaps $\mathbf{-6.19^{\pm 10.58}}$ $2.14^{\pm 1.00}$
-Table 5: Results: Further Benchmarking on ESC-50, Clotho-v2, AudioSet-eval, VGGSound
-Model Metric ($\uparrow$) Separation
-ESC-50 Clotho-v2 AudioSet-eval VGGSound
-AudioSep + dnr-v2 SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$
-ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$
-CodecSep + dnr-v2 SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$
-ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$
-Generalization across open-domain datasets.
-Table 4 reports generalization results on the AudioCaps-test set derived from AudioSet.
-AudioSep benefits from distributional alignment, having been trained on diverse datasets, including AudioSet, and consequently achieves the strongest separation performance.
-CodecSep+dnr-v2 generalizes well to AudioCaps and outperforms AudioSep+dnr-v2 in SI-SDR while maintaining competitive ViSQOL scores.
-When retrained on AudioCaps, CodecSep again outperforms AudioSep in separation quality, demonstrating strong cross-domain robustness.
-Table 1 further supports this trend on the more challenging dnr-v2 test set, where CodecSep+AudioCaps outperforms AudioSep+AudioCaps in SI-SDR across all sources while maintaining comparable perceptual quality.
-However, both models experience a performance drop on dnr-v2 due to its increased mixture complexity, often containing speech, music, and three or more overlapping SFX sources—making it significantly more challenging than the simpler mixtures seen during AudioCaps training.
-Further benchmarking on ESC-50, Clotho-v2, AudioSet-eval, and VGGSound (Table 5).
-Extending beyond AudioCaps and dnr-v2, we evaluate both systems on four additional open-domain benchmarks spanning environmental sound classification (ESC-50), audio captioning-style corpora (Clotho-v2), weakly labeled web-scale audio (AudioSet-eval), and visually grounded audio (VGGSound).
-Under matched training data and prompt protocols, CodecSep+dnr-v2 consistently outperforms AudioSep+dnr-v2 across all four datasets in both separation fidelity and perceptual quality.
-Taken together with the dnr-v2 analyses, these results indicate that CodecSep’s latent-domain masking—leveraging the codec’s structured, semantically aligned representations—confers robust cross-domain behavior while using $25\times$ less architecture-only compute than spectrogram separators.
-This combination of accuracy and efficiency is particularly attractive in practical deployments, where models must generalize to diverse sound taxonomies under strict compute and memory budgets.
-Robustness to prompt paraphrasing (Table 6) .
-To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}.
-This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference.
-We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts.
-Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
-Table 6: Results: Impact of prompt ambiguity (dnr-v2-test)
-Model Metric ($\uparrow$) Separation
-Music Speech
-Using Generic Prompts for Speech and Music
-AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$
-ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$
-ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$
-Using ambiguous prompts for Speech and Music
-AudioSep + dnr-v2 SI-SDR $6.43^{\pm 3.29}$ $4.14^{\pm 3.77}$
-ViSQOL $2.53^{\pm 0.57}$ $2.64^{\pm 0.47}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{-5.60^{\pm 3.61}}$ $\mathbf{4.19^{\pm 4.18}}$
-ViSQOL $\mathbf{2.56^{\pm 0.58}}$ $\mathbf{2.74^{\pm 0.51}}$
-Table 7: Full Inference Complexity (GMACs $\downarrow$)
-Model Input: Audio Stream Input: Code Stream Architecture-only
-AudioSep $33.5$ $73.6$ $33.5$
-CodecSep $41.45$ $\mathbf{1.35}$ ($\downarrow 54\times$) $\mathbf{1.35}$ ($\downarrow 25\times$)
-Codec GMACs used above: DAC Enc(.) =12.28=12.28, Dec(.) =27.82=27.82.
-Full inference complexity (Table 7) .
-We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep).
-Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep.
-Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space.
-Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting.
-Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth.
-Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
-Relative–gain summaries (Appendix A).
-For a compact view of percent improvements, Appendix A Table 8 reports relative gains of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text.
-In brief: on dnr-v2, CodecSep yields large improvements in SI-SDR with clear perceptual lifts; under paraphrased prompts, gains are smaller but remain positive, indicating robustness to lexical variation; across ESC-50, Clotho-v2, AudioSet-eval, and VGGSound, SI-SDR gains are consistent while ViSQOL deltas are modest; and when trained on AudioCaps, CodecSep maintains an advantage on AudioCaps-test and shows strong improvements when evaluated on dnr-v2.
-Reconstruction study (Appendix B).
-We probe source leakage with a single-source reconstruction diagnostic on dnr-v2-test : each input contains one target source and models are either prompted with the matching class/label (text-guided) or routed to the corresponding fixed head (non–text-guided).
-We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original.
-This is a leakage/consistency check—not a primary separation metric; full setup/results are in the Appendix B (Table 9).
-5 Conclusion
-[…]
-Limitations and clarifications.
-(1) Codec/bandwidth.
-Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic.
-The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning.
-(2) Data and prompts.
-Training data scale and prompt diversity are modest relative to open-domain audio.
-As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains.
-(3) Temporal prompting.
-While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction.
-(4) Perceptual SFX quality.
-In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work.
-(5) Perceptual evaluation.
-Following CodecFormer, we report ViSQOL as a proxy for listening quality (predicting MOS-LQO on a 1–5 scale).
-Although we did not conduct a formal multi-listener study, we include real-world audio outputs for both CodecSep and AudioSep in the supplementary material, together with a Markdown summary of an informal author listening pass.
-Future Works.
-We will scale CodecSep with higher-bandwidth/stereo codecs, broaden prompt coverage (including temporal and referring-expression cues), explore stronger self-alignment objectives, and investigate alternative conditioning (audio-only or multimodal) within the same FiLM-masking framework.
-Combined with its large compute reduction and code-stream compatibility, these extensions position CodecSep as a practical foundation for fast, universal, and deployable prompt-guided source separation.
-Appendix A Discussion of Relative–Gain Summaries.
+(b) Transfer to dnr-v2-test when trained on AudioCaps (zero-shot on dnr-v2).
+Model Metric ($\uparrow$) Music Speech Sfx
+AudioSep + AudioCaps SI-SDR $14.86^{\pm 23.08}$ $7.11^{\pm 25.80}$ $14.63^{\pm 23.26}$
+(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
+CodecSep + AudioCaps SI-SDR $8.46^{\pm 2.78}$ $2.47^{\pm 2.91}$ $5.93^{\pm 4.33}$
+(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
+Appendix E Discussion of Relative–Gain Summaries.
 Tables 8(a)8(d) consolidate relative improvements of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text.
-On DnR-v2 (Table 8(a)), CodecSep delivers large SI-SDR gains—especially for music and SFX—together with a clear perceptual lift.
-Under paraphrased prompts (Table 8(b)), gains are smaller but remain positive, indicating robustness to lexical variation.
-Across additional open-domain benchmarks (Table 8(c)), SI-SDR gains are consistent while ViSQOL deltas are modest, aligning with cross-domain trends reported earlier.
-Finally, when trained on AudioCaps (Table 8(d)), CodecSep maintains an advantage on AudioCaps-test and yields strong improvements on DnR-v2, supporting the claim that codec-latent masking generalizes well across datasets and prompt regimes.
+On dnr-v2 (cf.
+Table 8(a)), CodecSep delivers large SI-SDR gains—especially for music and SFX—together with a clear perceptual lift.
+Under paraphrased prompts (cf.
+Table 8(b)), gains are smaller but remain positive, indicating robustness to lexical variation.
+Across additional open-domain benchmarks (cf.
+Table 8(c)), SI-SDR gains are consistent while ViSQOL deltas are modest, aligning with cross-domain trends reported earlier.
+Finally, when trained on AudioCaps (cf.
+Table 8(d)), CodecSep maintains an advantage on AudioCaps-test and yields strong improvements on dnr-v2, supporting the claim that codec-latent masking generalizes well across datasets and prompt regimes.
 Table 8: Relative gains (%) of CodecSep over AudioSep under matched training/prompt settings. Each sub-table reports percent improvements for a specific evaluation setup.
@@
 ViSQOL 6.5-6.5 3.8-3.8 $\mathbf{+5.4}$ -4.2
-Appendix B Further Studies: Reconstruction Performance.
+Appendix F Further Studies: Reconstruction Performance.
 Table 9 assesses performance under a single-source reconstruction setting on the dnr-v2-test set, where each model is prompted to reproduce the input source.

### v3

@@
-Neural Audio Codecs for Prompt-Driven Universal Source Separation
+Neural Audio Codecs for Prompt-Driven Universal Sound Separation
 Abstract
@@
 1 Introduction
-We propose CodecSep, a text-conditioned universal source separation (USS) framework that marries the interpretability of prompt-driven extraction with the efficiency of neural audio codecs (NACs).
+We propose CodecSep, a text-conditioned universal sound separation (USS) framework that marries the interpretability of prompt-driven extraction with the efficiency of neural audio codecs (NACs).
 To our knowledge, CodecSep is the first system to bridge NACs with USS: it conditions a transformer masker on CLAP text embeddings Wu et al.
@@
 (2024).
-Prior codec–separation hybrids (CodecFormer, SDCodec) Yip et al.
-(2024b); Bie et al.
-(2024) demonstrate light-weight, high-fidelity separation but target fixed stems (e.g., speech separation or speech vs.
+Prior codec–separation hybrids (CodecFormer Yip et al.
+(2024b), SDCodec Bie et al.
+(2024)) are lightweight and high-fidelity but target fixed stems (e.g., speech separation or speech vs.
 music vs.
-SFX) rather than open-domain prompt-conditioned USS.
+SFX); extending them to open-domain, prompt-conditioned USS is non-trivial (cf.
+§2, para.
+3).
 CodecSep adopts a frozen DAC encoder–decoder backbone and inserts a FiLM-conditioned transformer masker that predicts a soft mask over codec latents.
@@
 Operating on compact codec features cuts memory traffic and MACs compared to spectrogram-domain pipelines, while preserving the codec’s inductive biases (periodicity, timbre, transients).
-In doing so, CodecSep delivers interpretable, prompt-guided separation with markedly lower compute, without sacrificing separation fidelity.
-We evaluate CodecSep for universal source separation under a comprehensive set of protocols: (i) dnr-v2 for in-domain text-guided separation; (ii) five open-domain corpora—AudioCaps, ESC-50, Clotho-v2, AudioSet-eval, and VGGSound—for cross-domain generalization; (iii) three prompt granularities (fixed-stem, generic three-stem {“speech”, “music”, “sfx”}, and a universal setting with fine-grained, compositional SFX descriptions); and (iv) paraphrase robustness via unseen synonyms for speech/music.
+In doing so, CodecSep delivers interpretable, prompt-guided separation with markedly lower compute without sacrificing separation fidelity.
+Crucially, conditioning via text embeddings enables open-vocabulary operation of NAC-based separation.
+We evaluate CodecSep for universal sound separation under a comprehensive set of protocols: (i) dnr-v2 for in-domain text-guided separation; (ii) five open-domain corpora—AudioCaps, ESC-50, Clotho-v2, AudioSet-eval, and VGGSound—for cross-domain generalization; (iii) three prompt granularities (fixed-stem, generic three-stem {“speech”, “music”, “sfx”}, and a universal setting with fine-grained, compositional SFX descriptions); and (iv) paraphrase robustness via unseen synonyms for speech/music.
 We also include an architectural ablation contrasting decoder-style source generation (CodecFormer Yip et al.
@@
 2 Related Work
-Classical audio source separation systems frequently adopt an encoder–masker–decoder design in which an encoder produces STFT-like latents, a masker predicts source-specific masks, and a decoder reconstructs waveforms.
+Classical sound separation systems frequently adopt an encoder–masker–decoder design in which an encoder produces STFT-like latents, a masker predicts source-specific masks, and a decoder reconstructs waveforms.
 Representative models include DPTNet Chen et al.
@@
 (2020)), underscoring the breadth of design choices across time and frequency domains.
-Moving from domain-specific separation to universal source separation (USS), supervised systems often rely on Permutation Invariant Training (PIT) Yu et al.
+Moving from domain-specific separation to universal sound separation (USS), supervised systems often rely on Permutation Invariant Training (PIT) Yu et al.
 (2017); Kavalerov et al.
@@
 However, both typically assume a fixed maximum number of sources and emit all estimates indiscriminately, necessitating a separate identification step.
-Query-Guided Source Separation (QSS) addresses this limitation by conditioning extraction on external queries—visual, exemplar audio, class labels, or text.
+Query-Guided Sound Separation (QSS) addresses this limitation by conditioning extraction on external queries—visual, exemplar audio, class labels, or text.
 Text queries are compact and expressive, capture high-level semantics, and avoid costly references.
@@
 Neural audio codecs (NACs) have recently been integrated into separation pipelines to improve efficiency.
-CodecFormer Yip et al.
-(2024b) performs separation directly in DAC Kumar et al.
-(2023) latent space with a transformer separator trained using negative SI-SDR, and CodecFormer-EL Yip et al.
-(2024a) further adds an embedding-level objective that aligns separator outputs with DAC encoder latents.
-SD-Codec Bie et al.
-(2024) embeds separation behavior inside the codec by assigning dedicated RVQ branches to speech, music, and SFX, summing their codes to form mixture-level representations while enabling per-stem decoding.
-In contrast, our work repurposes the CodecFormer transformer into a text-conditioned mask predictor operating in the NAC latent domain: CLAP-driven FiLM conditioning modulates intermediate transformer activations to estimate a single, prompt-guided mask, enabling efficient, controllable USS without fixed source heads.
+CodecFormer separates directly in DAC Kumar et al.
+(2023) latent space with a transformer trained using negative SI-SDR, and CodecFormer-EL Yip et al.
+(2024a) adds an embedding-level objective to align separator outputs with encoder latents.
+SDCodec Bie et al.
+(2024) embeds separation inside the codec by assigning dedicated RVQ branches to speech, music, and SFX and summing their codes to form mixture representations.
+However, neither design trivially extends to open-domain, prompt-conditioned USS: SDCodec’s hardwired, per-stem RVQ branches do not scale to open vocabularies (adding branches explodes parameters, while a mixture-invariant reformulation collapses into “three generic codecs” with no stem-specific RVQ specialization).
+MixIT-style training of CodecFormer still presupposes a maximum number of sources, conflicting with universal separation.
 Figure 1: An overview of CodecSep. (Left) The full pipeline for text-guided USS. (Right) The integration of text conditioning into intermediate layers of transformer masker via FiLM layers.
@@
 CodecSep adapts the 16 kHz DAC Kumar et al.
-(2023) codec backbone for text-driven universal source separation (USS).
+(2023) codec backbone for text-driven universal sound separation (USS).
 We use a transformer masker that estimates soft masks over codec latents and inject text conditioning via Feature-wise Linear Modulation (FiLM) Perez et al.
@@
 Compared to STFT-domain AudioSep, operating in compact codec latents yields markedly lower compute and is amenable to edge deployment.
-3.1 Our Model
-Backbone codec.
+3.1 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
+Descript Audio Codec (DAC) Backbone.
 We use DAC Kumar et al.
@@
 Given $Z$, $Quant(\cdot)$ yields discrete codes $A{=}[a_{t}\!\in\![1024]^{8}]$, which map to embeddings $e_{t}{=}\sum_{i=1}^{8}e^{i}_{t}$; $Dec(\cdot)$ upsamples $E{=}[e_{t}]$ back to waveform $y(t)$.
-Transformer masker and text-guided separation.
-We consider a mono mixture as $x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$, where $x(t)$ is the observed waveform, $\mathcal{S}$ is the (unbounded) set of source classes/instances present, and $y_{s}(t)$ is the waveform of source $s$.
-Passing $x(t)$ through the frozen DAC encoder yields codec latents $Z\in\mathbb{R}^{d\times T}$ ($d$: channel width, $T$: latent frames).
-The masker $Mask(\cdot)$ operates in this latent space to predict an element-wise mask that selects the target source.
-We adopt a CodecFormer-style transformer with L=16L{=}16 layers, width d=256d{=}256, and Snake activations.
-Given a natural-language query $\tau$ (e.g., “dog barking”, “speech and music”), we compute a CLAP text embedding $e_{\tau}\in\mathbb{R}^{d}$.
-A lightweight query network $query(.)$ maps $e_{\tau}$ to per-layer FiLM parameters $(\gamma^{l},\beta^{l})\in\mathbb{R}^{d}$ for $l\in\{2,\ldots,L{-}1\}$, applied channel-wise to intermediate activations $H^{l}\in\mathbb{R}^{d\times T}$: $\tilde{H}^{l}=\mathrm{FiLM}(H^{l};\gamma^{l},\beta^{l})=\gamma^{l}\odot H^{l}+\beta^{l}$.
-The final transformer output $H^{L}$ is then passed through a single 1D convolutional head to produce the prompt-conditioned mask $M_{s}\in[0,1]^{d\times T}$.
-We obtain source latents $\tilde{Z}_{s}=M_{s}\odot Z$ and decode with the frozen codec decoder to get the waveform estimate $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$, bypassing RVQ lookup.
-Training Objective
-[…]
-3.2 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
-Spectrogram-domain (AudioSep):
-$x\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
-where $g(\cdot,e_{\tau})$ is a FiLM-conditioned separator that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$ from $X$ and text embedding $e_{\tau}$.
-CodecSep (NAC latent-domain):
-$x\xrightarrow{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
-with frozen DAC backbone $Enc(\cdot),Dec(\cdot)$ and a FiLM-conditioned transformer masker $Mask(\cdot,e_{\tau})$ that estimates $M_{s}\!\in\![0,1]^{d\times T}$.
+Text-guided Sound Separation in Spectrogram-domain (AudioSep):
+$x(t)\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
+where $g(\cdot,e_{\tau})$ is a FiLM-conditioned masker that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$ given the complex STFT $X$ of audio $x(t)$ and text-embedding $e_{\tau}$.
+Text-guided Sound Separation in NAC latent-domain (CodecSep):
+$x(t)\xrightarrow{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
+with frozen DAC backbone $Enc(\cdot),Dec(\cdot)$ and a FiLM-conditioned transformer masker $Mask(\cdot,e_{\tau})$ that estimates mask $M_{s}\!\in\![0,1]^{d\times T}$ applied element-wise to DAC latent $Z$.
 Why NAC latents vs. spectrograms.
@@
 discrete $A{\to}E$ operation (and codes-in/out), and complexity/deployment comparisons.
+3.2 Our Model
+Masking (Not Generating): FiLM-Conditioned Transformer masker over NAC Latents for text-guided sound separation.
+We consider a mono mixture as $x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$, where $x(t)$ is the observed waveform, $\mathcal{S}$ is the (unbounded) set of source classes/instances present, and $y_{s}(t)$ is the waveform of source $s$.
+Passing $x(t)$ through the frozen DAC encoder yields codec latents $Z\in\mathbb{R}^{d\times T}$ ($d$: channel width, $T$: latent frames).
+The masker $Mask(\cdot)$ operates in this latent space to predict an element-wise mask that selects the target source.
+We adopt a CodecFormer-style transformer with L=16L{=}16 layers, width d=256d{=}256, and Snake activations.
+Given a natural-language query $\tau$ (e.g., “dog barking”, “speech and music”), we compute a CLAP text embedding $e_{\tau}\in\mathbb{R}^{d}$.
+A lightweight query network $query(.)$ maps $e_{\tau}$ to per-layer FiLM parameters $(\gamma^{l},\beta^{l})\in\mathbb{R}^{d}$ for $l\in\{2,\ldots,L{-}1\}$, applied channel-wise to intermediate activations $H^{l}\in\mathbb{R}^{d\times T}$: $\tilde{H}^{l}=\mathrm{FiLM}(H^{l};\gamma^{l},\beta^{l})=\gamma^{l}\odot H^{l}+\beta^{l}$.
+The final transformer output $H^{L}$ is then passed through a single 1D convolutional head to produce the prompt-conditioned mask $M_{s}\in[0,1]^{d\times T}$.
+We obtain source latents $\tilde{Z}_{s}=M_{s}\odot Z$ and decode with the frozen codec decoder to get the waveform estimate $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$, bypassing RVQ lookup.
+Training Objective
+[…]
 4 Experiments
@@
 For fair comparison, TDANet and CodecFormer are re-trained from scratch on dnr-v2 using our setup, AudioSep is evaluated both as a public checkpoint and when re-trained under matched protocols, and SDCodec is used as released by the authors.
-To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through the codec during both training and inference.
+To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through a full-band stereo-capable $48$ kHz EnCodec during both training and inference.
 Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix C.
-Table 1: Results: Separation Performance, Universal Source Separation (dnr-v2-test)
+Table 1: Results: Separation Performance, Universal Sound Separation (dnr-v2-test)
 Model Metric ($\uparrow$) Music Speech Sfx
@@
 4.1 Results and Discussions
-Tables 15 present universal source separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols.
+Tables 15 present universal sound separation results, prompt–granularity analyses, architectural ablations, cross-dataset generalization, paraphrase robustness, and full inference complexity under matched training/evaluation protocols.
 Table 1 reports dnr-v2 test results for speech, music, and SFX: text-guided models use generic prompts for speech/music and ground-truth compositional captions for SFX; we also include a masker ablation to isolate the role of the Transformer masker.
@@
 All models are evaluated against the original (uncompressed) ground truth; our methods are highlighted in bold; and we report mean and standard deviation ($1\sigma$).
-Source Separation Performance on dnr-v2 (cf.
+Sound Separation Performance on dnr-v2 (cf.
 Table 1).
@@
 While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
-Table 2: Results: Impact of SFX Prompt Granularity on Universal Source Separation (dnr-v2-test)
+Table 2: Results: Impact of SFX Prompt Granularity on Universal Sound Separation (dnr-v2-test)
 Model Metric ($\uparrow$) Music Speech Sfx
@@
 Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
-Additional experiments (cf. Appendix DF) .
+Additional experiments (cf. Appendix DG) .
 For readability, three extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs.
@@
 Table 8).
-(iii) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided).
+(iii) Higher bandwidth (cf.
+Appendix F): Swapping the frozen 16 kHz DAC for a 48 kHz (stereo–capable) EnCodec preserves the masking architecture and deployment path.
+Higher $F_{s}$ lengthens $T$ and adds high-frequency detail, making separation harder and lowering absolute SI-SDR/ViSQOL (cf.
+Table 9).
+Improvements hinge on backbone capacity/data—not on redesign.
+(iv) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided).
 We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original.
-This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix F (cf.
+This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix G (cf.
 Table 10).
@@
 Paired model outputs and reference stems are included in the supplementary materials for side-by-side listening.
-Extensions: Multimodal prompting and high-bandwidth codecs.
+Extension to multi-modal prompting.
 Because conditioning enters only via a fixed-dimensional query embedding $e_{\tau}$ that drives FiLM in the masker, the architecture is agnostic to the prompt modality.
-Concretely, one can replace the text encoder with (i) an audio encoder to accept exemplar audio prompts $\big(e_{\tau}^{\text{aud}}\big)$, (ii) an image/vision–language encoder (e.g., CLIP-like) to accept image prompts $\big(e_{\tau}^{\text{vis}}\big)$, or (iii) a lightweight fusion (e.g., gated additive or attention pooling) of $\big(e_{\tau}^{\text{text}},e_{\tau}^{\text{aud}},e_{\tau}^{\text{vis}}\big)$ to support mixed prompts—all without modifying the masker or the codec.
-Likewise, the separation stack is codec-agnostic: upgrading the frozen backbone from 16 kHz to full-band 44.1/48 kHz simply changes the latent shape $(d,T)$, primarily along the feature axis rather than the time dimension; the FiLM-conditioned masker and training objective remain unchanged.
+Concretely, one can replace the text encoder with (i) an audio encoder to accept audio prompts $\big(e_{\tau}^{\text{aud}}\big)$, (ii) an image/vision–language encoder (e.g., CLIP) to accept image prompts $\big(e_{\tau}^{\text{vis}}\big)$, or (iii) a lightweight fusion (e.g., gated additive or attention pooling) of $\big(e_{\tau}^{\text{text}},e_{\tau}^{\text{aud}},e_{\tau}^{\text{vis}}\big)$ to support mixed prompts—all without modifying the masker or the codec.
 5 Conclusion
@@
 Limitations and clarifications.
-(1) Codec/bandwidth.
-Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic.
-The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning.
 (1) Data and prompts.
@@
 Future Work.
-We will (i) scale the frozen codec backbone to higher bandwidths and stereo (44.1/48 kHz), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
+We will (i) extend separation to higher bandwidths and stereo/spatial audio (e.g., 48 kHz stereo Encodec; HO–DirAC Hold et al.
+(2024), SpatialCodec Xu et al.
+(2024)), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
 Appendix A Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
@@
 We also include SDCodec, using the official pre-trained models made available by the authors.
-To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing during training and inference.
+To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing with a full-band stereo-capable $48$ kHz EnCodec during training and inference.
 This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
@@
 However, both models experience a performance drop on dnr-v2 due to its increased mixture complexity, often containing speech, music, and three or more overlapping SFX sources—making it significantly more challenging than the simpler mixtures seen during AudioCaps training.
-Table 7: Generalization and transfer results for universal source separation.
+Table 7: Generalization and transfer results for universal sound separation.
 (a) Generalization on AudioCaps-test.
@@
 ViSQOL 6.5-6.5 3.8-3.8 $\mathbf{+5.4}$ -4.2
-Appendix F Further Studies: Reconstruction Performance.
+Appendix F Bandwidth scaling: extending CodecSep to 48 kHz full-band
+[…]
+Table 9: Results: Extending CodecSep to $48$ kHz full-band (dnr-v2-test)
+Model Sampling Rate Metric ($\uparrow$) Music Speech Sfx
+AudioSep $32$ kHz SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
+(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
+AudioSep + dnr-v2 $32$ kHz SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
+ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
+CodecSep + dnr-v2 $16$ kHz SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
+(DAC Backbone) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
+CodecSep + dnr-v2 $48$ kHz SI-SDR $2.8^{\pm 3.5}$ $5.36^{\pm 2.36}$ $0.46^{\pm 3.83}$
+(EnCodec Backbone) ViSQOL $2.39^{\pm 0.5}$ $2.63^{\pm 0.41}$ $2.37^{\pm 0.65}$
+Appendix G Further Studies: Reconstruction Performance.
 Table 10 assesses performance under a single-source reconstruction setting on the dnr-v2-test set, where each model is prompted to reproduce the input source.
@@
 However, strong mixture reconstruction despite poor separation suggests source leakage across separated outputs.
-Table 9: Results: Reconstruction Performance, Universal Source Separation (dnr-v2-test)
+Table 10: Results: Reconstruction Performance, Universal Sound Separation (dnr-v2-test)
 Model Metric ($\uparrow$) Reconstruction

### v4

@@
 Crucially, conditioning via text embeddings enables open-vocabulary operation of NAC-based separation.
-We evaluate CodecSep for universal sound separation under a comprehensive set of protocols: (i) dnr-v2 for in-domain text-guided separation; (ii) five open-domain corpora—AudioCaps, ESC-50, Clotho-v2, AudioSet-eval, and VGGSound—for cross-domain generalization; (iii) three prompt granularities (fixed-stem, generic three-stem {“speech”, “music”, “sfx”}, and a universal setting with fine-grained, compositional SFX descriptions); and (iv) paraphrase robustness via unseen synonyms for speech/music.
-We also include an architectural ablation contrasting decoder-style source generation (CodecFormer Yip et al.
-(2024b)) with our Transformer masker over codec latents, plus a reconstruction diagnostic (Appendix) to probe source leakage and mixture consistency.
 […]
@@
 (2020), Spleeter Hennequin et al.
-(2020)), underscoring the breadth of design choices across time and frequency domains.
-Moving from domain-specific separation to universal sound separation (USS), supervised systems often rely on Permutation Invariant Training (PIT) Yu et al.
+(2020)), underscoring the breadth of design choices.
+Moving from domain-specific separation to universal sound separation (USS), supervised systems typically rely on Permutation Invariant Training (PIT) Yu et al.
 (2017); Kavalerov et al.
-(2019), while unsupervised approaches such as MixIT Wisdom et al.
-(2020) learn from mixtures alone.
-However, both typically assume a fixed maximum number of sources and emit all estimates indiscriminately, necessitating a separate identification step.
-Query-Guided Sound Separation (QSS) addresses this limitation by conditioning extraction on external queries—visual, exemplar audio, class labels, or text.
-Text queries are compact and expressive, capture high-level semantics, and avoid costly references.
+(2019), while unsupervised methods such as MixIT Wisdom et al.
+(2020) learn directly from mixtures.
+Both paradigms assume a fixed maximum number of sources and output all estimates indiscriminately, requiring a post-hoc identification step (cf.
+Appendix A for detailed failure modes).
+A recent PIT-trained USS model is Sudo rm-rf Tzinis et al.
+(2022a), a parameter-efficient time-domain separator that is based on ConvTasNet-style encoder–masker–decoder architecture with adaptive encoder/decoder modules.
+It downsamples input waveform to STFT-like latents before separation and is PIT-trained to separate mixtures with up to four sources.
+Query-Guided Sound Separation (QSS) addresses this limitation of PIT or MixIT models by conditioning extraction on external queries—visual cues, audio, class labels, or text.
+Text queries are compact, expressive, and capture high-level semantics without requiring additional reference signals.
 AudioSep Liu et al.
-(2024) exemplifies this direction by injecting BERT Devlin et al.
+(2024) follows this direction by injecting BERT Devlin et al.
 (2019) or CLAP Wu et al.
 (2023) embeddings via FiLM Perez et al.
-(2018) at intermediate masker layers to steer separation toward the requested source.
+(2018) at intermediate masker layers to steer separation toward the query source.
+BiModalSS Mahmud et al.
+(2024) extends AudioSep with attention-based conditioning and more efficient training strategies.
+FiLM-conditioned variants of Sudo rm-rf Tzinis et al.
+(2022b, 2023) have also been explored for class-guided separation using one-hot or multi-hot conditioning vectors; however, such label-based conditioning does not extend naturally to open-domain text queries and cannot handle unseen classes.
 Neural audio codecs (NACs) have recently been integrated into separation pipelines to improve efficiency.
@@
 Residual vector quantization (RVQ) compresses encoder outputs with factorized codes and $\ell_{2}$-normalized codebooks.
-For a 1 s audio $x(t)$ at $F_{s}{=}24$ kHz compressed to R=6000R{=}6000 bps, the encoder $Enc(\cdot)$ downsamples by M=320M{=}320 to $T{=}{F_{s}}/{M}{=}75$ frames $Z{=}[z_{t}\!\in\!\mathbb{R}^{d}]_{t=1}^{T}$ with r=R/T=80r{=}R/T{=}80 bits/frame.
+For a 1 s audio $x(t)$ at $F_{s}{=}24$ kHz compressed to R=6000R{=}6000 bps, the encoder $Enc(\cdot)$ downsamples by M=320M{=}320 to $T{=}{F_{s}}/{M}{=}75$ frames of latents $Z{=}[z_{t}\!\in\!\mathbb{R}^{d}]_{t=1}^{T}$ ($d$: channel width) with r=R/T=80r{=}R/T{=}80 bits/frame.
 RVQ allocates $r_{i}{=}r/N_{q}{=}10$ bits across $N_{q}{=}8$ codebooks (size $2^{10}{=}1024$).
@@
 Text-guided Sound Separation in Spectrogram-domain (AudioSep):
-$x(t)\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
-where $g(\cdot,e_{\tau})$ is a FiLM-conditioned masker that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$ given the complex STFT $X$ of audio $x(t)$ and text-embedding $e_{\tau}$.
+$x(t)\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,Spec(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
+where $Spec(\cdot,e_{\tau})$ is a FiLM-conditioned masker that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ ($F$: frequency bins, $T_{spec}$: spectrogram frames) and a phase residual $\angle\hat{M}_{s}$ given the complex STFT $X$ of audio $x(t)$ and text-embedding $e_{\tau}$.
 Text-guided Sound Separation in NAC latent-domain (CodecSep):
-$x(t)\xrightarrow{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
+$x(t)\xrightarrow[\text{DAC}]{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow[\text{DAC}]{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
 with frozen DAC backbone $Enc(\cdot),Dec(\cdot)$ and a FiLM-conditioned transformer masker $Mask(\cdot,e_{\tau})$ that estimates mask $M_{s}\!\in\![0,1]^{d\times T}$ applied element-wise to DAC latent $Z$.
@@
 Crucially, $Enc(\cdot)$ organizes $Z$ on a discriminative, perceptually aligned manifold , making selection (masking) easier than representation learning from raw $X$.
-In spectrogram systems, the separator must first learn a high-level latent from $X$ (via CNN/UNet) and then separate, coupling abstraction and masking and inflating parameters/compute.
+In spectrogram systems $Spec(.)$, the separator must first learn a high-level latent from $X$ (via CNN/UNet) and then separate, coupling abstraction and masking and inflating parameters/compute.
+Waveform separators $Wave(.)$ such as Sudo rm-rf Tzinis et al.
+(2022a, b, 2023) likewise downsample the waveform into STFT-like intermediate latents via 1D convolutions before encoding, resulting in latents with similar dimensionality and thus inheriting the same challenges as spectrogram systems.
 Masking over codec latents (leveraging the codec prior).
@@
 We exploit this organization with a FiLM-conditioned transformer masker that predicts a soft mask $M_{s}\in[0,1]^{d\times T}$ and applies it element-wise, yielding source latent estimate $\tilde{Z}_{s}=M_{s}\odot Z$.
-In contrast to learning a generator $Sep(Z,e_{\tau}):Z\!\to\!\tilde{Z}_{s}$ as in CodecFormer, learning a mask $Mask(Z,e_{\tau}):Z\!\to\!M_{s}$ on the compact, semantically organized codec manifold both exploits the codec prior more effectively and yields a more stable optimization that converges faster.
+In contrast to learning a generator $Gen(Z,e_{\tau}):Z\!\to\!\tilde{Z}_{s}$ as in CodecFormer, learning a mask $Mask(Z,e_{\tau}):Z\!\to\!M_{s}$ on the compact, semantically organized codec manifold both exploits the codec prior more effectively and yields a more stable optimization that converges faster.
 Moreover, masking in the denoised, low-dimensional codec space is fundamentally easier than masking in the high-dimensional, noisy spectrogram domain.
@@
 CodecFormer-EL) in place of, or alongside, SI-SDR: $\mathcal{L}_{\mathrm{emb}}=\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$.
-In deployment, the variant simply replaces the masker input with $E$ and optionally re-quantizes for codes-out.
-Deployment advantages over spectrogram-domain separators.
-In realistic pipelines, edge devices already run a codec and send code streams.
-Spectrogram models must decode $\to$ separate on $X$ $\to$ re-encode, adding latency and energy; CodecSep masks in the codec domain and can return separated code streams directly:
-$\text{Edge: }Enc(\cdot)\Rightarrow\text{Code}\Rightarrow\text{Server: reconstruct }E\!\approx\!Z\ \&\ Mask(\cdot,e_{\tau})\Rightarrow\text{Code}\Rightarrow\text{Edge: }Dec(\cdot).$
-Concretely, with codec costs $C_{\mathrm{enc}},C_{\mathrm{dec}}$, spectrogram separator cost $C_{\mathrm{spec}}$, and CodecSep masker cost $C_{\mathrm{mask}}$:
-$\text{Code-stream input: }\ \text{AudioSep}=C_{\mathrm{dec}}{+}C_{\mathrm{spec}}{+}C_{\mathrm{enc}},\quad\text{CodecSep}=C_{\mathrm{mask}}.$
-Within the separator, working on $Z/E$ (small $d,T$ with $|Z|\!\ll\!|X|$) shrinks attention/MLP activations, enabling tighter batching and lower memory bandwidth.
-Interface compatibility is immediate when only codes $A$ are available: perform lookup to $E$, apply the same FiLM-conditioned masker, and optionally re-quantize to emit $\hat{A}_{s}$.
-Overall, CodecSep (i) eliminates redundant decode/re-encode in server workflows; (ii) preserves the codec manifold (periodicity, timbre, transients) by modulating $Z/E$ instead of rebuilding $X$; and (iii) keeps conditioning non-iterative (single forward pass), yielding low-latency, high-fidelity stems at scale.
-See Appendix A for a full rationale covering latent-domain masking vs.
-spectrogram-domain generation, the RVQ-induced semantic prior, FiLM placement inside the masker, continuous $Z$ vs.
-discrete $A{\to}E$ operation (and codes-in/out), and complexity/deployment comparisons.
+In deployment, the variant simply replaces the masker input with $E$ and optionally re-quantizes for codes-out as,
+$x(t)\;\xrightarrow[\text{On Edge}]{\,Quant(Enc(\cdot))\,}\;A\;\xrightarrow[\text{Codes In}]{\,lookup(A)\,}\;E\approx Z\;\xrightarrow[\text{On Server}]{\,Mask(E,e_{\tau})\,}\;\tilde{E}_{s}=M_{s}\!\odot\!E\;\xrightarrow[\text{Codes Out}]{\,Quant(\tilde{E}_{s})\,}\;\hat{A}_{s}.$
+Deployment advantages over audio-stream separators.
+In realistic pipelines, edge devices already run a codec and transmit code streams rather than raw audio.
+Traditional spectrogram-based $Spec(.)$ and waveform-based $Wave(.)$ separators, however, operate on the audio stream: they first convert audio to STFT or STFT-like representations (often via 1D convolutions), and then must decode $\to$ separate on $X$ $\to$ re-encode, incurring additional latency and energy cost.
+In contrast, CodecSep performs masking directly in the codec domain and can output code streams without any decode–re-encode cycle.
+Concretely, with codec costs $C_{{Enc}},C_{{Dec}}$, spectrogram or audio-stream separator (AudioSep) cost $C_{{Spec}}$, and CodecSep masker cost $C_{{Mask}}$:
+$\text{Compute Cost for Code-stream input: }\ \text{AudioSep}=C_{{Dec}}{+}C_{{Spec}}{+}C_{{Enc}},\quad\text{CodecSep}=C_{{Mask}}.$
+We treat the codebook lookup $C_{{lookup}}$ and quantization $C_{{Quant}}$ costs as negligible ($\approx 0$) and omit the CLAP text-encoder cost since it is shared across all models.
+Figure 2 illustrates a typical edge–server deployment and compares compute requirements for conventional audio-stream separators (audio in $\to$ codes out) versus CodecSep’s code-stream separator (codes in $\to$ codes out).
+As shown, the CodecSep masker operates on $Z/E$ with small $(d,T)$ where $|Z|\ll|X|$, dramatically reducing attention and MLP activations and enabling tighter batching and lower memory bandwidth.
+Interface compatibility is immediate when only codes $A$ are available: perform a lookup to obtain $E$, apply the FiLM-conditioned masker, and optionally re-quantize to produce $\hat{A}_{s}$.
+CodecSep thus eliminates redundant decode/re-encode cycles in server workflows yielding low-latency, high-fidelity separation at scale.
+See Appendix B for a full discussion covering all of the aforementioned rationale in §3.1.
+Figure 2: Typical edge–server deployment comparing compute requirements of conventional audio-stream separators (audio in $\to$ codes out) versus CodecSep discrete inference (codes in $\to$ codes out).
 3.2 Our Model
@@
 (2022) from fixed-label three-stem separation to universal, prompt-driven separation by replacing source labels with natural-language queries.
-Speech and music are queried with broad category prompts (e.g., “speech”, “music”), while SFX stems are queried using long-form, compositional prompts (>1>1 overlapping sources) synthesized from FSD50K’s hierarchical annotations Fonseca et al.
+Speech and music are queried with broad category prompts (e.g., “speech”, “music”), while SFX stems are queried using long-form, compositional prompts ($\geq 2$ overlapping sources) synthesized from FSD50K’s hierarchical annotations Fonseca et al.
 (2022), combining fine-grained classes with parent categories (e.g., “dog barking, Animal, engine rumbling, motor vehicle”).
@@
 (2020a).
-Dataset construction details, clip durations, split statistics, and segmentation rules are provided in the Appendix B.
+Dataset construction details, clip durations, split statistics, and segmentation rules are provided in the Appendix C.
 Evaluation.
-We compare CodecSep against representative spectrogram- and codec-domain baselines—TDANet Li et al.
+We compare CodecSep against representative spectrogram-, waveform- and codec-domain baselines—TDANet Li et al.
 (2023); Pons et al.
-(2024), CodecFormer Yip et al.
+(2024), Sudo rm-rf Tzinis et al.
+(2022a), CodecFormer Yip et al.
 (2024b), SDCodec Bie et al.
-(2024), and the text-guided spectrogram model AudioSep Liu et al.
-(2024).
+(2024), and the text-guided audio stream separators AudioSep Liu et al.
+(2024), BiModalSS Mahmud et al.
+(2024) and Sudo rm-rf + FiLM Tzinis et al.
+(2022b, 2023).
 We report objective signal fidelity via scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al.
@@
 Following prior work (e.g., SDCodec Bie et al.
-(2024)), we use ViSQOL as a proxy for perceptual listening quality and complement it with a human MOS-LQS study comparing real-world outputs from CodecSep (trained on dnr-v2) and the publicly released AudioSep.
-To quantify efficiency in a hardware-agnostic manner, we report multiply–accumulate operations (MACs) using torchprofile [^1] https://github.com/zhijian-liu/torchprofile under matched input durations (2 s) and batching (batch size 2).
+(2024)), we use ViSQOL as a proxy MOS score for perceptual listening quality and complement it with a human MOS-LQS study comparing real-world outputs from CodecSep (trained on dnr-v2) and the publicly released AudioSep.
+To quantify efficiency, we report multiply–accumulate operations (MACs), inference time, memory footprint using torchinfo [^1] https://github.com/tyleryep/torchinfo under matched input durations (2 s) and batching (batch size 2) on a 2424GB NVIDIA A30 GPU.
+Details on evaluation workflow for each benchmark are deferred to Appendix D.
 Training.
@@
 We produce two variants of CodecSep, trained separately on dnr-v2 and on AudioCaps, to study distributional effects; we denote them with the suffixes “+dnr-v2” and “+AudioCaps”.
-For fair comparison, TDANet and CodecFormer are re-trained from scratch on dnr-v2 using our setup, AudioSep is evaluated both as a public checkpoint and when re-trained under matched protocols, and SDCodec is used as released by the authors.
+For fair comparison, the 3-stem versions of TDANet, Sudo rm-rf, and CodecFormer are re-trained from scratch on dnr-v2 using our setup.
+AudioSep is evaluated both as the publicly released checkpoint and when re-trained under matched protocols.
+We similarly re-train Sudo rm-rf+FiLM under the same matched settings.
+Pretrained checkpoints of BiModalSS and SDCodec are used as released by the authors.
 To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through a full-band stereo-capable $48$ kHz EnCodec during both training and inference.
-Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix C.
+Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix E.
 Table 1: Results: Separation Performance, Universal Sound Separation (dnr-v2-test)
 Model Metric ($\uparrow$) Music Speech Sfx
-AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
-(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
-AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
-ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
-ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
-CodecSep + dnr-v2 SI-SDR $0.16^{\pm 3.55}$ $8.26^{\pm 2.60}$ $0.96^{\pm 4.20}$
-(codes in : codes out, zero-shot) ViSQOL $2.52^{\pm 0.52}$ $2.97^{\pm 0.44}$ $2.33^{\pm 0.67}$
-CodecSep + dnr-v2 SI-SDR $6.72^{\pm 2.77}$ $1.95^{\pm 2.84}$ $6.75^{\pm 3.83}$
-(ablate Masker) ViSQOL $2.48^{\pm 0.58}$ $2.58^{\pm 0.50}$ $2.08^{\pm 0.74}$
+AudioSep SI-SDR $2.5^{\pm 4.06}$ $4.9^{\pm 4.21}$ $0.3^{\pm 5.39}$
+(zero-shot) ViSQOL $2.9^{\pm 0.63}$ $3.1^{\pm 0.56}$ $\mathbf{2.6^{\pm 0.77}}$
+Bimodal SI-SDR $6.8^{\pm 2.73}$ $1.8^{\pm 2.78}$ $6.36^{\pm 3.57}$
+(zero-shot) ViSQOL $2.5^{\pm 0.57}$ $2.6^{\pm 0.51}$ $2.3^{\pm 0.72}$
+AudioSep + dnr-v2 SI-SDR $5.6^{\pm 2.89}$ $7.7^{\pm 3.0}$ $4.7^{\pm 3.68}$
+ViSQOL $2.6^{\pm 0.57}$ $2.5^{\pm 0.37}$ $2.3^{\pm 0.7}$
+Sudo rm-rf + FiLM + dnr-v2 SI-SDR $6.7^{\pm 2.62}$ $2.0^{\pm 2.76}$ $6.6^{\pm 3.71}$
+ViSQOL $2.7^{\pm 0.59}$ $2.9^{\pm 0.45}$ $2.3^{\pm 0.72}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{1.2^{\pm 3.29}}$ $\mathbf{10.0^{\pm 2.92}}$ $\mathbf{0.9^{\pm 4.22}}$
+ViSQOL $\mathbf{2.9^{\pm 0.57}}$ $\mathbf{3.2^{\pm 0.45}}$ $2.3^{\pm 0.73}$
+CodecSep + dnr-v2 SI-SDR $0.2^{\pm 3.55}$ $8.3^{\pm 2.60}$ $1.0^{\pm 4.20}$
+(codes in : codes out, zero-shot) ViSQOL $2.5^{\pm 0.52}$ $3.0^{\pm 0.44}$ $2.3^{\pm 0.67}$
+CodecSep + dnr-v2 SI-SDR $6.8^{\pm 2.77}$ $2.0^{\pm 2.84}$ $6.8^{\pm 3.83}$
+(ablate Masker) ViSQOL $2.5^{\pm 0.58}$ $2.6^{\pm 0.50}$ $2.1^{\pm 0.74}$
 4.1 Results and Discussions
@@
 Table 1).
-CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot) and retrained AudioSep+dnr-v2 across all categories, with sizable SI-SDR gains in speech (9.97 vs.
-4.92/7.68 dB), music (1.15 vs.
-2.46-2.46/5.55-5.55 dB), and SFX (0.89 vs.
-0.34-0.34/4.66-4.66 dB).
+CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot) and retrained AudioSep+dnr-v2 across all categories, with sizable SI-SDR gains in speech ($10.0$ vs.
+4.94.9/$7.7$ dB), music (1.2 vs.
+2.5-2.5/5.6-5.6 dB), and SFX ($0.9$ vs.
+0.3-0.3/4.7-4.7 dB).
 In ViSQOL, CodecSep matches or exceeds AudioSep in speech and music while slightly trailing on SFX, likely reflecting differences in SFX prompt distributions (AudioSep’s diverse training vs.
@@
 Importantly, our bitstream-native variant—CodecSep+dnr-v2 (codes in: codes out, zero-shot)—evaluates the CodecSep+dnr-v2 masker directly on reconstructed embeddings $E$ from code streams (§3.1, para.
-6677) without any fine-tuning: it incurs a modest drop relative to the continuous-latent path (about $1{\text{–}}2$ dB SI-SDR across sources; small ViSQOL deltas for music/speech and parity on SFX), yet still surpasses AudioSep+dnr-v2 on SI-SDR for all sources (e.g., music: 0.16-0.16 vs.
-5.55-5.55 dB; speech: $8.26$ vs.
-$7.68$ dB; SFX: 0.96-0.96 vs.
-4.66-4.66 dB).
+6677) without any fine-tuning: it incurs a modest drop relative to the continuous-latent path (about $1{\text{–}}2$ dB SI-SDR across sources; small ViSQOL deltas for music/speech and parity on SFX), yet still surpasses AudioSep+dnr-v2 on SI-SDR for all sources (e.g., music: 0.2-0.2 vs.
+5.6-5.6 dB; speech: $8.3$ vs.
+$7.7$ dB; SFX: 1.0-1.0 vs.
+4.7-4.7 dB).
 Compared to pretrained AudioSep (zero-shot), the codes-in:codes-out variant improves SI-SDR on speech/music but lags on SFX SI-SDR and ViSQOL.
@@
 As part of our ablation, the lightweight CodecSep+dnr-v2 (ablate Masker) removes the transformer masker and applies FiLM directly to the encoder; it attains SI-SDR comparable to AudioSep+dnr-v2 with better perceptual speech quality, but overall separation quality drops due to FiLM perturbing the mixture latents held by the encoder.
+Beyond AudioSep, CodecSep also outperforms the USS-pretrained BiModalSS model and the retrained text-conditioned Sudo rm-rf + FiLM baseline.
+The heavier attention-based conditioning used in BiModalSS does not generalize well to the open-domain mixtures in dnr-v2, leading to degraded performance.
+CodecSep also outperforms the Sudo rm-rf + FiLM model trained under the same universal setting.
+This gap is likely due to two factors.
+First, continuous CLAP embeddings provide far richer semantic conditioning than the fixed one-hot or multi-hot vectors that Sudo rm-rf + FiLM was designed for, causing the model to struggle with open-domain prompts.
+Second, applying FiLM across all U-Conv blocks while using only a single Conv1d layer for audio encoding can destabilize internal representations, especially when conditioned with high-dimensional continuous embeddings.
+Since AudioSep also outperforms both by a large margin, we use AudioSep as the primary baseline in our further experiments.
 Effect of SFX Prompt Granularity during training.
-We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec) (cf.
+We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, Sudo rm-rf, CodecFormer, SDCodec) (cf.
 Table 2), (ii) generic 3-stem prompts (“music/speech/sfx”) (cf.
@@
 For (ii) and (iii), we train and evaluate separate versions of both CodecSep and AudioSep on each prompt setting.
-Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality (cf.
+Across settings, CodecSep matches or exceeds fixed-stem baselines on SFX while maintaining comparable speech and music quality, and its USS variant outperforms Sudo rm-rf across all stems with modest margins (cf.
 Table 2).
@@
 3-Stem: Fixed stem baselines (no text-guidance)
-TDANet SI-SDR $1.84^{\pm 3.55}$ $10.18^{\pm 2.91}$ $1.36^{\pm 4.90}$
-ViSQOL $2.85^{\pm 0.58}$ $3.09^{\pm 0.43}$ $2.44^{\pm 0.72}$
-CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
-ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
-SDCodec SI-SDR $\mathbf{1.85^{\pm 3.68}}$ $\mathbf{11.32^{\pm 2.98}}$ $\mathbf{1.77^{\pm 4.08}}$
-ViSQOL $\mathbf{2.96^{\pm 0.56}}$ $\mathbf{3.49^{\pm 0.40}}$ $\mathbf{2.64^{\pm 0.73}}$
+TDANet SI-SDR $1.8^{\pm 3.55}$ $10.2^{\pm 2.91}$ $1.4^{\pm 4.90}$
+ViSQOL $2.9^{\pm 0.58}$ $3.1^{\pm 0.43}$ $2.4^{\pm 0.72}$
+Sudo rm-rf SI-SDR $0.9^{\pm 4.01}$ $9.0^{\pm 2.60}$ $0.6^{\pm 4.74}$
+ViSQOL $2.7^{\pm 0.59}$ $2.9^{\pm 0.45}$ $2.3^{\pm 0.72}$
+CodecFormer SI-SDR $5.7^{\pm 3.44}$ $2.3^{\pm 2.32}$ $6.5^{\pm 4.36}$
+ViSQOL $2.2^{\pm 0.47}$ $2.5^{\pm 0.49}$ $2.1^{\pm 0.67}$
+SDCodec SI-SDR $\mathbf{1.9^{\pm 3.68}}$ $\mathbf{11.3^{\pm 2.98}}$ $\mathbf{1.8^{\pm 4.08}}$
+ViSQOL $\mathbf{3.0^{\pm 0.56}}$ $\mathbf{3.5^{\pm 0.40}}$ $\mathbf{2.6^{\pm 0.73}}$
 3-Stem: {“music”,“speech”,“sfx”} as generic prompt
-AudioSep SI-SDR $\mathbf{-2.46^{\pm 4.06}}$ $4.92^{\pm 4.21}$ $6.65^{\pm 4.73}$
-(zero-shot) ViSQOL $\mathbf{2.86^{\pm 0.63}}$ $\mathbf{3.11^{\pm 0.56}}$ $2.08^{\pm 0.68}$
-AudioSep + dnr-v2 SI-SDR $6.22^{\pm 2.77}$ $\mathbf{7.71^{\pm 3.11}}$ $2.11^{\pm 3.90}$
-ViSQOL $2.55^{\pm 0.57}$ $2.47^{\pm 0.37}$ $2.39^{\pm 0.74}$
-CodecSep + dnr-v2 SI-SDR $7.71^{\pm 2.84}$ $4.63^{\pm 2.48}$ $\mathbf{0.58^{\pm 4.15}}$
-ViSQOL $2.45^{\pm 0.55}$ $2.70^{\pm 0.49}$ $\mathbf{2.39^{\pm 0.70}}$
+AudioSep SI-SDR $\mathbf{-2.5^{\pm 4.06}}$ $4.9^{\pm 4.21}$ $6.7^{\pm 4.73}$
+(zero-shot) ViSQOL $\mathbf{2.9^{\pm 0.63}}$ $\mathbf{3.1^{\pm 0.56}}$ $2.1^{\pm 0.68}$
+AudioSep + dnr-v2 SI-SDR $6.2^{\pm 2.77}$ $\mathbf{7.7^{\pm 3.11}}$ $2.1^{\pm 3.90}$
+ViSQOL $2.6^{\pm 0.57}$ $2.5^{\pm 0.37}$ $2.4^{\pm 0.74}$
+CodecSep + dnr-v2 SI-SDR $7.7^{\pm 2.84}$ $4.6^{\pm 2.48}$ $\mathbf{0.6^{\pm 4.15}}$
+ViSQOL $2.5^{\pm 0.55}$ $2.7^{\pm 0.49}$ $\mathbf{2.4^{\pm 0.70}}$
 Why use a Transformer masker instead of a decoder?(cf.
@@
 Model Metric ($\uparrow$) Music Speech Sfx
-CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
-ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
-CodecSep + dnr-v2 SI-SDR $1.15^{\pm 3.35}$ $9.90^{\pm 2.91}$ $0.82^{\pm 4.18}$
-(unguided, 3-stem) ViSQOL $2.75^{\pm 0.55}$ $3.09^{\pm 0.45}$ $\mathbf{2.48^{\pm 0.72}}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
-(text-guided) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
+CodecFormer SI-SDR $5.8^{\pm 3.44}$ $2.3^{\pm 2.32}$ $6.5^{\pm 4.36}$
+ViSQOL $2.2^{\pm 0.47}$ $2.5^{\pm 0.49}$ $2.1^{\pm 0.67}$
+CodecSep + dnr-v2 SI-SDR $1.2^{\pm 3.35}$ $10.0^{\pm 2.91}$ $0.9^{\pm 4.18}$
+(unguided, 3-stem) ViSQOL $2.8^{\pm 0.55}$ $3.1^{\pm 0.45}$ $\mathbf{2.5^{\pm 0.72}}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{1.2^{\pm 3.29}}$ $\mathbf{10.0^{\pm 2.92}}$ $\mathbf{0.9^{\pm 4.22}}$
+(text-guided) ViSQOL $\mathbf{2.9^{\pm 0.57}}$ $\mathbf{3.2^{\pm 0.45}}$ $2.3^{\pm 0.73}$
 Further benchmarking on ESC-50, Clotho-v2, AudioSet, VGGSound, & AudioCaps (cf. Table 4).
@@
 Under matched training data and prompt protocols, CodecSep+dnr-v2 consistently outperforms AudioSep+dnr-v2 across all five datasets in both separation fidelity and perceptual quality.
+[…]
+Additional experiments (cf. Appendix FJ) .
+For readability, five extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs.
+universal and paraphrased variants), training/evaluation splits, and metrics.
+(i) Robustness to prompt paraphrasing (cf.
+Appendix F): We test lexical robustness by replacing the generic speech and music prompts with three unseen paraphrases each (e.g., “spoken voice,” “people talking”; “instrumental music,” “band playing”).
+This zero-shot paraphrase test requires models trained on generic labels to handle synonymic descriptors at inference.
+While both CodecSep+dnr-v2 and AudioSep+dnr-v2 degrade under paraphrasing, CodecSep remains consistently stronger (Table 6).
+This experiment focuses on lexical variants; paraphrases involving temporal structure (e.g., “applause follows a song”) are left for future work.
+(ii) Generalization across open-domain datasets (cf.
+Appendix G): On AudioCaps (derived from AudioSet), the spectrogram baseline AudioSep benefits from distributional alignment and attains the strongest absolute scores, yet CodecSep+dnr-v2 generalizes competitively—surpassing AudioSep+dnr-v2 in SI-SDR at comparable ViSQOL (cf.
+Table 7).
+When both are retrained on AudioCaps, CodecSep+AudioCaps again outperforms AudioSep+AudioCaps in separation quality.
+The same trend holds on the more challenging dnr-v2 test set, where mixtures often contain speech, music, and multiple overlapping SFX; both models drop in absolute performance, but CodecSep retains its advantage in SI-SDR.
+(iii) Relative–gain summaries (cf.
+Appendix H): We report percent improvements of CodecSep over AudioSep under matched training data and prompt settings; CodecSep shows large SI-SDR gains on dnr-v2 with positive (though smaller) gains under paraphrased prompts, and consistent improvements across ESC-50, Clotho-v2, AudioSet, VGGSound, and AudioCaps, with modest ViSQOL deltas (cf.
+Table 8).
+(iv) Higher bandwidth (cf.
+Appendix I): Replacing the frozen 16 kHz DAC with a 48 kHz, stereo-capable EnCodec leaves the masking architecture unchanged but increases latent-frame count and high-frequency detail, making separation harder and reducing SI-SDR/ViSQOL (Table 9).
+At higher $F_{s}$, gains depend mainly on backbone capacity and data rather than architectural modifications.
+Evaluations of 24 kHz and 44.1 kHz DAC variants show the same pattern: with fixed masker capacity, performance declines monotonically as sampling rate increases, reflecting the difficulty of wide-bandwidth separation, while DAC consistently outperforms EnCodec at matched bandwidths.
+(v) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided).
+We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original.
+This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix J (cf.
+Table 10).
 Table 4: Results: Benchmarking on ESC-50, Clotho-v2, AudioSet, VGGSound, AudioCaps
 Model Metric ($\uparrow$) ESC-50 Clotho-v2 AudioSet VGGSound AudioCaps
-AudioSep SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$ $6.44^{\pm 11.48}$
-+ dnr-v2 ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$ $\mathbf{2.33^{\pm 1.08}}$
-CodecSep SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$ $\mathbf{-6.09^{\pm 11.62}}$
-+ dnr-v2 ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$ $2.24^{\pm 1.16}$
-Robustness to prompt paraphrasing (cf.
-Table 5) .
-To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}.
-This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference.
-We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts.
-Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
-Full inference complexity (cf.
-Table 6) .
-We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep).
-Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep.
-Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space.
-Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting.
-Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth.
-Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
-Additional experiments (cf. Appendix DG) .
-For readability, three extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs.
-universal and paraphrased variants), training/evaluation splits, and metrics.
-(i) Generalization across open-domain datasets (cf.
-Appendix D): On AudioCaps (derived from AudioSet), the spectrogram baseline AudioSep benefits from distributional alignment and attains the strongest absolute scores, yet CodecSep+dnr-v2 generalizes competitively—surpassing AudioSep+dnr-v2 in SI-SDR at comparable ViSQOL.
-When both are retrained on AudioCaps, CodecSep+AudioCaps again outperforms AudioSep+AudioCaps in separation quality.
-The same trend holds on the more challenging dnr-v2 test set, where mixtures often contain speech, music, and multiple overlapping SFX; both models drop in absolute performance, but CodecSep retains its advantage in SI-SDR.
-(ii) Relative–gain summaries (cf.
-Appendix E): We report percent improvements of CodecSep over AudioSep under matched training data and prompt settings; CodecSep shows large SI-SDR gains on dnr-v2 with positive (though smaller) gains under paraphrased prompts, and consistent improvements across ESC-50, Clotho-v2, AudioSet, VGGSound, and AudioCaps, with modest ViSQOL deltas (cf.
-Table 8).
-(iii) Higher bandwidth (cf.
-Appendix F): Swapping the frozen 16 kHz DAC for a 48 kHz (stereo–capable) EnCodec preserves the masking architecture and deployment path.
-Higher $F_{s}$ lengthens $T$ and adds high-frequency detail, making separation harder and lowering absolute SI-SDR/ViSQOL (cf.
-Table 9).
-Improvements hinge on backbone capacity/data—not on redesign.
-(iv) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided).
-We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original.
-This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix G (cf.
-Table 10).
-Table 5: Results: Using ambiguous prompts for Speech and Music (dnr-v2-test)
-Model Metric ($\uparrow$) Music Speech
-AudioSep + dnr-v2 SI-SDR $6.43^{\pm 3.29}$ $4.14^{\pm 3.77}$
-ViSQOL $2.53^{\pm 0.57}$ $2.64^{\pm 0.47}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{-5.60^{\pm 3.61}}$ $\mathbf{4.19^{\pm 4.18}}$
-ViSQOL $\mathbf{2.56^{\pm 0.58}}$ $\mathbf{2.74^{\pm 0.51}}$
-Table 6: Full Inference Complexity (GMACs $\downarrow$)
-Model Input: Audio Stream Input: Code Stream Architecture-only
+AudioSep SI-SDR $7.8^{\pm 14.46}$ $8.6^{\pm 17.0}$ $7.6^{\pm 11.42}$ $7.0^{\pm 12.65}$ $6.4^{\pm 11.48}$
+[…]
+CodecSep SI-SDR $\mathbf{-5.9^{\pm 11.55}}$ $\mathbf{-6.0^{\pm 11.10}}$ $\mathbf{-6.4^{\pm 10.53}}$ $\mathbf{-6.1^{\pm 12.12}}$ $\mathbf{-6.1^{\pm 11.62}}$
++ dnr-v2 ViSQOL $\mathbf{2.3^{\pm 1.13}}$ $\mathbf{2.3^{\pm 1.09}}$ $\mathbf{2.2^{\pm 1.0}}$ $\mathbf{2.3^{\pm 1.11}}$ $2.2^{\pm 1.16}$
+Table 5: Full Inference Compute Benchmarking Across Six Settings on 2424GB NVIDIA A-30 GPU
+(a) Inference GMACs ($\downarrow$)
+Model Audio Stream I/O Code Stream I/O Architecture-only
 AudioSep $33.5$ $73.6$ $33.5$
-CodecSep $41.45$ $\mathbf{1.35}$ ($\downarrow 54\times$) $\mathbf{1.35}$ ($\downarrow 25\times$)
-Codec GMACs used above: DAC Enc(.) =12.28=12.28, Dec(.) =27.82=27.82.
+Sudo rm-rf $\mathbf{16.44}$ $56.54$ $16.44$
+CodecSep $41.45$ $\mathbf{1.35}$ $\mathbf{1.35}$
+Codec GMACs: Enc=$12.28$, Dec=$27.82$
+(b) Inference Time (s) ($\downarrow$)
+Model Audio Stream I/O Code Stream I/O Architecture-only
+AudioSep $\mathbf{0.33}$ $1.63$ $0.33$
+Sudo rm-rf $0.51$ $1.81$ $0.51$
+CodecSep $1.49$ $\mathbf{0.19}$ $\mathbf{0.19}$
+Codec Inference Times (s): Enc=$1.16$, Dec=$0.14$
+(c) Parameter count (in million) ($\downarrow$)
+Model Audio Stream I/O Code Stream I/O Architecture-only
+AudioSep $39$ $112.8$ $39$
+Sudo rm-rf $\mathbf{5}$ $78.8$ $\mathbf{5}$
+CodecSep $90.1$ $\mathbf{16.3}$ $16.3$
+Codec Parameter Count (in million): Enc=$21.5$, Dec=$52.3$
+(d) Parameter-only Memory Footprint (MB) ($\downarrow$)
+Model Audio Stream I/O Code Stream I/O Architecture-only
+AudioSep $156.13$ $447.62$ $156.13$
+Sudo rm-rf $\mathbf{20.07}$ $311.25$ $\mathbf{20.07}$
+CodecSep $339.89$ $\mathbf{48.4}$ $48.4$
+Codec Parameter-only Memory Footprint (MB): Enc=$85.1$, Dec=$206.39$
+(e) Forward/Backward Pass Memory Footprint (MB) ($\downarrow$)
+Model Audio Stream I/O Code Stream I/O Architecture-only
+AudioSep $\mathbf{804.3}$ $1580.6$ $804.3$
+Sudo rm-rf $2032.13$ $2836.43$ $2032.13$
+CodecSep $804.4$ $\mathbf{28.06}$ $\mathbf{28.06}$
+Codec Forward/Backward Pass Memory Footprint (MB): Enc=$310.48$, Dec=$465.82$
+(f) Full Memory Footprint (MB) ($\downarrow$)
+Model Audio Stream I/O Code Stream I/O Architecture-only
+AudioSep $\mathbf{960.43}$ $2028.22$ $960.43$
+Sudo rm-rf $2052.2$ $3147.68$ $2052.2$
+CodecSep $1144.25$ $\mathbf{76.46}$ $\mathbf{76.46}$
+Codec Full Memory Footprint (MB): Enc=$395.58.$, Dec=$672.21$
 Subjective evaluation (MOS–LQS).
@@
 Future Work.
-We will (i) extend separation to higher bandwidths and stereo/spatial audio (e.g., 48 kHz stereo Encodec; HO–DirAC Hold et al.
+We plan to (i) extend separation to higher bandwidths and stereo/spatial audio (e.g., 48 kHz stereo Encodec; HO–DirAC Hold et al.
 (2024), SpatialCodec Xu et al.
-(2024)), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
-Appendix A Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
+(2024)); (ii) broaden prompt coverage to include temporal, relational, and referring-expression cues; (iii) explore stronger separator objectives; (iv) extend conditioning beyond text to audio, image, and mixed prompts through the same FiLM interface; and (v) pursue interpretability analyses.
+In particular, the predicted masks and FiLM responses indicate how CodecSep selectively activates latent components, suggesting a clear direction for future work involving mask visualization and reconstruction-consistency checks.
+Appendix A Failure Modes of PIT and MixIT for Universal Sound Separation
+Permutation-Invariant Training (PIT) and Mixture-Invariant Training (MixIT) have historically been effective for closed-domain separation tasks where the number of underlying sources is known, fixed, or varies within a narrow, well-defined range.
+However, their underlying assumptions lead to structural limitations when applied to open-domain universal source separation (USS), where mixtures may contain an arbitrary and potentially large number of heterogeneous sound events.
+In this section, we summarize the key failure modes observed when training with PIT/MixIT to extend fixed-stem models to open-domain mixtures.
+A core limitation of PIT/MixIT is the requirement to specify a maximum number of output sources, denoted by $N$.
+During training, the model produces exactly $N$ outputs for every mixture, and the PIT or MixIT objective establishes a correspondence between these outputs and the underlying reference sources (or intermediate MixIT partitions).
+This design is brittle in scenarios where the true number of sources varies widely.
+When the mixture contains more than $N$ sources—a frequent occurrence in open-domain audio—the model has no mechanism to create additional outputs.
+Instead, it suffers source collision and collapses multiple sources into a single output stem, resulting in unavoidable leakage, loss of fine structure, and a sharp degradation in separation quality.
+The post-hoc identification step cannot recover the missing sources, because the model never produced separate representations for them in the first place; those sources simply do not exist within the model’s output space.
+Conversely, when the mixture contains fewer than $N$ sources, the model is still obligated to return $N$ outputs.
+This mismatch introduces new problems: several outputs correspond to no actual source and become “inactive” stems, while others may capture residual background energy or hallucinated content.
+These false positives degrade metrics such as SI-SDR and create ambiguity during evaluation because the model does not encode which stems are meant to be meaningful.
+Such outputs also make deployment difficult, as downstream systems must decide which stems to trust and which to ignore.
+The reliance on a fixed maximum number of sources $N$ also places a heavy burden on both training stability and computational cost.
+As $N$ increases, the permutation space in PIT expands combinatorially, and MixIT assignments become increasingly complex, making training slow, unstable, and in many cases prone to divergence.
+In open-domain datasets such as dnr-v2, mixtures may contain eight or more concurrent sound events, forcing PIT/MixIT baselines to adopt impractically large values of $N$ to avoid source collisions.
+In practice, such configurations are computationally prohibitive and empirically unreliable.
+These limitations collectively illustrate why PIT and MixIT, despite their historical success in speech separation and other closed-set tasks, are poorly suited for open-domain universal separation.
+Their fixed-output architecture is fundamentally mismatched to real-world mixtures that contain highly variable and unpredictable numbers of sources.
+In contrast, CodecSep bypasses this bottleneck entirely through free-form text-guided inference: the model extracts only the requested source category, emits no unused stems, and scales naturally to mixtures with arbitrary levels of overlap.
+This flexibility enables CodecSep to support both closed-set and open-domain use cases, while also providing a foundation for future extensions to fine-grained extraction of individual speaker stems, instrument stems, or sound-effect stems.
+Appendix B Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
 Problem setup and pipeline contrast.
@@
 Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs.
-A.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
+B.1 Why NAC latents (vs. spectrograms) and how the codec prior enables separation?
 STFT vs.
@@
 In contrast, spectrogram systems must learn a task-specific latent from $X$ and perform separation jointly, which increases parameter count and MACs, slows convergence, and can lead to overfitting in the absence of the codec’s inductive bias.
-A.2 Transformer-based masker vs. decoder-style generation (specific design choice).
+B.2 Transformer-based masker vs. decoder-style generation (specific design choice).
 In CodecSep, we replace decoder-style source generation (as in CodecFormer, which directly predicts the target waveform/spectrogram) with a Transformer-based, FiLM-conditioned masker that outputs a soft mask over the codec latents.
@@
 FiLM conditioning steers the transformer to decide where/how much information to pass—not what to generate—so parameters and compute in the $Q/K/V$ and MLP projections are concentrated on selection and attenuation.
-A.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
+B.3 Why We Integrated FiLM Conditioning into NAC-Based Separation
 We deliberately integrate Feature-wise Linear Modulation (FiLM) Perez et al.
@@
 Empirically this reduces training variance and mitigates long-horizon inconsistencies common with generator-style heads.
-A.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
+B.4 Why report results on $Z$ (continuous latents) and how to extend to code streams $A\!\to\!E$
 Why we evaluate on continuous latents $Z$.
@@
 The deployment-ready path to discrete code streams is immediate: switch the masker input to $E$ via lookup from $A$ and train/fine-tune accordingly, with optional distribution alignment and bitrate-robustness augmentation.
-A.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
+B.5 Deployment advantages of CodecSep over spectrogram-domain separators (AudioSep)
 Central motivation.
@@
 Sec.
-A), reducing activation memory and bandwidth throughout attention and MLP blocks.
+B), reducing activation memory and bandwidth throughout attention and MLP blocks.
 Interface compatibility with codec bitstreams.
@@
   • Eliminates redundant codec cycles in server workflows: With code streams, we reconstruct embeddings $E\approx Z$ by codebook lookup (Sec.
-A.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
+B.4), avoiding server-side decode/re-encode; only the edge decodes the final stems.
   • Smaller working representation during separation: masking in $Z/E$ (e.g., d=64d{=}64) reduces intermediate activations, lowering memory bandwidth and enabling tighter batching.
@@
   • Maintains codec manifold structure: by modulating $Z/E$ rather than rebuilding $X$, CodecSep preserves periodicity/timbre/transients already organized by the codec, supporting stable, high-fidelity stems at deployment.
-Appendix B Dataset Details
-B.1 Divide and Remaster v 2.0 (dnr-v2)
+Appendix C Dataset Details
+C.1 Divide and Remaster v 2.0 (dnr-v2)
 dnr-v2 Petermann et al.
@@
 This results in long-form, compositional queries that reflect the structure of the mixture (e.g., “dog barking, Animal, engine rumbling, motor vehicle”).
-B.2 Open-Domain Benchmarks
+C.2 Open-Domain Benchmarks
 We benchmark on five open-domain datasets spanning captioned audio, environmental sounds, and large multi-event corpora: AudioCaps Kim et al.
@@
 For AudioCaps we form both training and testing mixtures (same scale of test data as dnr-v2) by summing three clips (validation segmented into $5$ s, test preserves clips up to $20$ s), while for ESC-50, Clotho-v2, AudioSet-eval, and VGGSound we construct test-only mixtures using the same three-clip protocol.
-Appendix C Training Details
+Appendix D Evaluation Details
+Figure 3: Evaluation workflow for dnr-v2.
+Each mixture contains multi-source stems: speech (often multi-speaker), music (multi-instrument), and SFX ($\geq 3$ overlapping events).
+Fixed-stem baselines predict a fixed set of outputs (e.g., 3 stems), whereas CodecSep and other text-guided models generate only the prompted source.
+Speech and music are evaluated using generic prompts, while SFX uses long-form compositional prompts listing all SFX events in each mixture.
+Extracted signals are compared with ground-truth category stems using SI-SDR and ViSQOL.
+D.1 Divide and Remaster v 2.0 (dnr-v2)
+The dnr-v2 benchmark presents a challenging open-domain separation setting: although the dataset provides three category labels—speech, music, and sound effects—each category represents a multi-source stem.
+A single mixture frequently contains five to ten underlying acoustic sources, including overlapping speakers, multiple musical instruments, and several sound-effect events occurring either concurrently or in sequence.
+The three reported stems are therefore semantic groupings that support interpretability and reproducibility, rather than an indication that the mixture contains only three sources.
+Any evaluation methodology must respect this structure.
+Figure 3 provides an overview of our dnr-v2 evaluation workflow and highlights how these multi-source stems are handled across fixed-stem un-guided and text-guided models.
+For fixed-stem unguided architectures, evaluation is performed by mapping each predicted output stem to one of the three ground-truth stems and computing SI-SDR and ViSQOL on a per-category basis.
+Importantly, we do not employ PIT or MixIT training objectives for these baselines; instead, we train dedicated three-stem models that directly predict speech, music, and SFX stems.
+Text-guided models, including CodecSep, follow a fundamentally different inference and evaluation paradigm.
+Speech and music stems are recovered using generic prompts (“speech,” “music”), which reliably capture their multi-source content.
+In contrast, SFX stems require mixture-specific prompts because sound effects span a wide and open-domain label space.
+For each mixture, we use a long-form compositional prompt enumerating all SFX events present in the ground truth.
+This ensures that the model has sufficient semantic context to extract the full SFX stem.
+The number of sfx events present in a mixture varies considerably.
+We additionally perform an ambiguous-prompt evaluation, where deliberately underspecified prompts for speech and music are used to assess robustness to vague or incomplete semantic queries.
+After inference, the extracted waveform for each category is directly compared with the corresponding ground-truth stem using SI-SDR and ViSQOL.
+This evaluation design ensures fairness between fixed-stem and text-guided systems while faithfully reflecting the multi-source structure of dnr-v2 mixtures
+Figure 4: Evaluation workflow for the standardized three-source benchmarks (AudioCaps, ESC-50, Clotho, VGGSound, and AudioSet-eval).
+Following prior USS protocols, each mixture is constructed by combining three isolated events drawn from distinct classes.
+For each class, the corresponding textual prompt is supplied to the separator (e.g., “dog barking,” “gun shot,” “motor vehicle”), and the extracted signal is compared with the ground-truth isolated source using SI-SDR and ViSQOL.
+D.2 Open-Domain Benchmarks
+For AudioCaps, ESC-50, Clotho, VGGSound, and AudioSet-eval, we adopt the standardized three-source mixture protocol.
+Following established practice, each mixture is constructed by combining three isolated events drawn from different classes.
+Each source is then extracted by the text-guided models using its corresponding textual caption as prompt, and evaluation metrics are computed against the ground-truth isolated audio.
+Figure 4 illustrates the evaluation workflow used for these benchmarks, highlighting how class-specific prompts are applied and how the resulting predictions are matched against the ground-truth isolated sources.
+Although these datasets do not reflect the complex multi-source structure of dnr-v2, the standardized 3-way protocol enables direct benchmarking against prior work (AudioSep) under consistent conditions.
+Appendix E Training Details
 The complete model, including the query module $query(.)$, is trained for $400K$ iterations with DAC Kumar et al.
@@
 We refer to these models using the suffixes +dnr-v2 and +AudioCaps, respectively, to indicate which dataset each model was trained on.
-Since TDANet and CodecFormer were originally developed for speech separation, we re-trained newly initialized versions on the dnr-v2 training set using the same configuration as CodecSep.
-For AudioSep, we evaluate the publicly available pre-trained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps using our setup for consistency.
-We also include SDCodec, using the official pre-trained models made available by the authors.
-To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing with a full-band stereo-capable $48$ kHz EnCodec during training and inference.
+Since TDANet and CodecFormer were originally designed for speech separation, we re-train newly initialized 3-stem versions on the dnr-v2 training set using the same configuration as CodecSep.
+We also train a 3-stem Sudo rm-rf model to compare against compute-efficient separators.
+For AudioSep, we evaluate both the publicly available pretrained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps for consistency.
+We include SDCodec using the official pretrained checkpoints released by the authors.
+Finally, we incorporate the USS-pretrained variant of BiModalSS and re-train SudoRmRf+FiLM on dnr-v2 with CLAP text conditioning.
+To ensure a fair comparison, all inputs to TDANet, Sudo rm-rf, AudioSep, BiModalSS, and Sudo rm-rf+FiLM undergo codec processing with a full-band stereo-capable $48$ kHz EnCodec during training and inference.
 This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
-Appendix D Cross-Benchmark Performance: AudioCaps and dnr-v2.
+Appendix F Robustness to prompt paraphrasing.
+To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}.
+This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference.
+We observe a consistent qualitative pattern (cf.
+Table 6): (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts.
+Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
+Table 6: Results: Using ambiguous prompts for Speech and Music (dnr-v2-test)
+Model Metric ($\uparrow$) Music Speech
+AudioSep + dnr-v2 SI-SDR $6.4^{\pm 3.29}$ $4.1^{\pm 3.77}$
+ViSQOL $2.5^{\pm 0.57}$ $2.6^{\pm 0.47}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{-5.6^{\pm 3.61}}$ $\mathbf{4.2^{\pm 4.18}}$
+ViSQOL $\mathbf{2.6^{\pm 0.58}}$ $\mathbf{2.7^{\pm 0.51}}$
+Appendix G Cross-Benchmark Performance: AudioCaps and dnr-v2.
 Table 7(a) reports generalization results on the AudioCaps-test set derived from AudioSet.
@@
 SI-SDR ($\uparrow$) ViSQOL ($\uparrow$)
-AudioSep $\mathbf{-2.51^{\pm 12.14}}$ $\mathbf{2.44^{\pm 1.08}}$
-AudioSep + dnr-v2 (zero-shot) $6.44^{\pm 11.48}$ $2.33^{\pm 1.08}$
-CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.09^{\pm 11.62}}$ $2.24^{\pm 1.16}$
-AudioSep + AudioCaps $9.17^{\pm 18.71}$ $2.29^{\pm 1.11}$
-CodecSep + AudioCaps $\mathbf{-6.19^{\pm 10.58}}$ $2.14^{\pm 1.00}$
+AudioSep $\mathbf{-2.5^{\pm 12.14}}$ $\mathbf{2.4^{\pm 1.08}}$
+AudioSep + dnr-v2 (zero-shot) $6.4^{\pm 11.48}$ $2.3^{\pm 1.08}$
+CodecSep + dnr-v2 (zero-shot) $\mathbf{-6.1^{\pm 11.62}}$ $2.2^{\pm 1.16}$
+AudioSep + AudioCaps $9.2^{\pm 18.71}$ $2.3^{\pm 1.11}$
+CodecSep + AudioCaps $\mathbf{-6.2^{\pm 10.58}}$ $2.1^{\pm 1.00}$
 (b) Transfer to dnr-v2-test when trained on AudioCaps (zero-shot on dnr-v2).
 Model Metric ($\uparrow$) Music Speech Sfx
-AudioSep + AudioCaps SI-SDR $14.86^{\pm 23.08}$ $7.11^{\pm 25.80}$ $14.63^{\pm 23.26}$
-(zero-shot) ViSQOL $2.36^{\pm 0.71}$ $2.43^{\pm 0.70}$ $2.15^{\pm 0.79}$
-CodecSep + AudioCaps SI-SDR $8.46^{\pm 2.78}$ $2.47^{\pm 2.91}$ $5.93^{\pm 4.33}$
-(zero-shot) ViSQOL $2.27^{\pm 0.53}$ $2.56^{\pm 0.47}$ $2.06^{\pm 0.72}$
-Appendix E Discussion of Relative–Gain Summaries.
+AudioSep + AudioCaps SI-SDR $14.9^{\pm 23.08}$ $7.1^{\pm 25.80}$ $14.6^{\pm 23.26}$
+(zero-shot) ViSQOL $2.4^{\pm 0.71}$ $4^{\pm 0.70} 2.2$$2^{\pm 0.79}$
+CodecSep + AudioCaps SI-SDR $8.5^{\pm 2.78}$ $2.5^{\pm 2.91}$ $5.9^{\pm 4.33}$
+(zero-shot) ViSQOL $2.3^{\pm 0.53}$ $2.6^{\pm 0.47}$ $2.1^{\pm 0.72}$
+Appendix H Discussion of Relative–Gain Summaries.
 Tables 8(a)8(d) consolidate relative improvements of CodecSep over AudioSep under matched training data and prompt protocols, complementing the absolute results in the main text.
@@
 ViSQOL 6.5-6.5 3.8-3.8 $\mathbf{+5.4}$ -4.2
-Appendix F Bandwidth scaling: extending CodecSep to 48 kHz full-band
+Appendix I Bandwidth scaling: extending CodecSep to full-band audio
 […]
@@
 Model Sampling Rate Metric ($\uparrow$) Music Speech Sfx
-AudioSep $32$ kHz SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
-(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
-AudioSep + dnr-v2 $32$ kHz SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
-ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
-CodecSep + dnr-v2 $16$ kHz SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
-(DAC Backbone) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
-CodecSep + dnr-v2 $48$ kHz SI-SDR $2.8^{\pm 3.5}$ $5.36^{\pm 2.36}$ $0.46^{\pm 3.83}$
-(EnCodec Backbone) ViSQOL $2.39^{\pm 0.5}$ $2.63^{\pm 0.41}$ $2.37^{\pm 0.65}$
-Appendix G Further Studies: Reconstruction Performance.
+AudioSep $32$ kHz SI-SDR $2.5^{\pm 4.06}$ $4.9^{\pm 4.21}$ $0.3^{\pm 5.39}$
+(zero-shot) ViSQOL $2.9^{\pm 0.63}$ $3.1^{\pm 0.56}$ $\mathbf{2.6^{\pm 0.77}}$
+AudioSep + dnr-v2 $32$ kHz SI-SDR $5.6^{\pm 2.89}$ $7.7^{\pm 3.0}$ $4.5^{\pm 3.68}$
+ViSQOL $2.6^{\pm 0.57}$ $2.5^{\pm 0.37}$ $2.3^{\pm 0.7}$
+CodecSep + dnr-v2 $16$ kHz SI-SDR $\mathbf{1.2^{\pm 3.29}}$ $\mathbf{10.0^{\pm 2.92}}$ $\mathbf{0.9^{\pm 4.22}}$
+(DAC Backbone) ViSQOL $\mathbf{2.9^{\pm 0.57}}$ $\mathbf{3.1^{\pm 0.45}}$ $2.3^{\pm 0.73}$
+CodecSep + dnr-v2 $24$ kHz SI-SDR $0.2^{\pm 3.3}$ $8.8^{\pm 2.9}$ $0.6^{\pm 4.2}$
+(DAC Backbone) ViSQOL $2.7^{\pm 0.56}$ $3.0^{\pm 0.44}$ $2.3^{\pm 0.72}$
+CodecSep + dnr-v2 $44.1$ kHz SI-SDR $2.3^{\pm 3.27}$ $5.9^{\pm 2.48}$ $0.3^{\pm 3.79}$
+(DAC Backbone) ViSQOL $2.5^{\pm 0.46}$ $2.7^{\pm 0.42}$ $2.4^{\pm 0.68}$
+CodecSep + dnr-v2 $48$ kHz SI-SDR $2.8^{\pm 3.5}$ $5.4^{\pm 2.36}$ $0.5^{\pm 3.83}$
+(EnCodec Backbone) ViSQOL $2.4^{\pm 0.5}$ $2.6^{\pm 0.41}$ $2.4^{\pm 0.65}$
+Appendix J Further Studies: Reconstruction Performance.
 Table 10 assesses performance under a single-source reconstruction setting on the dnr-v2-test set, where each model is prompted to reproduce the input source.
@@
 3-Stem: Fixed Stem, Non Text-guided
-TDANet SI-SDR $3.26^{\pm 7.78}$ $\mathbf{7.96^{\pm 5.29}}$ $\mathbf{11.07^{\pm 3.32}}$ $\mathbf{4.70^{\pm 5.11}}$
-ViSQOL $3.93^{\pm 0.35}$ $\mathbf{4.20^{\pm 0.46}}$ $\mathbf{4.51^{\pm 0.32}}$ $\mathbf{4.09^{\pm 0.39}}$
-CodecFormer SI-SDR $47.55^{\pm 9.51}$ $47.11^{\pm 10.97}$ $47.77^{\pm 9.65}$ $48.21^{\pm 9.87}$
-ViSQOL $1.02^{\pm 0.07}$ $4^{\pm 0.12} 1.01±0$$^{\pm 0.06} 1.1$$16^{\pm 0.47}$
-CodecSep + dnr-v2 SI-SDR $3.42^{\pm 1.85}$ $4.12^{\pm 3.97}$ $6.15^{\pm 2.87}$ $0.83^{\pm 5.16}$
-(unguided, 3-stem) ViSQOL $3.23^{\pm 0.20}$ $3.02^{\pm 0.33}$ $3.47^{\pm 0.24}$ $3.23^{\pm 0.46}$
-SDCodec SI-SDR $\mathbf{6.98^{\pm 2.49}}$ $7.65^{\pm 4.60}$ $8.28^{\pm 3.26}$ $2.54^{\pm 5.65}$
-ViSQOL $\mathbf{4.29^{\pm 0.15}}$ $03^{\pm 0.28} 4.4$$44^{\pm 0.15}$ $3.98^{\pm 0.34}$
+TDANet SI-SDR $3.3^{\pm 7.78}$ $\mathbf{8.0^{\pm 5.29}}$ $\mathbf{11.1^{\pm 3.32}}$ $\mathbf{4.7^{\pm 5.11}}$
+ViSQOL $3.9^{\pm 0.35}$ $\mathbf{4.2^{\pm 0.46}}$ $\mathbf{4.5^{\pm 0.32}}$ $\mathbf{4.1^{\pm 0.39}}$
+CodecFormer SI-SDR $47.6^{\pm 9.51}$ $47.1^{\pm 10.97}$ $47.8^{\pm 9.65}$ $48.2^{\pm 9.87}$
+ViSQOL $^{\pm 0.07} 1.0±0.1$$^{\pm 0.12}$ $1.0^{\pm 0.06}$ $1.2^{\pm 0.47}$
+CodecSep + dnr-v2 SI-SDR $3.4^{\pm 1.85}$ $4.1^{\pm 3.97}$ $6.2^{\pm 2.87}$ $0.8^{\pm 5.16}$
+(unguided, 3-stem) ViSQOL $3.2^{\pm 0.20}$ $3.0^{\pm 0.33}$ $3.5^{\pm 0.24}$ $3.2^{\pm 0.46}$
+SDCodec SI-SDR $\mathbf{7.0^{\pm 2.49}}$ $7.7^{\pm 4.60}$ $8.3^{\pm 3.26}$ $2.5^{\pm 5.65}$
+ViSQOL $\mathbf{4.3^{\pm 0.15}}$ $0^{\pm 0.28} 4.4$$4^{\pm 0.15}$ $4.0^{\pm 0.34}$
 Text-guided
-AudioSep SI-SDR $5.53^{\pm 1.96}$ $4.69^{\pm 5.36}$ $10.97^{\pm 2.99}$ $1.99^{\pm 5.68}$
-(zero-shot) ViSQOL $4.06^{\pm 0.38}$ $3.75^{\pm 0.65}$ $4.57^{\pm 0.13}$ $3.19^{\pm 0.77}$
-AudioSep + dnr-v2 SI-SDR $6.47^{\pm 2.26}$ $7.99^{\pm 4.55}$ $8.13^{\pm 3.35}$ $2.29^{\pm 5.95}$
-ViSQOL $4.20^{\pm 0.18}$ $4.12^{\pm 0.21}$ $3.03^{\pm 0.29}$ $3.81^{\pm 0.47}$
-CodecSep + dnr-v2 SI-SDR $4.06^{\pm 2.06}$ $3.91^{\pm 3.93}$ $6.10^{\pm 2.86}$ $0.67^{\pm 5.29}$
-ViSQOL $74^{\pm 0.22} 3.3$$37^{\pm 0.33}$ $3.83^{\pm 0.24}$ $3.54^{\pm 0.44}$
-CodecSep + dnr-v2 SI-SDR $\mathbf{12.24^{\pm 2.42}}$ $\mathbf{12.58^{\pm 3.81}}$ $\mathbf{13.59^{\pm 2.59}}$ $\mathbf{8.67^{\pm 4.17}}$
-(ablate Masker) ViSQOL $\mathbf{4.44^{\pm 0.14}}$ $\mathbf{4.13^{\pm 0.31}}$ $\mathbf{3.85^{\pm 0.34}}$ $3.76^{\pm 0.54}$
-AudioSep + AudioCaps SI-SDR $6.67^{\pm 2.52}$ $8.10^{\pm 4.63}$ $8.39^{\pm 3.21}$ $2.42^{\pm 6.12}$
-(zero-shot) ViSQOL $4.23^{\pm 0.19}$ $4.12^{\pm 0.21}$ $4.17^{\pm 0.21}$ $\mathbf{3.82^{\pm 0.46}}$
-CodecSep + AudioCaps SI-SDR $0.59^{\pm 1.89}$ $0.24^{\pm 5.15}$ $11.03^{\pm 5.21}$ $1.23^{\pm 4.84}$
-(zero-shot) ViSQOL $3.33^{\pm 0.23}$ $2.91^{\pm 0.64}$ $1.69^{\pm 0.48}$ $3.44^{\pm 0.41}$
+AudioSep SI-SDR $5.5^{\pm 1.96}$ $4.7^{\pm 5.36}$ $11.0^{\pm 2.99}$ $2.0^{\pm 5.68}$
+(zero-shot) ViSQOL $4.1^{\pm 0.38}$ $3.8^{\pm 0.65}$ $4.6^{\pm 0.13}$ $3.2^{\pm 0.77}$
+AudioSep + dnr-v2 SI-SDR $6.5^{\pm 2.26}$ $8.0^{\pm 4.55}$ $8.1^{\pm 3.35}$ $2.3^{\pm 5.95}$
+ViSQOL $4.2^{\pm 0.18}$ $4.1^{\pm 0.21}$ $3.0^{\pm 0.29}$ $3.8^{\pm 0.47}$
+CodecSep + dnr-v2 SI-SDR $4.1^{\pm 2.06}$ $3.9^{\pm 3.93}$ $6.1^{\pm 2.86}$ $0.7^{\pm 5.29}$
+ViSQOL $3.7^{\pm 0.22}$ $3.4^{\pm 0.33}$ $3.8^{\pm 0.24}$ $3.5^{\pm 0.44}$
+CodecSep + dnr-v2 SI-SDR $\mathbf{12.2^{\pm 2.42}}$ $\mathbf{12.6^{\pm 3.81}}$ $\mathbf{13.6^{\pm 2.59}}$ $\mathbf{8.7^{\pm 4.17}}$
+(ablate Masker) ViSQOL $\mathbf{4.4^{\pm 0.14}}$ $\mathbf{4.1^{\pm 0.31}}$ $\mathbf{3.9^{\pm 0.34}}$ $3.8^{\pm 0.54}$
+AudioSep + AudioCaps SI-SDR $6.7^{\pm 2.52}$ $8.1^{\pm 4.63}$ $8.4^{\pm 3.21}$ $2.4^{\pm 6.12}$
+(zero-shot) ViSQOL $4.2^{\pm 0.19}$ $4.1^{\pm 0.21}$ $4.2^{\pm 0.21}$ $\mathbf{3.8^{\pm 0.46}}$
+CodecSep + AudioCaps SI-SDR $0.6^{\pm 1.89}$ $0.2^{\pm 5.15}$ $11.^{\pm 5.21}$ $1.2^{\pm 4.84}$
+(zero-shot) ViSQOL $3.3^{\pm 0.23}$ $2.9^{\pm 0.64}$ $1.7^{\pm 0.48}$ $3.4^{\pm 0.41}$